### v1

# The Unpaid Toll: Quantifying the Public Health Impact of AI

Abstract
The surging demand for AI has led to a rapid expansion of energy-intensive data centers, impacting the environment through escalating carbon emissions and water consumption. While significant attention has been paid to AI’s growing environmental footprint, the public health burden, a hidden toll of AI, has been largely overlooked. Specifically, AI’s lifecycle, from chip manufacturing to data center operation, significantly degrades air quality through emissions of criteria air pollutants such as fine particulate matter, substantially impacting public health. This paper introduces a methodology to model pollutant emissions across AI’s lifecycle, quantifying the public health impacts. Our findings reveal that training an AI model of the Llama-3.1 scale can produce air pollutants equivalent to more than 10,000 round trips by car between Los Angeles and New York City. The total public health burden of U.S. data centers in 2030 is valued at up to more than \$20 billion per year, double that of U.S. coal-based steelmaking and comparable to that of on-road emissions of California. Further, the public health costs unevenly impact economically-disadvantaged communities, where the per-household health burden could be 200x more than that in less-impacted communities. We recommend adopting a standard reporting protocol for criteria air pollutants and the public health costs of AI, paying attention to all impacted communities, and implementing health-informed AI to mitigate adverse effects while promoting public health equity.
1 Introduction
The rise of artificial intelligence (AI) has numerous potentials to play a transformative role in addressing grand societal challenges, including air quality and public health [1, 2]. For example, by integrating multimodal data from various sources, AI can provide effective tools and actionable insights for pandemic preparedness, disease prevention, healthcare optimization, and air quality management [3, 1]. However, the surging demand for AI — particularly generative AI, as exemplified by the recent popularity of large language models (LLMs) — has driven a rapid increase in computational needs, fueling the unprecedented expansion of energy-intensive AI data centers. According to McKinsey projections, under a medium-growth scenario [4], the U.S. data centers are anticipated to account for 11.7% of national electricity consumption in 2030, a substantial increase from their current share of less than 4% in 2023.
The growing electricity demand of AI data centers has not only created significant stress on power grid stability [5, 6], but also increasingly impacts the environment through escalating carbon emissions [7, 8] and water consumption [9]. These environmental impacts are driven primarily by the “expansion of AI products and services,” as recently acknowledged by Google in its latest sustainability report [10]. To mitigate the challenges posed to both power grids and the environment, a range of strategies have been explored, including grid-integrated data centers [6, 11], energy-efficient hardware and software [12, 13, 14], and the adoption of carbon-aware and water-efficient computing practices [15, 9, 16, 17], among others.
The hidden toll of AI.
While the environmental footprint of AI has garnered attention, the public health burden, a hidden toll of AI, has been largely overlooked. Across its entire lifecycle — from chip manufacturing to data center operation — AI contributes substantially to air quality degradation and public health costs through the emission of various criteria air pollutants. These include fine particulate matter (PM2.5, particles measuring 2.5 micrometers or smaller in diameter that can penetrate deep into lungs and cause serious health effects), sulfur dioxide (SO2), and nitrogen dioxide (NO2). Concretely, the AI hardware manufacturing process [18], electricity generation from fossil fuels to power AI data centers, and the maintenance and usage of diesel backup generators to ensure continuous AI data center operation all produce significant amounts of criteria air pollutants. Moreover, the distinct spatial-temporal heterogeneities of emission sources suggest that focusing solely on reducing AI’s carbon footprints may not minimize its emissions of criteria air pollutants or the resulting public health impacts (Section 5).
Exposure to criteria air pollutants is directly and causally linked to various adverse health outcomes, [^2] While we focus on public health, we note that the impacts of criteria air pollutants extend beyond humans and include harms to environmentally sensitive areas, such as some national parks and wilderness areas which, classified as “Class 1 areas” under the Clean Air Act, require special air protection [19]. including premature mortality, lung cancer, asthma, heart attacks, cardiovascular diseases, strokes, and even cognitive decline, especially for the elderly and vulnerable individuals with pre-existing conditions [20, 21, 22, 23]. Moreover, even short-term (hours to days) PM2.5 exposure is harmful and deadly, accounting for approximately 1 million premature deaths per year from 2000 to 2019 and representing 2% of total global deaths [24].
Globally, 4.2 million deaths were attributed to ambient (i.e., outdoor) air pollution in 2019 [25]. Air pollution has become the second highest risk factor for noncommunicable diseases [26]. Notably, according to the latest Global Burden of Disease report [27], along with high blood pressure and high blood sugar, ambient particulate matter is placed among the leading risk factors for disease burden globally in every socio-demographic group.
While the U.S. has generally better air quality than many other countries, 4 in 10 people in the U.S. still live with unhealthy levels of air pollution, according to the “State of the Air 2024” report published by the American Lung Association [28]. In 2019 (the latest year of data provided by the World Health Organization, or WHO, as of November 2024), an estimate of 93,886 deaths in the U.S. were attributed to ambient air pollution [29]. In fact, even compliance with the U.S. Environmental Protection Agency (EPA) air quality standards does not necessarily guarantee healthy air that meets the WHO guidelines. Concretely, the EPA’s recently tightened primary standard for PM2.5 sets an annual average limit of 9 $\mu g/m^{3}$, considerably higher than the WHO’s recommended level of 5 $\mu g/m^{3}$ [30, 31]. In addition, the EPA projects that 53 U.S. counties, including 23 in the most populous state of California, would fail to meet the revised national annual PM2.5 standard in 2032 [32].
Further, criteria air pollutants are not confined to the immediate vicinity of their emission sources; they can travel hundreds of miles through a dispersion process (i.e., cross-state air pollution) [33, 34], impacting public health across vast regions — pollutants from the 2024 Canadian wildfires significantly degraded air quality across much of the U.S. and reached as far as Mexico and Europe [35].
Importantly, along with transportation and industrial activities, electricity generation is a major contributor to ambient air pollution with substantial public health impacts [26, 36, 37]. For example, a recent study [38] shows that, between 1999 and 2020, a total of 460,000 excess deaths were attributed to PM2.5 generated by coal-fired power plants alone in the U.S. As highlighted by the U.S. EPA [36], despite years of progress, “fossil fuel-based power plants remain a leading source of air, water, and land pollution that affects communities nationwide.” Moreover, according to the U.S. Energy Information Administration (EIA) projection [39], the coal consumption by the electricity sector in 2050 will still be about 30% of the 2024 level in the baseline reference case, and the number will exceed 50% in the high zero-carbon technology cost case. Indeed, the growing energy demands of AI are already delaying the decommissioning of coal-fired power plants and increasing fossil-fuel plants in the U.S. as well as around the world [6, 40, 41].
The public health outcomes of AI due to its emission of criteria air pollutants lead to various losses, such as hospitalizations, medication usage, emergency room visits, school loss days, and lost workdays. Moreover, these losses can be further quantified in economic costs based on epidemiology and economics research for the corresponding health endpoints [42, 22]. In contrast, the environmental impacts of AI, e.g., carbon emission from fossil fuels and water consumption for data center cooling, often do not cause the same immediate health impacts. For instance, while anthropogenic carbon emissions could also pose risks to public health, such impacts are often second- or third-order effects through long-term climate change which can then threaten the human well-being by affecting the food people eat and facilitating the spreading of pests, among others [43]. Nonetheless, despite their immediate and tangible impacts on public health, the criteria air pollutants of AI have remained under the radar, entirely omitted from today’s AI risk assessments and sustainability reports [44, 45, 10].
Quantifying the public health costs of AI.
In this paper, we uncover and quantify the hidden public health impacts of AI. We introduce a general methodology to model the emission of criteria air pollutants associated with AI tasks across three distinct scopes: emissions from the maintenance and operation of backup generators (Scope 1), emissions from fossil fuel combustion for electricity generation (Scope 2), and emissions resulting from the manufacturing of server hardware (Scope 3). Then, we analyze the dispersion of criteria air pollutants and the resulting public health impacts across different regions.
Our main results (Section 4) focus on the scope-2 health impacts of U.S. data centers and, specifically, LLM training. [^3] Our study focuses on the 48 contiguous U.S. states plus Washington D.C. because the EPA data does not include other regions [46]. If located in countries with higher population densities or less strict air quality standards, the same AI task and data centers would likely contribute to significantly more deaths and other adverse health effects. We recommend further research on the public health impact of AI outside the U.S. Using the reduced-complexity modeling tool COBRA (CO-Benefits Risk Assessment) provided by the EPA [46], our screening analysis demonstrates that driven by the growing demand for AI, the U.S. data centers could contribute to, among others, approximately 600,000 asthma symptom cases and 1,300 premature deaths in 2030, exceeding 1/3 of asthma deaths in the U.S. each year [47]. The overall public health costs could reach more than \$20 billion, double that of the U.S. coal-based steelmaking industry [48], and rival or even top those of on-road emissions of the largest U.S. states such as California with $\sim$35 million registered vehicles [49]. Moreover, depending on the location, training an AI model of the Llama-3.1 scale can produce an amount of air pollutants equivalent to driving a car for more than 10,000 round trips between Los Angeles and New York City (LA-NYC), resulting in a health cost that even exceeds 120% of the training electricity cost.
Critically, the health costs are unevenly distributed across counties and communities, disproportionately affecting low-income counties (e.g., Meigs County, Ohio) where the per-household health burden could be equivalent to nearly 8 months of electricity bills and more than 200x compared to that in other counties.
In addition, to highlight the importance of scope-1 and scope-3 health impacts, we consider data center backup generators in Virginia (Scope 1) and semiconductor manufacturing plants in Arizona and Ohio (Scope 3). Our analysis shows that, assuming the actual emissions are only 10% of the permitted level, the data center backup generators registered in Virginia (mostly in Loudoun, Prince William, and Fairfax) could already cause 14,000 asthma symptom cases among other health outcomes and a total public health burden of \$220-300 million per year, impacting residents in multiple surrounding states and as far as Florida (Section 2.2.1). If these data centers emit air pollutants at the maximum permitted level, the total public health cost will become 10-fold and reach \$2.2-3.0 billion per year. The scope-3 health impact of AI is also substantial. For example, just a single semiconductor facility in Arizona can cause an annual public health cost of \$26-39 million, with \$14-21 million attributed to the facility’s on-site emissions of criteria air pollutants (Section 2.2.2). Furthermore, relocating the same facility to a planned site in Ohio could almost quadruple the public health cost to \$94-156 million, with \$23-36 million resulting from on-site emissions.
Finally, we provide recommendations to address the increasing public health impact of AI (Section 5). Specifically, we recommend technology companies adopt a standard reporting protocol for criteria air pollutants and public health impacts in their AI model cards and sustainability reports, implement health-informed AI to proactively minimize the adverse health effects of AI data centers, pay attention to all impacted communities, and prioritize reducing the health impact on disadvantaged communities to promote public health equity.
To summarize, our study sheds light on and quantifies the overlooked public health impact of AI. It can inform the public, policymakers, and technology companies in conducting a more comprehensive cost-benefit analysis. We also urge further research to comprehensively address the public health implications when developing powerful and truly responsible AI in the future, ensuring that the growth of AI does not exacerbate the health burden or outweigh the potential benefits AI can provide to improve public health.
2 Background on the Air Quality Impact of AI
This section presents an overview of AI’s impact on air quality and contribution to criteria air pollutants throughout its lifecycle, beginning with background on criteria air pollutants and U.S. air quality policies.
2.1 Criteria Air Pollutants
Criteria air pollutants, including PM2.5, SO2 and NO2, are a group of airborne contaminants that are emitted from various sources such as industrial activities and vehicle emissions. The direct emission of PM2.5 is called primary PM2.5, while precursor pollutants such as SO2, NOx, and VOCs, can form secondary PM2.5 and/or ozones [50]. These air pollutants can travel a long distance (a.k.a. cross-state air pollution), posing direct and significant risks to public health over large areas, particularly for vulnerable populations including the elderly and individuals with respiratory conditions [33, 34].
Long-term exposure to PM2.5, even at a low level, are directly linked to numerous health outcomes, including premature mortality, heart attacks, asthma, stroke, lung cancer, and even cognitive decline [21, 22]. These health effects result in various losses, such as hospitalizations, medication usage, emergency room visits, school loss days, and lost workdays, which can be further quantified in economic costs based on public health research for various health endpoints [42]. In addition, short-term (hours to days) PM2.5 exposure is also dangerous, contributing to approximately 1 million premature deaths per year globally from 2000 to 2019 [24].
[…]
While CO2 is broadly classified by the EPA as an air pollutant following the U.S. Supreme Court ruling in 2007 [53] and contributes to long-term climate change, it often does not cause the same immediate health impacts as criteria pollutants. In the U.S., CO2 and other greenhouse gases are subject to different EPA regulations from those for criteria air pollutants. Thus, for the sake of presentation in this paper, we use “air pollutants” to solely refer to criteria air pollutants wherever applicable.
2.2 AI’s Contribution to Air Pollutants
To understand the impact of AI on air quality, we focus on the three scopes over which AI contributes to criteria air pollutants as well as other toxic materials. The scoping definition in this paper parallels the well-established greenhouse gas protocol [54].
2.2.1 Scope 1
The scope-1 public health impact of AI primarily comes from the emission of operating on-site backup generators. Data centers are mission-critical facilities that are designed to operate with high availability and uptime guarantees. As a result, to maintain operation during emergencies such as grid outages, AI data centers require highly reliable backup power sources [10, 45]. Diesel generators are known to emit significant amounts of air pollutants and even hazardous emissions during operation. For example, they emit 200-600 times more NOx than new or controlled existing natural gas-fired power plants for each unit of electricity produced [55]. Nonetheless, there is limited experience with cleaner backup alternatives that can provide comparable reliability in real-world settings, as highlighted by the U.S. Department of Energy in its recent recommendations regarding AI data center infrastructures [6]. Consequently, AI data centers, including those newly built by major technology companies, primarily depend on on-site diesel generators for backup power [6, 10, 45, 56]. For example, in northern Virginia (mostly in Loudoun, Prince William, and Fairfax), the number of permits for data center diesel generators has increased by about 70% since 2023 compared to the total number of permits issued between 2000 and 2022 [56].
(a) Health cost map
(b) CDF of county-wide health cost
State County Health Cost (million \$) MD Montgomery 19.9 (17.3, 22.4) VA Fairfax 18.9 (16.6, 21.2) MD Prince Georges 8.9 (7.5, 10.4) MD Baltimore 8.3 (7.0, 9.6) DC District of Columbia 7.6 (6.2, 9.0) MD Anne Arundel 6.3 (5.5, 7.2) MD Baltimore City 6.0 (4.8, 7.1) VA Loudoun 5.4 (4.7, 6.1) VA Prince William 5.0 (4.4, 5.7) MD Frederick 4.6 (3.9, 5.2)
(c) Top-10 counties by health cost
Figure 1: The county-level total scope-1 health cost of data center backup generators operated in Virginia (mostly in Loudoun County, Fairfax County, and Prince William County) [57]. The backup generators are assumed to emit air pollutants at 10% of the permitted levels per year. The total annual public health cost is \$220-300 million, including \$190-260 million incurred in Virginia, West Virginia, Maryland, Pennsylvania, New York, New Jersey, Delaware, and Washington D.C. (a) County-level health cost in Virginia, West Virginia, Maryland, Pennsylvania, New York, New Jersey, Delaware, and Washington D.C. Counties with data centers are marked in orange, except for Loudoun County (marked in yellow). (b) CDF of the county-level cost. (c) Top-10 counties by the total health cost.
While diesel generators need to comply with air quality regulations and typically do not operate over extended periods of time, regular maintenance and testing are essential to ensure their operational reliability. In addition, capacity redundancy is typically followed for diesel generator installations to ensure high availability [58]. Thus, diesel generators represent a major source of on-site air pollutants for data centers and pose a significant health risk to the public [59]. For instance, the total permitted annual emission limits for data centers in northern Virginia are approximately 13,000 tons of NOx, 1,400 tons of VOCs, 50 tons of SO2, and 600 tons of PM2.5, all in U.S. short tons. Assuming that the actual emissions are only 10% of the permitted level, these backup generators could already cause 14,000 asthma symptom cases and 13-19 deaths each year among other health implications, resulting in a total annual public health burden of \$220-300 million throughout the U.S. This includes \$190-260 million in Virginia, West Virginia, Maryland, Pennsylvania, Delaware, New Jersey, New York, and Washington D.C. We show the county-level health cost and the top-10 counties in Figure 1, while deferring the details of calculations to Appendix A.3.
Moreover, due to power grid capacity constraints in many U.S. states, AI data centers are increasingly pressured to vary their loads subject to the grid’s operating conditions, i.e., grid-integrated data centers [60, 6]. This trend may necessitate extended reliance on backup generators, e.g., possibly 15 days per year [6]. Such prolonged usage of diesel generators could substantially elevate AI’s scope-1 air pollution, creating even higher public health costs. Concretely, if the backup generators in northern Virginia emit air pollutants at the maximum permitted level, the total public health cost could reach \$2.2-3.0 billion per year.
What further adds to the public health threat is that many data center generators in a region may operate simultaneously for demand response during grid capacity shortages, potentially resulting in a short-term spike in PM2.5 and NOx emissions that can be particularly harmful [6, 24, 31].
2.2.2 Scope 2
While technology companies have started implementing various initiatives — such as purchasing renewable energy credits and nuclear power from small modular reactors [5, 10, 61] — to lower their (market-based) carbon emissions, the vast majority of U.S. data centers remain directly powered by local power grids with a substantial portion of fossil fuel-based energy sources [10]. Thus, just as AI is accountable for scope-2 carbon emissions, it also contributes to scope-2 air pollution through its electricity usage.
The combustion of fossil fuels for electricity production is a major emitter of criteria air pollutants, releasing large amounts of PM2.5, SO2, NOx, VOCs, and others. [^4] Wet cooling towers, including those used by data centers [9, 10] and carbon-free nuclear power plants, rely on water evaporation for heat rejection and produce PM2.5 due to spray drift droplets [62, 63]. Nonetheless, because of limited data available, we exclude the cooling tower PM2.5 emission from our analysis unless other specified. Critically, the growing energy demands of AI are already delaying the decommissioning of coal-fired power plants and increasing fossil-fuel plants in the U.S. and other countries [6, 40]. For example, in addition to keeping 2,099 MW coal generation capacity until 2039 (more than 80% of the 2024 level), Virginia Electric and Power Company plans to install 5,934 MW gas-fired plants to meet the growing energy demand driven by AI data centers [41]. At the national level, per the EIA’s projection, [39], the 2050 natural gas consumption for U.S. electricity generation will be about 80% of the 2024 level in the baseline reference case, and even exceed the 2024 level by 20% if the zero-carbon technology cost is high; for coal consumption by the electricity sector in 2050, the numbers will also be considerably high, about 30% and over 50% of the 2024 level in the baseline reference case and in the high zero-carbon technology cost case, respectively. These projections were published by the EIA at the very beginning of the generative AI boom in early 2023. More recently, it has been reported that AI data centers could even be primarily powered by coal power plants in some countries [40]. As a result, AI’s scope-2 air pollution is expected to remain at a high level for a substantially long time into the future.
We also note that the practice of using various credits to offset scope-2 carbon emissions [10] may not be effective for mitigating the scope-2 public health impact. The reason is that the public health impact of using grid electricity is highly location-dependent, e.g., the impact in a populated region may not be mitigated by renewable energy generated elsewhere.
2.2.3 Scope 3
The surging demand for AI necessitates large quantities of computational hardware, including graphics processing units (GPUs), thus intensifying the supply chain requirements [64]. However, semiconductor manufacturing generates various criteria air pollutants, wastewater, toxic materials, and hazardous air emissions [18]. Moreover, the energy-intensive nature of semiconductor production further contributes to pollutants from power plants. Combined with other pollution sources such as transportation and electronic waste recycling [65], the supply chain activities form a large portion of AI’s scope-3 impact on public health.
Although semiconductor manufacturing facilities are subject to air quality regulations [66], they still pose significant risks, affecting populations across large regions. Maricopa County, AZ, has been an EPA-designated non-attainment area for several years due to its failures to meet federal air quality standards [67]. The establishment of multiple semiconductor facilities in such areas could further exacerbate air quality issues. In 2023–2024, the estimated annual public health impact of a single semiconductor facility was \$26-39 million, with \$14-21 million attributed to direct on-site emissions of air pollutants from the facility, based on COBRA estimates [18, 46]. Moreover, relocating the facility to a planned site in Licking County, Ohio, could nearly quadruple public health costs to \$94-156 million, with \$23-36 million resulting from direct on-site emissions. This increase is partly due to Ohio’s weather conditions and higher reliance on coal-based power [68]. The details of calculations are available in Appendix A.4. Importantly, the global demand for AI chips in 2030 is projected to be tens of times of the overall production capacity of this single facility [69], further magnifying the overall scope-3 public health impact of AI. It is also worth noting that additional pollutants, including hazardous air pollutants like hydrogen fluoride, may further elevate public health costs but are not included in this analysis.
3 Quantifying the Public Health Impact of AI
To quantify the public health impact of AI, we present a general methodology that quantifies AI’s criteria air pollutants at the emission source, models its dispersed air pollutants at different receptors (i.e., destination regions), and finally obtains the public health impact and cost at each receptor.
For an AI task (e.g., AI model training), we consider $M$ types of criteria air pollutants, $N$ receptor regions of interest (e.g., all the U.S. counties), $H$ types of public health impacts (e.g., mortality, asthma symptoms, school loss days, etc.). We use $p^{s}=(p_{1}^{s},\cdots,p_{M}^{s})$ and $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ denote the quantities for $M$ types of air pollutants attributed to the task at the emission source and at the receptor $i$, respectively, for $i=1,\cdots,N$. Additionally, we use $h_{i}=(h_{i,1},\cdots,h_{i,H})$ and $c_{i}=(c_{i,1},\cdots,c_{i,H})$ to denote the incidences and economic costs associated with $H$ types of health impacts at receptor $i$, respectively, for $i=1,\cdots,N$. With a slight abuse of notations, we reuse these symbols when modeling AI’s public health impacts across the three different scopes.
3.1 Criteria Air Pollutants at the Source
We first model AI’s criteria air pollutants at the source across the three different scopes in Section 2.2.
3.1.1 Scope 1
On-site backup diesel generators are sized based on the data center power capacity and routinely tested to ensure a high availability of the entire data center. Thus, the overall scope-1 air pollutants should be attributed to each computing task based on its power allocation and duration. Suppose that the overall scope-1 emission by an AI data center under consideration is $\bar{p}^{s}=(\bar{p}_{1}^{s},\cdots,\bar{p}_{M}^{s})$, for $M$ types of air pollutants, over a timespan of $\overline{T}$ (e.g., one year). Considering an AI task that is allocated a fraction of $x\in(0,1]$ of the overall data center power capacity and lasts for a duration of $T$, we express the scope-1 air pollutants attributed to the AI task as
$p^{s}=\frac{x\cdot T}{\overline{T}}\cdot\bar{p}^{s},$ (1)
which attributes the overall emission $\bar{p}^{s}$ to the task in proportion to its allocated power and duration.
3.1.2 Scope 2
AI’s scope-2 air pollutants come from its usage of electricity generated from fossil fuels. Suppose that the power grid serving the AI data center has an emission rate of $\gamma=(\gamma_{1},\cdots,\gamma_{M})$ for $M$ types of air pollutants to produce each unit of electricity. In practice, the power grid consists of multiple interconnected power plants to supply electricity to many customers over a wide area (e.g., a balancing area [70]). Thus, similar to carbon footprint accounting [71], the air pollutant emission rate $\gamma$ can be calculated based on either the weighted average emission rate of all the power plants (i.e., $\gamma=\frac{\sum_{k}\gamma_{k}\cdot b_{k}}{\sum_{k}\cdot b_{k}}$ where $\gamma_{k}$ and $b_{k}$ are the emission rate and electricity generation of the power plant $k$) or the emission rate of the marginal power plant (i.e., the power plant dispatched in response to the next electricity demand increment), which are referred to as average emission rate or marginal emission rate, respectively. The average emission represents a proportional share of the overall air pollutant emission by an electricity consumer, while the marginal emission is useful for quantifying the additionality of air pollutants due to a consumer’s electricity usage.
Suppose that the electricity consumption by the AI task is $e$, including the data center overhead captured by the power usage effectiveness. Then, we can write the scope-2 air pollutants as
$p^{s}=e\cdot\gamma,$ (2)
which is either based on either average attribution or marginal attribution. While the marginal emission is typically associated with a single marginal power plant, the average emission is spread across all the interconnected power plants within a wide area such as a power balancing area [71, 70]. Thus, when considering the average attribution method, we split the energy consumption $e$ over all the power plants in proportion to their contributions to the grid’s supply and calculate the corresponding per-plant emission. In other words, each involved power plant is an individual pollution source, and the air pollutant emission at the $k$-th power plant is $p_{k}^{s}=e\cdot\frac{b_{k}}{\sum_{k}b_{k}}\cdot\gamma_{k}$, where $b_{k}$ is the electricity generation of the $k$-th power plant.
Since both the average and marginal air pollutant emission rates vary over time and locations to meet the supply-demand balance, we can also refine the calculation of scope-2 air pollutants in (2) by considering the summation of air pollutants over multiple time slots over the AI task’s duration.
3.1.3 Scope 3
Following the attribution method for scope-3 carbon emission and water consumption [13, 9], we attribute the AI hardware’s air pollutants during the manufacturing process to a specific task based on the task duration. Specifically, let the AI hardware’s expected lifespan be $\overline{T}_{0}$ and the AI task lasts a duration of $T$. Considering that the $M$ types of air pollutants for manufacturing the AI hardware are $\bar{p}_{0}^{s}=(\bar{p}_{0,1}^{s},\cdots,\bar{p}_{0,M}^{s})$ and excluding other miscellaneous pollutants (e.g., transportation), we obtain AI’s scope-3 air pollutants as
$p^{s}=\frac{T}{\overline{T}_{0}}\cdot\bar{p}_{0}^{s}.$ (3)
As an AI server cluster includes multiple hardware components (e.g., GPU and CPU) manufactured in different locations, we apply (3) to estimate the scope-3 air pollutants for each component manufactured in a different location.
3.2 Air Quality Dispersion Modeling
Once emitted from their sources, criteria air pollutants can travel long distances, impacting multiple states along their paths. Unlike carbon emissions that have a similar effect on climate change regardless of the emission source locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source. Generally, the closer a receptor is to the source, the greater the air quality impact. Furthermore, the dispersion of air pollutants is influenced by meteorological conditions, such as wind speed and direction.
The movement of air pollutants can be modeled using mathematical equations to simulate the atmospheric processes governing the dispersion, known as dispersion modeling. By incorporating emission data and meteorological inputs, dispersion modeling can predict pollutant concentrations at selected receptor locations [72]. We consider a general dispersion model $(p^{r}_{1},\cdots,p^{r}_{N})=D_{\theta}(p^{s})$, which yields the amount of $M$ types of air pollutants $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ at the receptor region $i=1,\cdots,N$. The parameter $\theta$ captures the geographical conditions, emission source characteristics (e.g., height), and meteorological data if applicable [73]. We apply the dispersion model to each scope of air pollutants (Section 3.1) to estimate the corresponding pollutant concentrations at receptor regions.
In practice, many dispersion modeling tools are available, including AERMOD, CTDMPLUS, PCAPS and InMAP with a reduced complexity [72, 22, 74, 75]. For example, PCAPS (Pattern Constructed Air Pollution Surfaces), an advanced reduced-complexity model that provides representations of both primarily emitted PM2.5 and secondarily formed PM2.5 and ozone, is used in COBRA as a quick assessment of otherwise lengthy iterations and simulations of various pollution scenarios in terms of the annual average PM2.5 and seasonal average maximum daily average 8-hour ozone [75, 22]. Even compared with state-of-the-science photochemical grid models, PCAPS provides similar prediction accuracies and can realistically capture the change in air pollution due to changing emissions [75]. More specifically, for electric power sectors and on-road/highway vehicle sectors (the two sectors we consider in Section 4), the prediction results of PCAPS compare very well with photochemical model predictions, with Pearson correlation coefficients of 0.92 and 0.94, respectively [22, 75].
3.3 Converting Health Outcomes to Economic Costs
By assessing pollutant levels $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ and population size at each receptor region $i$, we can estimate the incidences of health outcomes $h_{i}=(h_{i,1},\cdots,h_{i,H})$ and the corresponding public health cost $c_{i}=(c_{i,1},\cdots,c_{i,H})$. The relations between $p_{i}^{r}$ and $h_{i}$ and between $h_{i}$ and $c_{i}$ can be established based on epidemiology research [22]. For example, the premature mortality rate can be modeled as a log-linear function in terms of the PM2.5 level [23].
Further, by summing up the economic costs, we obtain quantitative estimates of the public health burden at both regional and national levels. It is important to note that the public health cost is not necessarily an out-of-pocket expense incurred by each individual, but rather reflects the estimated economic burden on a population to mitigate the adverse effects of pollutants within a specific region. Therefore, it is a quantitative scalar measure of the public health impact resulting from a particular pollutant-producing activity.
3.4 Implementation
We now briefly describe the specific implementation we use to study the public health impact of U.S. data centers and AI training. The details are available in Appendix A.
Due to the limited data available for scope-1 and scope-3 impacts, we mainly focus on the scope-2 health impacts from electricity consumption. To account for future uncertainties, we use the U.S. data center electricity consumption data provided by EPRI [5] and McKinsey [4] under various growth-rate scenarios, excluding cryptocurrency servers. Unless otherwise specified, we consider the average attribution method by default, i.e., attribute the overall health impact within an electricity region to data centers in proportion to their electricity consumption.
To model the air pollutant dispersion and quantify health impacts, we use the latest COBRA (Desktop v5.1, as of October 2024) provided by the U.S. EPA [46]. COBRA integrates reduced-complexity air dispersion modeling (including both primarily emitted PM2.5 and secondly formed PM2.5 and ozone [75]) with various concentration-response functions [22], offering a quantitative screening analysis particularly suitable for large-scale health impacts. The same or similar reduced-complexity modeling tools have been commonly used in the literature to examine the health impacts of various industries over a large area [76, 74], including electric vehicles [77], bitcoin mining [78], and inter-region electricity imports [79], among others. While each health impact model used by COBRA considers 95% confidence intervals, the high-end and low-end estimates provided by COBRA are based on different models instead of the 95% confidence interval of a single model [22]. COBRA provides data for county-level population, health incidence, and valuation projections in 2030, but the baseline emissions are missing [46]. Thus, to account for model uncertainties, we estimate the 2030 baseline emission by extrapolating the COBRA data for 2016, 2023, and 2028 using three extrapolation methods (Linear, Exponential, and Unchanged) as detailed in Appendix A.1.
We only consider the contiguous U.S. and simply refer to it as the U.S. For consistency with COBRA, cities considered county-equivalents for census purposes are also referred to as “counties” in our paper. All our monetary values are for one year (or one AI task if applicable) and in 2023 U.S. dollars.
4 Results
We now present our estimates of the public health impacts caused by the U.S. data centers in aggregate and by training a large generative AI model at specific locations. Our results demonstrate that in 2030, the scope-2 pollutants of U.S. data centers alone could cause, among others, approximately 600,000 asthma symptom cases and 1,300 premature deaths, exceeding 1/3 of asthma deaths in the U.S. each year [47]. The overall public health costs of U.S. data centers could rival or even exceed those of on-road emissions of the largest U.S. states such as California. Moreover, depending on the locations, training an AI model of the Llama-3.1 scale can produce an amount of air pollutants equivalent to driving a passenger car for more than 10,000 LA-NYC round trips, resulting in a health cost that even exceeds 120% of the training electricity cost. Importantly, the health costs are disproportionately distributed across counties and communities, particularly affecting low-income counties that could experience more than 200x per-household health costs than others.
4.1 Public Health Impact of U.S. Data Centers in 2023
We first show in Table 1 the public health cost of U.S. data centers in 2023 as a reference. [^5] We use the “mid (low, high)” format to represent the midrange, low and high estimates offered by COBRA. When presenting a single value or a ratio (e.g., health-to-electricity cost ratio), we use the midrange by default. Even at the beginning of the generative AI boom, the U.S. data centers have already resulted in a total public health cost of about \$5.6 billion, or \$39.7 per household, in 2023. This is equivalent to 43% of the data centers’ total electricity cost. By considering marginal attribution, the U.S. data centers’ public health cost increases to about \$7.6 billion in 2023, due to the heavy reliance on fossil fuels by many marginal generators [70]. This suggests that, by powering the U.S. data centers using alternative energy sources (e.g., geothermal) off the main grid, the U.S. could have seen a public health benefit of \$7.6 billion in 2023. Additional results can be found in Appendix B, including county-wide total and per-household health costs that demonstrate the uneven distribution of health impacts across different communities.
Table 1: The public health cost of U.S. data centers in 2023.
Attribution Method Electricity (TWh) Electricity Cost (billion \$) Mortality Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$) Months of Electricity Bill % of CA On-road Health Cost
Average 152.1 13.0 360 (270, 460) 5.6 (4.2, 7.0) 43% 39.7 (29.6, 49.8) 0.29 (0.22, 0.36) 35%
Marginal 152.1 13.0 490 (360, 620) 7.6 (5.7, 9.4) 58% 53.8 (40.5, 67.2) 0.39 (0.30, 0.49) 47%
Mobile sources, including vehicles, marine engines, and generators, collectively account for more than half of the air pollutants in the U.S., with vehicles being a primary contributor [80, 81]. Thus, we contextualize the data centers’ public health cost by comparing it to that of on-road emissions of California, which has about 35 million registered vehicles and exhibits the highest public health cost resulting from on-road emissions among all the U.S. states [49, 46]. On-road emissions are categorized as the “Highway Vehicles” sector in COBRA and include both tailpipe exhaust and tire and brake wear. The details of calculating on-road emissions and the corresponding health costs are available in Appendix A.1. We see from Table 1 that in 2023, the total public health cost of U.S. data centers exceeds 1/3 of that of California’s on-road emissions.
4.2 Public Health Impact of U.S. Data Centers in 2030
This section presents our projections of the public health cost of the U.S. data centers in 2030.
We first show in Fig. 2 the health costs of U.S. data centers and compare them with top-3 state on-road emissions in 2030 by using different extrapolation methods. More detailed results are available in Table 2. Due to the tightening air pollutant regulations [82], the health costs of on-road emissions — a primary source of air pollutants in the U.S. — have generally decreased from 2016 to 2030. In contrast, the surging demand for AI data centers in the U.S. has outweighed the power plant emission efficiency improvement, potentially quadrupling the public health cost from 2023 to 2030. Under McKinsey’s projection with a medium growth rate, the scope-2 pollutants of U.S. data centers in 2030 alone could cause, among others, approximately 600,000 asthma symptom cases and 1,300 deaths, exceeding 1/3 of asthma deaths in the U.S. each year [47]. Importantly, the public health costs of U.S. data centers could rival or even exceed those of on-road emissions of the largest U.S. states including California, suggesting a need for urgent attention to the health impact of U.S. data centers beyond on-road emissions.
(a) Linear extrapolation
(b) Exponential extrapolation
(c) 2028-to-2030 unchanged
Figure 2: The health costs of U.S. data centers and top-3 state on-road emissions from 2016 to 2030 based on different extrapolations for 2030 baseline emissions.
Table 2: The public health cost of U.S. data centers in 2030 based on EPRI’s energy demand projection [5]. “” denotes McKinsey’s projection under a medium growth rate [4].
Extrapolation Method Projected Growth Electricity (TWh) Electricity Cost (billion \$) Mortality Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$) Months of Electricity Bill % of CA On-road Health Cost
Unchanged Low 196.3 16.8 490 (360, 610) 8.3 (6.3, 10.3) 49% 55.3 (41.9, 68.8) 0.40 (0.31, 0.50) 43%
Moderate 214.0 18.3 530 (390, 660) 9.0 (6.8, 11.1) 49% 59.9 (45.4, 74.5) 0.44 (0.33, 0.54) 47%
High 296.4 25.4 710 (530, 890) 12.1 (9.2, 15.1) 48% 80.9 (61.3, 100.6) 0.59 (0.45, 0.74) 63%
Higher 403.9 34.6 940 (700, 1180) 16.0 (12.1, 19.8) 46% 106.7 (80.8, 132.6) 0.78 (0.59, 0.97) 84%
Medium 519.3 44.5 1210 (900, 1510) 20.5 (15.5, 25.5) 46% 137.2 (103.9, 170.5) 1.00 (0.76, 1.25) 107%
Linear Low 196.3 16.8 500 (370, 630) 8.5 (6.4, 10.5) 50% 56.5 (42.6, 70.5) 0.41 (0.31, 0.51) 61%
Moderate 214.0 18.3 540 (400, 680) 9.2 (6.9, 11.4) 50% 61.3 (46.2, 76.3) 0.45 (0.34, 0.56) 66%
High 296.4 25.4 730 (540, 920) 12.4 (9.3, 15.4) 49% 82.7 (62.4, 103.1) 0.60 (0.46, 0.75) 90%
Higher 403.9 34.6 960 (710, 1210) 16.3 (12.3, 20.3) 47% 109.0 (82.2, 135.9) 0.80 (0.60, 0.99) 118%
Medium 519.3 44.5 1230 (920, 1550) 21.0 (15.8, 26.1) 47% 140.2 (105.7, 174.7) 1.02 (0.77, 1.28) 152%
Exponential Low 196.3 16.8 510 (380, 640) 8.7 (6.6, 10.8) 52% 58.0 (43.8, 72.3) 0.42 (0.32, 0.53) 53%
Moderate 214.0 18.3 550 (410, 700) 9.4 (7.1, 11.7) 51% 62.9 (47.4, 78.3) 0.46 (0.35, 0.57) 57%
High 296.4 25.4 750 (560, 940) 12.7 (9.6, 15.8) 50% 84.9 (64.1, 105.7) 0.62 (0.47, 0.77) 78%
Higher 403.9 34.6 990 (730, 1240) 16.7 (12.6, 20.9) 48% 111.9 (84.4, 139.4) 0.82 (0.62, 1.02) 102%
Medium 519.3 44.5 1270 (940, 1590) 21.5 (16.2, 26.8) 48% 143.9 (108.6, 179.2) 1.05 (0.79, 1.31) 132%
(a) Per-household health cost map
(b) CDF of county-level per-household health cost
State County Per-household Health Cost (\$) Months of Electricity Bills County-to-nation Income Ratio WV Mason 986.4 (782.3, 1190.5) 7.1 (5.6, 8.6) 0.71 OH Meigs 981.1 (744.5, 1217.7) 7.9 (6.0, 9.8) 0.62 WV Marion 977.0 (783.2, 1170.9) 7.1 (5.7, 8.5) 0.80 WV Marshall 974.0 (759.0, 1189.0) 7.0 (5.5, 8.6) 0.77 OH Gallia 939.6 (710.7, 1168.6) 7.5 (5.7, 9.4) 0.74 WV Taylor 890.1 (722.9, 1057.2) 6.4 (5.2, 7.6) 0.70 PA Fayette 874.9 (694.0, 1055.7) 6.1 (4.8, 7.4) 0.74 WV Brooke 861.6 (660.5, 1062.7) 6.2 (4.8, 7.7) 0.69 PA Greene 837.9 (687.2, 988.5) 5.9 (4.8, 6.9) 0.88 WV Jackson 753.2 (610.5, 895.9) 5.4 (4.4, 6.5) 0.73
(c) Top-10 counties by per-household health cost
County Per-household Health Cost (\$) Months of Electricity Bills County-to-nation Income Ratio Emporia City 510.4 (369.2, 651.6) 3.6 (2.6, 4.6) 0.55 Lancaster 509.5 (405.5, 613.6) 3.6 (2.9, 4.3) 0.83 Staunton City 502.7 (386.3, 619.2) 3.5 (2.7, 4.4) 0.79 Buena Vista City 490.5 (378.8, 602.3) 3.5 (2.7, 4.3) 0.65 Highland 466.0 (367.9, 564.1) 3.3 (2.6, 4.0) 0.76 Buchanan 463.2 (362.3, 564.0) 3.3 (2.6, 4.0) 0.53 Fairfax City 462.8 (344.9, 580.8) 3.3 (2.4, 4.1) 1.71 Dickenson 454.0 (354.2, 553.9) 3.2 (2.5, 3.9) 0.53 Tazewell 441.8 (344.8, 538.8) 3.1 (2.4, 3.8) 0.62 Colonial Heights City 435.5 (340.3, 530.7) 3.1 (2.4, 3.7) 0.96
(d) Top-10 counties in Virginia by per-household health cost
Figure 3: The county-level per-household health cost of U.S. data centers in 2030 based on exponential extrapolation of baseline emissions (McKinsey’s medium-growth forecast). The income data is based on the 2018-2022 American Community Survey 5-year estimates provided by [83].
Next, we show in Fig. 3 the county-level per-household health cost of U.S. data centers in 2030 based on exponential extrapolation under McKinsey’s medium-growth forecast. We see that the health cost is highly disproportionately distributed across different counties and communities, particularly affecting low-income communities. The ratio of the highest county-level per-household health cost to the lowest cost could be more than 200. Crucially, all the top-10 counties in the U.S. and 9 out of top-10 counties in Virginia (which has the largest concentration of data centers in the U.S. [5, 4]) have lower median household incomes than the national median value. Moreover, many of the hardest-hit communities do not have large data centers or directly receive economic benefits from AI data centers such as tax revenues. Yet, compared to the national average of about 1 month of electricity bill per year, the households in these communities could each suffer from health impacts equivalent to up to $\sim$8 months of their electricity bills. The high degree of disparity across different communities in terms of the public health cost suggests that we must examine the local and regional health impacts of AI data centers and improve public health equity to enable truly responsible AI.
(a) Health cost map
(b) CDF of county-level health cost
State County Health Cost (million \$) IL Cook 309.6 (214.4, 404.7) PA Allegheny 266.4 (204.7, 328.1) TX Harris 262.0 (182.9, 341.1) MI Wayne 171.8 (123.0, 220.6) PA Philadelphia 169.4 (118.0, 220.8) OH Hamilton 166.5 (127.3, 205.7) NY Kings 165.2 (105.2, 225.2) OH Franklin 152.0 (112.2, 191.8) OH Cuyahoga 151.1 (110.6, 191.5) NY Queens 138.6 (90.7, 186.5)
(c) Top-10 counties by total health cost
Figure 4: The county-level total health cost of U.S. data centers in 2030 based on exponential extrapolation of baseline emissions (McKinsey’s medium-growth forecast).
We also show the county-level total public health cost in Fig. 4. Compared to the per-household health cost distribution in Fig. 3, the county-level total health cost distribution is more aligned with the population distribution — despite the low per-household health cost, populous counties in California have a high total health cost. Nonetheless, some less populous counties (e.g., Hamilton County, Ohio) near coal and/or natural gas power plants are still significantly impacted and even more so than those (e.g., Loudoun County, Virginia) that have high concentrations of data centers.
Table 3: The public health cost of training an AI model of the Llama-3.1 scale in Meta’s U.S. data centers.
Location Electricity Price (¢/kWh) Electricity Cost (million \$) Health Cost (million \$) % of Electricity Cost Emission (Metric Ton)
PM2.5 (LA-NYC) NOx (LA-NYC) SO2
Huntsville, AL 7.11 2.1 0.70 (0.54, 0.87) 33% 0.61 (13800) 2.80 (2500) 2.72
Stanton Springs, GA 6.88 2.0 0.85 (0.65, 1.04) 41% 0.69 (15500) 3.37 (3000) 3.35
DeKalb, IL 8.20 2.4 1.92 (1.41, 2.42) 79% 1.25 (28100) 7.31 (6600) 7.83
Altoona, IA 6.91 2.1 2.51 (1.84, 3.17) 122% 1.52 (34000) 11.78 (10600) 14.76
Sarpy, NE 7.63 2.3 1.54 (1.16, 1.92) 68% 1.13 (25300) 13.5 (12200) 18.51
Los Lunas, NM 5.75 1.7 0.73 (0.56, 0.90) 43% 0.78 (17500) 8.36 (7500) 9.84
Forest City, NC 7.15 2.1 1.07 (0.85, 1.30) 50% 0.72 (16200) 5.72 (5200) 3.27
New Albany, OH 7.03 2.1 1.61 (1.20, 2.03) 77% 1.13 (25200) 5.15 (4600) 4.44
Prineville, OR 7.52 2.2 0.23 (0.19, 0.28) 10% 0.59 (13300) 4.67 (4200) 2.40
Gallatin, TN 6.23 1.9 0.32 (0.24, 0.40) 17% 0.41 (9200) 1.21 (1100) 0.93
Fort Worth, TX 6.60 2.0 0.51 (0.38, 0.65) 26% 0.47 (10500) 3.02 (2700) 3.81
Eagle Mountain, UT 6.99 2.1 0.24 (0.19, 0.29) 12% 0.60 (13300) 4.82 (4300) 2.52
Henrico, VA 8.92 2.7 1.61 (1.20, 2.03) 61% 1.13 (25200) 5.15 (4600) 4.44
4.3 Public Health Impact of Generative AI Training
We now study the health impact of training a generative AI model. Specifically, we consider the training of an LLM and assume that the electricity consumption is the same as training Llama-3.1 recently released by Meta [84]. While we use Meta’s Llama-3.1 training electricity consumption and U.S. data center locations as an example, our results should be interpreted as the estimated public health impact of training a general LLM with a comparable scale of Llama-3.1.
We show the results in Table 3. It can be seen that the total health cost can even exceed 120% of the electricity cost and vary widely depending on the training data center locations. For example, the total health cost is only \$0.23 million in Oregon, whereas the cost will increase dramatically to \$2.5 million in Iowa due to various factors, such as the wind direction and the pollutant emission rate for electricity generation [70]. Additionally, depending on the locations, training an AI model of the Llama-3.1 scale can produce an amount of air pollutants equivalent to more than 10,000 LA-NYC round trips by car.
The results highlight that the public health impact of AI model training is highly location-dependent. Combined with the spatial flexibility of model training, they suggest that AI model developers should take into account potential health impacts when choosing data center locations for training.
5 Our Recommendations
We provide our recommendations to address the increasing public health impact of AI.
Recommendation 1: Standardization of Reporting Protocols
Despite their immediate and tangible impacts on public health, criteria air pollutants have been entirely overlooked in AI model cards and sustainability reports published by technology companies [45, 44, 10]. The absence of such critical information adds substantial challenges to accurately identifying specific AI data centers as a key root cause of public health burdens and could potentially pose hidden risks to public health. To enhance transparency and lay the foundation for truly responsible AI, we recommend standardization of reporting protocols for criteria air pollutants and the public health impacts across different regions. Concretely, criteria air pollutants can be categorized into three different scopes (Section 2.2), and reported following the greenhouse gas protocol widely adopted by technology companies [45, 10, 85].
Just as addressing scope-2 and scope-3 carbon emissions is important for mitigating climate change, it is equally crucial to address scope-2 and scope-3 criteria air pollutants to promote public health throughout the power generation and hardware manufacturing processes in support of AI. For instance, power plants are dispatched based on real-time energy demand to ensure grid stability. As a result, only focusing on regulating scope-2 air pollutants at the power plant level fails to address the root cause — electricity consumption — and overlooks the potential of demand-side solutions. In contrast, recognizing scope-2 air pollutants and their associated public health impacts enables novel opportunities for health-informed AI, which, as detailed below, taps into demand-side flexibilities to holistically reduce AI’s adverse public health impacts.
Recommendation 2: Health-informed AI
Data centers, including those operated by major technology companies [10, 45], predominantly rely on grid electricity due to the practical challenges of installing on-site low-pollutant and low-carbon energy sources at scale. However, the spatial-temporal variations of scope-2 health costs (Fig. 5) open up new opportunities to reduce the public health impact by exploiting the high scheduling flexibilities of AI training and inference workloads. For example, as further supported by EPRI’s recent initiative on maximizing data center flexibility for demand response [11], AI training can be scheduled in more than one data center, while multiple AI models with different resource-performance tradeoffs are often available to serve AI inference requests. To date, the existing data centers have mostly exploited such scheduling flexibilities for reducing electricity costs [86], carbon emissions [15], water consumption [87], and/or environmental inequity [88]. Nonetheless, the public health impact of AI significantly differs from these environmental costs or metrics.
Concretely, despite sharing some common sources (e.g., fossil fuels) with carbon emissions, the public health impact resulting from the dispersion of criteria air pollutants is highly dependent on the emission source location and only exhibits a weak correlation with carbon emissions. For example, the same quantity of carbon emissions generally results in the same climate change impacts regardless of the emission source; in contrast, criteria air pollutants have substantially greater public health impacts if emitted in densely populated regions compared to sparsely populated or unpopulated regions, emphasizing the importance of considering spatial variability.
(a) Normalized IQR (110)
(b) Normalized STD (90)
(c) Average
Figure 5: Analysis of marginal scope-2 carbon emission rates and public health costs over 114 U.S. regions between October 1, 2023 and September 30, 2024 [71]. (a) In 110 out of the 114 U.S. regions (96%), the normalized IQR of marginal health cost is higher than that of marginal carbon intensity. (b) In 90 out of the 114 U.S. regions (79%), the normalized standard deviation of marginal health cost is higher than that of marginal carbon intensity. (c) The Pearson correlation between the per-region yearly average marginal health cost and carbon intensity is 0.292.
To further confirm this point, we analyze the scope-2 marginal carbon intensity and public health cost for each unit of electricity generation across all the 114 U.S. regions between October 1, 2023, and September 30, 2024, provided by [71]. [^6] The health cost signal provided by [71] only considers mortality from PM2.5, while COBRA includes a variety of health outcomes including asthma, lung cancer, and mortality from ozone, among others [22]. The time granularity for data collection is 5 minutes. We show in Fig. 5a the region-wise normalized interquartile ranges (IQR divided by the yearly average) for both public health costs and carbon emissions. The normalized IQR measures the spread of the time-varying health and carbon signals. Specifically, in 110 out of the 114 U.S. regions (96%), the normalized IQR of health cost is higher than that of the carbon intensity for each unit of electricity consumption. Moreover, the normalized IQR for carbon emissions is less than 0.2 in most of the regions. This implies that health costs exhibit a greater temporal variation than carbon emissions in 110 out of the 114 U.S. regions. Likewise, in Fig. 5b, the greater temporal variation of health costs is also supported by its greater normalized standard deviation (STD divided by the yearly average) in 90 out of the 114 U.S. regions (79%). Next, we show in Fig. 5c the weak spatial correlation (Pearson correlation coefficient: 0.292) between the yearly average health cost and carbon intensity across the 114 regions. Furthermore, the normalized IQR of the health cost spatial distribution is 3.62x that of carbon emission spatial distribution (1.05 vs. 0.29), while the health-to-carbon ratio in terms of the spatial distribution’s normalized STD is 3.37 (0.64 vs. 0.19). In other words, the health cost has a greater spatial spread than the carbon emission.
These findings highlight that leveraging spatial-temporal variations in a health-aware manner could significantly reduce AI’s public health costs while still maintaining climate benefits. As a result, we advocate for a new research direction — health-informed AI. Specifically, decisions regarding the siting of AI data centers and the runtime scheduling of AI tasks should explicitly address their public health impacts. By judiciously accounting for and exploiting the spatial-temporal diversity of health costs, AI data centers can be optimized to minimize adverse public health impacts while supporting sustainability goals.
Additionally, as the public health awareness serves as an effective implicit incentive (e.g., as demonstrated in the context of residential energy conservation [89]), AI data center operators can also leverage this approach by informing end users about the public health impacts of their AI usage. This may help extract additional user-side demand flexibilities as part of the recent efforts to maximize the overall data center load flexibility [11].
Recommendation 3: Attention to All
Counties and communities located near AI data centers or supplying electricity to them often experience most significant health burdens. Nonetheless, these health impacts can extend far beyond the immediate vicinity, affecting communities hundreds of miles away [34, 33]. For example, the health impact of backup generators in northern Virginia can affect several surrounding states (Fig. 1a) and even reach as far as Florida.
While the health impact on communities where data centers operate is increasingly recognized, there has been very little, if any, attention paid to other impacted communities that bear substantial public health burdens. This disconnect leaves those communities to shoulder the public health cost of AI silently without receiving adequate support. To fulfill their commitment to social responsibility, we recommend technology companies holistically evaluate the cross-state public health burden imposed by their operations on all impacted communities, when deciding where they build data centers, where they get electricity for their data centers, and where they install renewables.
Additionally, to quantify the health effects on impacted communities with greater accuracy for potential regulatory actions, we recommend further interdisciplinary research such as cross-state air quality dispersion, health economics, and health-informed computing.
Recommendation 4: Promoting Public Health Equity
The public health impact of AI is highly unevenly distributed across different counties and communities in the U.S., often disproportionately affecting low-income communities and potentially exacerbating socioeconomic inequities [90, 37]. For example, as shown in Table 3c and 3d, all the top-10 counties in the U.S. and 9 out of top-10 counties in Virginia have lower median household incomes than the national median value. The ratio of the highest county-level per-household health cost to the lowest cost could be more than 200. Critically, minimizing the total health cost without considering equity can even reinforce existing inequities, similar to the way environmental inequities have been amplified [88]. Therefore, it is imperative to address the substantial health impact disparities across communities and ensure that AI fosters public health equity rather than exacerbating inequities.
6 Conclusion
In this paper, we uncover and quantify the overlooked public health impact of AI. We present a general methodology to model air pollutant emissions across AI’s lifecycle, from chip manufacturing to data center operation. Our findings demonstrate that under McKinsey’s projection with a medium-growth scenario, the U.S. data centers in 2030 could contribute to nearly 1,300 deaths annually, resulting in a public health burden of more than \$20 billion which could even exceed that of on-road emissions of California. Importantly, these public health costs are unevenly distributed and disproportionately impact low-income communities, where the per-household health burden could be equivalent to nearly 8 months of electricity bills and 200x compared to other less-impacted counties. We recommend adopting a standard reporting protocol for criteria air pollutants and public health costs, paying attention to impacted communities, and implementing health-informed AI to mitigate these effects while promoting public health equity.
Our study provides novel insights for the public, policymakers, and technology companies, enabling a more comprehensive cost-benefit analysis of AI’s impacts on society. We also call for further research to fully address the public health implications when developing powerful and responsible AI in the future. It is crucial to prioritize public health and ensure that the growth of AI does not exacerbate health burdens or negate the potential benefits AI can bring in improving public health.
Acknowledgement
The authors would like to thank Susan Wierman for providing comments on an initial draft of this paper.
Appendix
Appendix A Implementation Details
We describe the evaluation methodology used for our empirical analysis. We use the latest COBRA (Desktop v5.1, as of October 2024) provided by the U.S. EPA [46] to study the public health impact of U.S. data centers in both 2023 and 2030. While COBRA uses a reduced-complexity air quality dispersion model based on a source-receptor matrix for rapid evaluation, its accuracy has been validated and the same or similar model has been commonly adopted in the literature for large-area air quality and health impact analysis [76, 78, 74, 79]. We consider county-level air pollutant dispersion throughout the contiguous U.S., which is the area currently supported by COBRA [46]. Note that cities considered county-equivalents for census purposes are also referred to as “counties” in COBRA. Throughout the paper, we use “county” without further specification.
All the monetary values are presented in the 2023 U.S. dollars unless otherwise stated. We set the discount rate as 2% in COBRA as recommended by the EPA based on the U.S. Office of Management and Budget Circular No. A-4 guidance [46]. When presenting a single value or a ratio (e.g., health-to-electricity cost ratio) if applicable, we use the midrange of the low and high estimates provided by COBRA.
A.1 Estimation of 2030 Baseline Emissions
For estimates in 2030, COBRA provides data for county-level population, health incidence, and valuation, but the baseline emissions are missing [46]. Thus, we estimate the 2030 baseline emission by extrapolating the data for 2016, 2023, and 2028 provided by COBRA. Specifically, we consider three different extrapolation methods as follows.
$\bullet$ Linear: For each pollutant type (e.g., PM2.5, SO2, and NOx) at each source, we apply a linear model $y=a\cdot t+b$, where $t$ is the year, to fit the 2016, 2023, and 2028 values and use the linear model to estimate the 2030 value. We also calculate the coefficient of determination, or $R^{2}$ score for each linear model. If $R^{2}$ is less than 0.5, we set the predicted 2030 value equal to the 2028 value. In addition, if the value is missing for a pollutant type at a source for any of the three years (2016, 2023, and 2028), we directly use the 2028 value as the 2030 value.
$\bullet$ Exponential: The exponential extrapolation method is similar to the linear method. When the model $y=a\cdot(1+r)^{t}$ shows an exponentially decreasing trend from 2016 to 2028 (i.e., $r<0$), we apply the model to estimate the 2030 value. Nonetheless, when the trend from 2016 to 2028 is increasing (i.e., $r>0$), we roll back to a linear model for conservative estimates to avoid over-estimates resulting from an exponential model.
$\bullet$ Unchanged: We directly apply the 2028 baseline emission data to 2030.
We show in Table 4 and Table 5 the estimated total baseline emissions of air pollutants for electricity generation and on-road traffic in 2030 using different extrapolation methods. We also show the baseline emissions for 2016, 2023, and 2028 as provided by COBRA [46]. By reducing a state’s on-road emissions to zero in COBRA, we obtain the corresponding public health cost in that state.
Table 4: U.S. electricity generation baseline emissions from 2016 to 2030
Year Electricity Generation Emission (Metric Ton)
NOx SO2 PM2.5 VOC
2016 1100575.41 1369417.44 111604.62 30250.76
2023 711746.94 717409.25 110878.22 34311.54
2028 695495.34 733437.11 110279.40 34446.71
2030 (Linear) 682541.75 726267.77 119326.10 36903.77
2030 (Exponential) 707846.63 751245.61 120870.45 37488.27
Table 5: U.S. and California on-road baseline emissions from 2016 to 2030
Year U.S. On-road Emission (Metric Ton) California On-road Emission (Metric Ton)
NOx SO2 PM2.5 VOC NOx SO2 PM2.5 VOC
2016 3293579.05 25001.53 106828.36 1680342.17 202427.66 1438.07 10197.26 89087.60
2023 1588423.83 11325.07 65742.16 996965.92 98095.76 1280.27 8144.83 54141.57
2028 1130369.84 10616.37 53455.43 758508.40 86573.30 1154.27 8276.27 44586.45
2030 (Linear) 594848.10 6402.84 40555.54 545983.16 52560.36 1109.71 7583.01 33840.13
2030 (Exponential) 925971.64 9009.44 47978.65 653737.61 68881.37 1122.56 7910.51 38536.35
On-road emissions are categorized as the “Highway Vehicles” sector in COBRA and include both tailpipe exhaust and tire and brake wear. Thus, following the EPA and U.S. Department of Transportation classification [22, 91], PM2.5 resulting from road dust is not counted as emissions of highway vehicles in our study. If the PM2.5 from paved road dust (categorized as “Miscellaneous $\to$ Other Fugitive Dust $\to$ Paved Roads” in COBRA) is considered, California is still projected to have the highest state-wide public health cost of on-road vehicles among all the U.S. states in 2030. For example, by assuming exponential extrapolation and including \$7.6 billion attributed to paved road dust PM2.5, California is projected to have a total health cost of \$23.9 billion. Nonetheless, even by including paved road dust PM2.5, our finding still indicates that the public health cost of U.S. data centers (e.g., \$21.5 billion based on McKinsey’s projection) could be comparable to that of California’s on-road emissions in 2030.
A.2 Evaluation of AI’s Public Health Impact (Scope 2)
Due to the limited data available for scope-1 and scope-3 impacts, we mainly focus on the scope-2 health impacts unless otherwise specified. Thus, the locations of emission sources depend on the power plants supplying electricity to data centers. To evaluate the public health impacts of U.S. data centers, we consider both average attribution and marginal attribution methods for 2023. Nonetheless, since it is difficult, if not impossible, to obtain the marginal emission rate without knowing the actual dispatch decisions for the future, we only use the average attribution method for 2030. The two attribution methods are described as follows.
$\bullet$ Average attribution: We first calculate the total data center electricity consumption $e_{DC}$ and the overall electricity consumption (including non-data center loads) $e_{Total}$ within each electricity region. The U.S. electricity grid is divided into 14 regions following the AVoided Emissions and geneRation Tool (AVERT, the latest version v4.3 as of October 2024) provided by the EPA [70]. We use the state-level electricity consumption data for 2023 and 2030 provided by EPRI [5], and distribute state-level electricity consumption to relevant electricity regions following the state-to-region electricity apportionment used by AVERT. Note that the actual state-to-region electricity apportionment in 2030 may vary from the assumption in AVERT. Thus, we also consider an alternative apportionment to further evaluate the public health impact of U.S. data centers. Specifically, we consider a state-level electricity apportionment scenario in which each state is viewed as an electricity region. The evaluation results are shown in Appendix C and further reinforce our key finding that the health impact of U.S. data centers could rival that of on-road emissions in some of the largest U.S. states such as California.
We calculate the percentage $x\%=\frac{e_{DC}}{e_{Total}}$ of the data center electricity consumption with respect to the overall electricity consumption for each electricity region. The relationship between the health impact and emission reduction in COBRA is approximately linear. Thus, we apply a reduction by $x\%$ to the baseline emissions of all the power plants within the respective electricity region in COBRA and estimate the corresponding county-level health impacts, including health outcomes and costs.
When assessing the health impact of generative AI training, we follow the same approach, except for changing the total data center electricity consumption to the AI model training electricity consumption.
Assuming a medium growth rate, McKinsey projects that the U.S. data center electricity demand (excluding cryptocurrency) will reach 606 TWh, or 11.7% of the U.S. national electricity demand, in 2030 [4]. When using McKinsey’s projection, we only use its projected percentage of 11.7%. That is, we consider the EPRI’s projection of non-data center loads and scale up the EPRI’s projection of data center electricity demand to match the percentage of 11.7%. As a result, the 2030 U.S. data center electricity demand is 519 TWh, instead of 606 TWh, in our study under McKinsey’s projection. Nonetheless, as we apply a reduction by $x\%$ to the baseline emissions in COBRA, what matters most is the percentage, rather than the absolute electricity consumption by data centers.
$\bullet$ Marginal attribution: We only consider marginal attribution for 2023. Specifically, we use the state-level data center electricity consumption [5] and run AVERT to calculate the resulting county-level marginal air pollutant reduction [70]. AVERT allows a maximum of 15% electricity reduction within an electricity region during each hour. For regions where the data center electricity demand exceeds the 15% reduction threshold for certain hours in 2023, we cap the reduction at 15%, which results in a conservative estimate (i.e., the actual health impact of data centers is slightly higher). The county-level emission reduction data provided by AVERT is then applied to COBRA to estimate the county-level health outcomes and costs.
Electricity price. When estimating the electricity cost for data centers in 2023 and 2030, we use the state-level average price for industrial users in [92]. The projected U.S. nominal electricity price for industrial users remains nearly the same from 2023 to 2030 (24.96 \$/MMBtu in 2023 vs. 23.04 \$/MMBTu in 2030) in the baseline case per the EIA’s Energy Outlook 2023 [39]. Thus, our estimated health-to-electricity cost ratio will be even higher if we further adjust inflation. Similarly, to estimate the household electricity bills, we use the state-level average price for residential users and county-level average household electricity consumption in [92].
Location-based emission. There are two types of scope-2 carbon emissions associated with electricity consumption: location-based and market-based [10]. Specifically, location-based carbon emissions refer to the physical carbon emissions attributed to an electricity consumer connected to the power grid, while market-based carbon emissions are net emissions after applying reductions due to contractual arrangements and other credits (e.g., renewable energy credits). In this paper, similar to location-based carbon emissions commonly studied in the literature [8], we focus on criteria air pollutants for AI data centers without considering market-based pollution reduction mechanisms.
While data centers, including large technology companies, often use various credits to reduce their market-based carbon emissions [10], it is likely less effective to apply this practice to mitigate the public health impact. The reason is that, unlike carbon emissions that have a similar effect on climate change regardless of the emission source locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source. For example, the public health impact of using grid power from a populated region may not be effectively mitigated by the renewable energy credits generated elsewhere.
A.3 Public Health Impact of Backup Generators in Virginia
Virginia has issued a total of 174 air quality permits for data center backup generators as of December 1, 2024 [56]. More than half of the data center sites are within Loudoun County. We collect a dataset of the air quality permits: permits issued before January 1, 2023, from [57], and permits issued between January 1, 2023 and December 1, 2024, from [56]. The total permitted site-level annual emission limits are approximately 13,000 tons of NOx, 1,400 tons of VOCs, 50 tons of SO2, and 600 tons of PM2.5, all in U.S. short tons. By assuming that the actual emissions are 10% of the permitted level, the data centers in Virginia could already cause approximately 14,000 asthma symptom cases and 13-19 deaths each year, among other health implications, resulting in a total annual public health burden of \$220-300 million, including \$190-260 million incurred in Virginia, West Virginia, Maryland, Pennsylvania, New York, New Jersey, Delaware, and Washington D.C., as estimated by COBRA under the “Fuel Combustion: Industrial” sector.
A.4 Public Health Impact of a Semiconductor Facility
We consider a semiconductor manufacturing facility located in Ocotillo, a neighborhood in Chandler, Arizona [93]. By averaging the rolling 12-month air pollutant emission levels listed in the recent air quality monitoring report (as of October, 2024) [18], we obtain the annual emissions as follows: 150.4 tons of NOx, 82.7 tons of VOCs, 1.1 tons of SO2, and 28.9 tons of PM2.5. By applying these on-site emissions to COBRA under the “Other Industrial Processes” sector, we obtain a total public health cost of \$14-21 million. Additionally, the total annual energy consumption by the facility is 2074.88 million kWh as of Q2, 2024 [93]. Assuming 84.2% of the energy comes from the electricity based on the company’s global average [94], we obtain the facility’s annual electricity consumption as 1746.63 million kWh. By using the average attribution method, we further obtain an estimated health cost of \$12-17 million associated with the electricity consumption. Thus, the total health cost of the facility is \$26-39 million.
By relocating the facility from Chandler, Arizona, to a planned site in Licking County, Ohio, and assuming the same emission level and electricity consumption, we can obtain the total health cost of \$94-156 million, including \$23-36 million attributed to direct on-site emissions and \$70-120 million attributed to electricity consumption.
A.5 Energy Consumption for Training a Generative AI Model
We consider Llama-3.1 as an example generative AI model. According to the model card [44], the training process of Llama-3.1 (including 8B, 70B, and 405B) utilizes a cumulative of 39.3 million GPU hours of computation on H100-80GB hardware, and each GPU has a thermal design power of 700 watts. Considering Meta’s 2023 PUE of 1.08 [45] and excluding the non-GPU overhead for servers, we estimate the total training energy consumption as approximately 30 GWh.
A.6 Average Emission for Each LA-NYC Round Trip by Car
We use the 2023 national average emission rate for light-duty vehicles (gasoline) provided by the U.S. Department of Transportation [91]. The emission rate accounts for tailpipe exhaust, tire wear and brake wear. Specifically, the average PM2.5 emission rate is 0.008 grams/mile (including 0.004 grams/mile for exhaust, 0.003 grams/mile for brake wear, and 0.001 grams/mile for tire wear), and the average NOx emission rate is 0.199 grams/mile for exhaust. We see that half of PM2.5 for light-duty vehicles comes from brake and tire wear (0.004 gram/miles), which are also produced by other types of vehicles including electric vehicles. The distance for a round-trip between Los Angeles, California, and New York City, New York, is about 5,580 miles. Thus, the average auto emissions for each LA-NYC round trip are estimated as 44.64 grams of PM2.5 and 1110.42 grams of NOx.
Appendix B Public Health Impact of U.S. Data Centers in 2023
We show in Fig. 6 the state-wide data center electricity consumption in 2023 [5]. It can be seen that Virginia, Texas and California have the highest data center electricity consumption in 2023.
Next, we show in Fig. 7 the county-level per-household (scope-2) health cost caused by the U.S. data centers in 2023. We see that the health cost is highly disproportionately distributed across different counties and communities, particularly affecting low-income communities. The ratio of the highest county-level per-household health cost to the lowest cost is more than 100. Crucially, all the top-10 counties in the U.S. have lower median household incomes than the national median value. Moreover, by comparing Fig. 7 and Fig. 6, we see that many of the hardest-hit communities do not have large data centers or directly receive economic benefits from AI data centers such as tax revenues. We also show in Fig. 8 the county-level total health costs of U.S. data centers in 2023.
(a) Electricity consumption map
State Electricity Consumption (TWh) State Electricity Consumption (TWh) State Electricity Consumption (TWh) VA 33.85 OH 2.36 ID 0.15 TX 21.81 SC 2.02 WI 0.15 CA 9.33 WY 1.86 MD 0.10 IL 7.45 KY 1.62 LA 0.08 OR 6.41 CO 1.51 SD 0.07 AZ 6.25 AL 1.49 ME 0.03 IA 6.19 FL 1.38 NH 0.02 GA 6.18 TN 1.33 RI 0.02 WA 5.17 OK 1.23 KS $<$ 0.01 PA 4.59 MA 1.06 AR $<$ 0.01 NY 4.07 MO 0.97 DE $<$ 0.01 NJ 4.04 MN 0.82 DC $<$ 0.01 NE 3.96 MT 0.58 MS $<$ 0.01 ND 3.92 MI 0.53 VT $<$ 0.01 NV 3.42 NM 0.40 WV $<$ 0.01 NC 2.67 CT 0.26 UT 2.56 IN 0.19
(b) Electricity consumption by state (descending order)
Figure 6: State-level electricity consumption of U.S. data centers in 2023 [5].
(a) Per-household health cost map
(b) CDF of county-level per-household health cost
State County Per-household Health Cost  (\$) Months of Electricity Bills County-to-nation Income Ratio WV Marion 306.0 (244.9, 367.1) 2.2 (1.8, 2.6) 0.80 WV Mason 299.4 (235.6, 363.1) 2.2 (1.7, 2.6) 0.71 OH Meigs 294.4 (220.0, 368.8) 2.4 (1.8, 3.0) 0.62 OH Gallia 289.9 (216.5, 363.3) 2.3 (1.7, 2.9) 0.74 WV Marshall 280.6 (215.6, 345.7) 2.0 (1.6, 2.5) 0.77 WV Taylor 266.6 (215.4, 317.7) 1.9 (1.6, 2.3) 0.70 PA Fayette 256.1 (201.9, 310.3) 1.8 (1.4, 2.2) 0.74 PA Greene 245.4 (200.2, 290.5) 1.7 (1.4, 2.0) 0.88 WV Brooke 235.7 (177.9, 293.5) 1.7 (1.3, 2.1) 0.69 WV Jackson 227.1 (183.3, 270.9) 1.6 (1.3, 2.0) 0.73
(c) Top-10 counties by per-household health cost
Figure 7: The county-level per-household health cost of U.S. data centers in 2023.
(a) Health cost map
(b) CDF of county-level health cost
State County Health Cost (million \$) IL Cook 81.9 (55.8, 107.9) PA Allegheny 80.1 (61.1, 99.2) TX Harris 68.3 (46.4, 90.2) OH Hamilton 49.0 (37.1, 60.9) PA Philadelphia 48.1 (33.4, 62.7) MI Wayne 46.8 (33.0, 60.6) OH Cuyahoga 43.1 (31.1, 55.2) NY Kings 41.7 (26.1, 57.3) OH Franklin 41.5 (30.3, 52.7) NY Queens 35.2 (22.5, 47.8)
(c) Top-10 counties by health cost
Figure 8: The county-level health cost of U.S. data centers in 2023.
We show in Fig. 9 the per-household health cost of U.S. data centers in 2023 by considering the marginal attribution method. The health cost using marginal attribution means the public health burden resulting from the additional loads of the U.S. data centers connected to the grid in 2023. In other words, if the U.S. data centers had been powered using off-grid sources (e.g., on-site renewables) in 2023, the per-household public health benefit would be valued at up to \$319 and the total public health benefit would be \$7.6 billion.
Table 6: The public health cost of U.S. data centers in 2030. “” denotes McKinsey’s projection under a medium growth rate (excluding energy consumption for cryptocurrency) [4]. State-level electricity apportionment.
Extrapolation Method Projected Growth Electricity (TWh) Electricity Cost (billion \$) Mortality Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$) Months of Electricity Bill % of CA On-road Health Cost
Unchanged Low 196.3 16.8 370 (270, 460) 6.3 (4.8, 7.8) 37% 42.0 (31.8, 52.1) 0.31 (0.23, 0.38) 33%
Moderate 214.0 18.3 400 (300, 500) 6.8 (5.1, 8.4) 37% 45.4 (34.4, 56.3) 0.33 (0.25, 0.41) 36%
High 296.4 25.4 530 (400, 660) 9.0 (6.9, 11.2) 36% 60.4 (45.8, 75.0) 0.44 (0.33, 0.55) 47%
Higher 403.9 34.6 690 (510, 860) 11.8 (8.9, 14.6) 34% 78.6 (59.6, 97.6) 0.57 (0.44, 0.71) 62%
Medium 519.3 44.5 890 (660, 1110) 15.1 (11.5, 18.8) 34% 101.0 (76.6, 125.4) 0.74 (0.56, 0.92) 79%
Linear Low 196.3 16.8 370 (280, 460) 6.3 (4.8, 7.8) 38% 42.2 (32.0, 52.4) 0.31 (0.23, 0.38) 46%
Moderate 214.0 18.3 400 (300, 500) 6.8 (5.2, 8.5) 37% 45.6 (34.5, 56.6) 0.33 (0.25, 0.41) 49%
High 296.4 25.4 530 (400, 670) 9.1 (6.9, 11.3) 36% 60.7 (46.0, 75.5) 0.44 (0.34, 0.55) 66%
Higher 403.9 34.6 690 (520, 870) 11.8 (9.0, 14.7) 34% 79.0 (59.8, 98.2) 0.58 (0.44, 0.72) 86%
Medium 519.3 44.5 890 (660, 1120) 15.2 (11.5, 18.9) 34% 101.6 (76.9, 126.2) 0.74 (0.56, 0.92) 110%
Exponential Low 196.3 16.8 380 (290, 480) 6.5 (4.9, 8.1) 39% 43.6 (33.0, 54.2) 0.32 (0.24, 0.40) 40%
Moderate 214.0 18.3 410 (310, 520) 7.0 (5.3, 8.8) 38% 47.1 (35.7, 58.5) 0.34 (0.26, 0.43) 43%
High 296.4 25.4 550 (410, 690) 9.4 (7.1, 11.7) 37% 62.7 (47.5, 78.0) 0.46 (0.35, 0.57) 57%
Higher 403.9 34.6 720 (530, 900) 12.2 (9.2, 15.2) 35% 81.6 (61.8, 101.4) 0.60 (0.45, 0.74) 75%
Medium 519.3 44.5 920 (690, 1160) 15.7 (11.9, 19.5) 35% 104.9 (79.4, 130.4) 0.77 (0.58, 0.95) 96%
(a) Per-household health cost map
(b) CDF of county-level per-household health cost
State County Per-household Health Cost (\$) Moths of Electricity Bills County-to-nation Income Ratio WV Mason 319.8 (258.1, 381.5) 2.3 (1.9, 2.8) 0.71 OH Meigs 308.5 (235.2, 381.8) 2.5 (1.9, 3.1) 0.62 OH Gallia 299.6 (229.0, 370.2) 2.4 (1.8, 3.0) 0.74 WV Brooke 285.3 (213.3, 357.2) 2.1 (1.5, 2.6) 0.69 WV Marshall 270.7 (204.2, 337.2) 2.0 (1.5, 2.4) 0.77 PA Fayette 254.3 (195.8, 312.8) 1.8 (1.4, 2.2) 0.74 WV Marion 252.8 (194.8, 310.8) 1.8 (1.4, 2.2) 0.80 WV Jackson 252.2 (206.8, 297.7) 1.8 (1.5, 2.1) 0.73 WV Hancock 252.1 (193.1, 311.1) 1.8 (1.4, 2.2) 0.77 WV Roane 241.4 (196.8, 286.0) 1.7 (1.4, 2.1) 0.55
(c) Top-10 counties by per-household health cost
Figure 9: The county-level per-household health cost of U.S. data centers in 2023. Marginal attribution.
Appendix C Public Health Impact of U.S. Data Centers in 2030 (State-level Electricity Apportionment)
AVERT [70] divides the U.S. electricity grid into 14 regions. Since the actual state-to-region electricity apportionment in 2030 may vary from the assumption in AVERT, we now consider an alternative apportionment to further evaluate the public health impact of U.S. data centers in 2030. Specifically, we hypothesize a state-level electricity apportionment scenario in which each state is viewed as an electricity region (i.e., data centers are powered by in-state electricity). We show the results in Table 6, Fig. 10, and Fig. 11. While the actual values slightly differ from those in Section 4.2, the key message remains the same: the health impact of U.S. data centers could rival that of on-road emissions in some of the largest U.S. states such as California, and disproportionately affect low-income communities. As we consider in-state electricity to power data centers, 9 out of 10 most-effected counties in terms of the per-household public health burden are in Virginia which has the largest concentration of data centers [5].
(a) Linear extrapolation
(b) Exponential extrapolation
(c) 2028-to-2030 unchanged
Figure 10: The health costs of U.S. data centers and top-3 state on-road emissions from 2016 to 2030 based on different extrapolations for 2030 baseline emissions. (State-level electricity apportionment.)
(a) Per-household health cost map
(b) CDF of county-level per-household health cost
State County Per-household Health Cost (\$) Months of Electricity Bills County-to-nation Income Ratio VA Emporia City 882.6 (623.3, 1141.8) 6.2 (4.4, 8.1) 0.55 VA Colonial Heights City 661.4 (516.2, 806.5) 4.7 (3.6, 5.7) 0.96 VA Brunswick 621.2 (441.8, 800.6) 4.4 (3.1, 5.7) 0.70 VA Greensville 607.1 (408.1, 806.1) 4.3 (2.9, 5.7) 0.69 VA Hopewell City 603.8 (466.9, 740.7) 4.3 (3.3, 5.2) 0.67 VA Hanover 593.7 (455.7, 731.6) 4.2 (3.2, 5.2) 1.39 VA Lancaster 580.3 (465.9, 694.7) 4.1 (3.3, 4.9) 0.83 VA Petersburg City 576.0 (429.3, 722.7) 4.1 (3.0, 5.1) 0.62 ND McLean 565.7 (451.1, 680.2) 4.8 (3.8, 5.8) 1.07 VA Sussex 563.1 (397.7, 728.4) 4.0 (2.8, 5.1) 0.79
(c) Top-10 counties by per-household health cost
Figure 11: The county-level per-household health cost of U.S. data centers in 2030 based on exponential extrapolation of baseline emissions (McKinsey’s medium-growth forecast). The income data is based on the 2018-2022 American Community Survey 5-year estimates provided by [83]. (State-level electricity apportionment.)
Appendix D Health-informed AI
We now provide additional results to highlight the importance of health-informed AI.
D.1 Correlation Analysis of Marginal Carbon Intensity and Health Impact for Google’s U.S. Data Center Locations
In addition to the analysis in Section 5, we study the scope-2 marginal carbon intensity and public health cost for each unit of electricity generation across Google’s U.S. data center between October 1, 2023, and September 30, 2024, provided by [71]. The health cost signal provided by [71] only considers mortality from PM2.5, while COBRA includes a variety of health outcomes including asthma, lung cancer, and mortality from Ozone, among others [22]. The time granularity for data collection is 5 minutes.
We present the results Table 7, further confirming that carbon intensities and health impacts are not always aligned and that health impacts vary more significantly than carbon intensities in almost all the locations. This suggests that, by judiciously accounting for and exploiting the spatial-temporal diversity of health costs, AI data centers can be optimized to minimize adverse public health impacts while supporting sustainability goals.
Location Pearson Correlation Normalized IQR Normalized STD
Health Carbon $\frac{\text{Health}}{\text{Carbon}}$ Ratio Health Carbon $\frac{\text{Health}}{\text{Carbon}}$ Ratio
Loudoun County, VA 0.427 0.158 0.065 2.409 0.131 0.059 2.222
Central Ohio, OH 0.479 0.160 0.065 2.441 0.137 0.066 2.064
The Dalles, OR 0.326 0.957 0.099 9.614 0.546 0.103 5.296
Douglas County, GA 0.756 0.507 0.093 5.418 0.293 0.075 3.913
Montgomery County, TN 0.760 0.289 0.067 4.320 0.195 0.046 4.236
Papillion, NE 0.736 0.748 0.840 0.891 0.487 0.553 0.881
Storey County, NV 0.584 0.178 0.057 3.132 0.168 0.042 4.004
Ellis County, TX 0.474 0.196 0.082 2.384 0.232 0.361 0.641
Berkeley County, SC 0.416 0.156 0.054 2.911 0.105 0.044 2.405
Council Bluffs, IA 0.361 0.185 0.111 1.671 0.129 0.311 0.415
Henderson, NV 0.584 0.178 0.057 3.132 0.168 0.042 4.004
Jackson County, AL 0.760 0.289 0.067 4.320 0.195 0.046 4.236
Lenoir, NC 0.240 0.176 0.059 2.982 0.129 0.046 2.800
Mayes County, OK 0.617 0.122 0.049 2.495 0.171 0.222 0.772
Table 7: Correlation analysis of marginal carbon emissions and health impacts for Google’s U.S. data center locations between October 1, 2023, and September 30, 2024 [71]. According to the region classification of WattTime [95], the two data centers in Storey County, NV, and Henderson, NV, belong to the same power grid region, and so do those in Jackson County, AL, and Montgomery County, TN.
D.2 Location-dependent Public Health Impact
We now show the location-dependent public health impacts of two technology companies based on Google’s and Meta’s U.S. data center locations in 2023, excluding their leased colocation data centers whose locations are proprietary. Due to the lack of information about the per-data center electricity consumption, we uniformly distribute Google’s North America electricity consumption over its U.S. data center locations based on Google’s latest sustainability report [10]. Meta discloses its per-location electricity usage [45]. We consider criteria air pollutants without accounting for renewable energy credits these two companies apply to offset their grid electricity consumption (see “Location-based emission” in Appendix A.2). As a consequence, although we consider the U.S. data center locations of Google and Meta, our results should not be interpreted as a quantitative evaluation of these two specific companies’ actual public health impacts. We also emphasize that our goal is to highlight the locational dependency of public health impacts and to motivate the need for health-informed siting of data centers. In our results, we refer to Google and Meta as Company A and Company B, respectively, to avoid potential misunderstandings.
We first see from Table 8 that while the two companies have different public health costs due to their different electricity consumption, their health-to-electricity cost ratios are similar at the national level. Nonetheless, we notice from Fig. 12 that the two companies have significant differences in terms of the per-household health cost distribution and most-affected counties. This is primarily due to the two companies’ different data center locations, and highlights the locational dependency of public health impacts. That is, unlike carbon emissions that have a similar effect on climate change regardless of the emission source locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source. Thus, technology companies should account for public health impacts when deciding where they build data centers, where they get electricity for their data centers, and where they install renewables in order to best mitigate the adverse health effects while promoting equity.
Table 8: The public health costs based on two technology companies’ U.S. data center electricity consumption in 2023.
Company (Attribution) Electricity (TWh) Electricity Cost (billion \$) Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$)
A (Average) 18.5 1.4 0.63 (0.47, 0.78) 45% 4.5 (3.4, 5.5)
A (Marginal) 18.5 1.4 0.97 (0.75, 1.20) 70% 6.9 (5.3, 8.6)
B (Average) 10.6 0.8 0.38 (0.29, 0.48) 51% 2.7 (2.0, 3.4)
B (Marginal) 10.6 0.8 0.53 (0.41, 0.66) 71% 3.8 (2.9, 4.7)
(a) Per-household health cost (Company A)
(b) Per-household health cost (Company B)
(c) CDF of per-household health cost (Company A)
(d) CDF of per-household health cost (Company B)
State County Per-household Health Cost (\$) County-to-nation Income Ratio TX Marion 33.8 (27.2, 40.4) 0.64 VA Mecklenburg 23.7 (19.3, 28.1) 0.68 VA Halifax 23.5 (18.8, 28.1) 0.65 NC Person 23.5 (19.4, 27.6) 0.81 VA Martinsville City 22.5 (18.6, 26.5) 0.52 VA Danville City 21.7 (17.3, 26.1) 0.55 TX Cass 21.2 (16.5, 25.8) 0.72 GA Pickens 21.1 (17.3, 24.9) 0.97 WV Marion 20.5 (16.3, 24.6) 0.80 VA Henry 20.4 (15.9, 24.8) 0.58
(e) Top-10 counties by per-household health cost (Company A)
State County Per-household Health Cost (\$) County-to-nation Income Ratio TX Marion 21.1 (17.0, 25.3) 0.64 TX Cass 13.3 (10.4, 16.3) 0.72 WV Marion 12.4 (9.9, 15.0) 0.80 GA Pickens 12.4 (10.1, 14.6) 0.97 WV Marshall 12.1 (9.2, 14.9) 0.77 WV Mason 12.0 (9.4, 14.6) 0.71 TX Gregg 12.0 (9.6, 14.5) 0.85 TX Harrison 12.0 (9.7, 14.3) 0.84 TX Morris 12.0 (9.3, 14.6) 0.69 OH Gallia 11.9 (8.9, 14.9) 0.74
(f) Top-10 counties by per-household health cost (Company B)
Figure 12: The county-level per-household health cost of two companies in 2023. The income data is based on the 2018-2022 American Community Survey 5-year estimates provided by [83]. Average attribution.

### v2

@@
-The Unpaid Toll: Quantifying the Public Health Impact of AI
+The Unpaid Toll: Quantifying and Addressing the Public Health Impact of Data Centers
 Abstract
 The surging demand for AI has led to a rapid expansion of energy-intensive data centers, impacting the environment through escalating carbon emissions and water consumption.
-While significant attention has been paid to AI’s growing environmental footprint, the public health burden, a hidden toll of AI, has been largely overlooked.
-Specifically, AI’s lifecycle, from chip manufacturing to data center operation, significantly degrades air quality through emissions of criteria air pollutants such as fine particulate matter, substantially impacting public health.
-This paper introduces a methodology to model pollutant emissions across AI’s lifecycle, quantifying the public health impacts.
-Our findings reveal that training an AI model of the Llama-3.1 scale can produce air pollutants equivalent to more than 10,000 round trips by car between Los Angeles and New York City.
-The total public health burden of U.S.
-data centers in 2030 is valued at up to more than \$20 billion per year, double that of U.S.
-coal-based steelmaking and comparable to that of on-road emissions of California.
-Further, the public health costs unevenly impact economically-disadvantaged communities, where the per-household health burden could be 200x more than that in less-impacted communities.
-We recommend adopting a standard reporting protocol for criteria air pollutants and the public health costs of AI, paying attention to all impacted communities, and implementing health-informed AI to mitigate adverse effects while promoting public health equity.
+While significant attention has been paid to data centers’ growing environmental footprint, the public health burden, a hidden toll of data centers, has been largely overlooked.
+Specifically, data centers’ lifecycle, from chip manufacturing to operation, can significantly degrade air quality through emissions of criteria air pollutants such as fine particulate matter, substantially impacting public health.
+This paper introduces a principled methodology to model lifecycle pollutant emissions for data centers and computing tasks, quantifying the public health impacts.
+Our findings reveal that training a large AI model comparable to the Llama-3.1 scale can produce air pollutants equivalent to more than 10,000 round trips by car between Los Angeles and New York City.
+The growing demand for AI is projected to push the total annual public health burden of U.S.
+data centers up to more than \$20 billion in 2028, rivaling that of on-road emissions of California.
+Further, the public health costs are more felt in disadvantaged communities, where the per-household health burden could be 200x more than that in less-impacted communities.
+Finally, we propose a health-informed computing framework that explicitly incorporates public health risk as a key metric for scheduling data center workloads across space and time, which can effectively mitigate adverse health impacts while advancing environmental sustainability.
+More broadly, we also recommend adopting a standard reporting protocol for the public health impacts of data centers and paying attention to all impacted communities.
 1 Introduction
@@
 For example, by integrating multimodal data from various sources, AI can provide effective tools and actionable insights for pandemic preparedness, disease prevention, healthcare optimization, and air quality management [3, 1].
-However, the surging demand for AI — particularly generative AI, as exemplified by the recent popularity of large language models (LLMs) — has driven a rapid increase in computational needs, fueling the unprecedented expansion of energy-intensive AI data centers.
-According to McKinsey projections, under a medium-growth scenario [4], the U.S.
-data centers are anticipated to account for 11.7% of national electricity consumption in 2030, a substantial increase from their current share of less than 4% in 2023.
-The growing electricity demand of AI data centers has not only created significant stress on power grid stability [5, 6], but also increasingly impacts the environment through escalating carbon emissions [7, 8] and water consumption [9].
-These environmental impacts are driven primarily by the “expansion of AI products and services,” as recently acknowledged by Google in its latest sustainability report [10].
+However, the surging demand for AI — particularly generative AI, as exemplified by the recent popularity of large language models (LLMs) — has driven a rapid increase in computational needs, fueling the unprecedented expansion of energy-intensive data centers.
+According to the recent Lawrence Berkeley National Lab report [4], AI training and inference are projected to become the dominant workloads and push the U.S.
+data center electricity consumption to account for 6.7–12.0% of the national total in 2028, up from 4.4% in 2023.
+The growing electricity demand of data centers has not only created significant stress on power grid stability [5, 6], but also increasingly impacts the environment through escalating carbon emissions [7, 8] and water consumption [9].
+These environmental impacts are driven primarily by the “expansion of AI products and services,” as recently acknowledged by technology companies in their sustainability reports [10].
 To mitigate the challenges posed to both power grids and the environment, a range of strategies have been explored, including grid-integrated data centers [6, 11], energy-efficient hardware and software [12, 13, 14], and the adoption of carbon-aware and water-efficient computing practices [15, 9, 16, 17], among others.
-The hidden toll of AI.
-While the environmental footprint of AI has garnered attention, the public health burden, a hidden toll of AI, has been largely overlooked.
-Across its entire lifecycle — from chip manufacturing to data center operation — AI contributes substantially to air quality degradation and public health costs through the emission of various criteria air pollutants.
+The hidden toll of data centers.
+While the environmental footprint of data centers has garnered attention, the public health burden, a hidden toll of data centers, has been largely overlooked.
+Across its entire lifecycle — from chip manufacturing to operation — a data center contributes substantially to air quality degradation and public health costs through the emission of various criteria air pollutants.
 These include fine particulate matter (PM2.5, particles measuring 2.5 micrometers or smaller in diameter that can penetrate deep into lungs and cause serious health effects), sulfur dioxide (SO2), and nitrogen dioxide (NO2).
-Concretely, the AI hardware manufacturing process [18], electricity generation from fossil fuels to power AI data centers, and the maintenance and usage of diesel backup generators to ensure continuous AI data center operation all produce significant amounts of criteria air pollutants.
-Moreover, the distinct spatial-temporal heterogeneities of emission sources suggest that focusing solely on reducing AI’s carbon footprints may not minimize its emissions of criteria air pollutants or the resulting public health impacts (Section 5).
+Concretely, the server manufacturing process [18], electricity generation from fossil fuels to power data centers, and the maintenance and usage of diesel backup generators to ensure continuous data center operation all produce significant amounts of criteria air pollutants.
+Moreover, the distinct spatial-temporal heterogeneities of emission sources suggest that focusing solely on reducing data centers’ carbon footprints may not minimize its emissions of criteria air pollutants or the resulting public health impacts (Section 6.2).
 Exposure to criteria air pollutants is directly and causally linked to various adverse health outcomes, [^2] While we focus on public health, we note that the impacts of criteria air pollutants extend beyond humans and include harms to environmentally sensitive areas, such as some national parks and wilderness areas which, classified as “Class 1 areas” under the Clean Air Act, require special air protection [19].
-including premature mortality, lung cancer, asthma, heart attacks, cardiovascular diseases, strokes, and even cognitive decline, especially for the elderly and vulnerable individuals with pre-existing conditions [20, 21, 22, 23].
+including premature mortality, lung cancer, asthma, heart attacks, cardiovascular diseases, and even cognitive decline, especially for the elderly and vulnerable individuals with pre-existing conditions [20, 21, 22].
 Moreover, even short-term (hours to days) PM2.5 exposure is harmful and deadly, accounting for approximately 1 million premature deaths per year from 2000 to 2019 and representing 2% of total global deaths [23].
+Criteria air pollutants are not confined to the immediate vicinity of their emission sources; they can travel hundreds of miles through a dispersion process (i.e., cross-state air pollution) [24, 25], impacting public health across vast regions.
+Further, PM2.5 is considered “non-threshold,” i.e., there is no absolutely safe exposure level [26].
+Thus, compliance with the national/regional air quality standards does not necessarily ensure the air is healthy.
 Globally, 4.2 million deaths were attributed to ambient (i.e., outdoor) air pollution in 2019 [27].
@@
 Notably, according to the latest Global Burden of Disease report [29], along with high blood pressure and high blood sugar, ambient particulate matter is placed among the leading risk factors for disease burden globally in every socio-demographic group.
-While the U.S.
-has generally better air quality than many other countries, 4 in 10 people in the U.S.
-still live with unhealthy levels of air pollution, according to the “State of the Air 2024” report published by the American Lung Association [28].
-In 2019 (the latest year of data provided by the World Health Organization, or WHO, as of November 2024), an estimate of 93,886 deaths in the U.S.
-were attributed to ambient air pollution [29].
-In fact, even compliance with the U.S.
-Environmental Protection Agency (EPA) air quality standards does not necessarily guarantee healthy air that meets the WHO guidelines.
-Concretely, the EPA’s recently tightened primary standard for PM2.5 sets an annual average limit of 9 $\mu g/m^{3}$, considerably higher than the WHO’s recommended level of 5 $\mu g/m^{3}$ [30, 31].
-In addition, the EPA projects that 53 U.S.
-counties, including 23 in the most populous state of California, would fail to meet the revised national annual PM2.5 standard in 2032 [32].
-Further, criteria air pollutants are not confined to the immediate vicinity of their emission sources; they can travel hundreds of miles through a dispersion process (i.e., cross-state air pollution) [33, 34], impacting public health across vast regions — pollutants from the 2024 Canadian wildfires significantly degraded air quality across much of the U.S.
-and reached as far as Mexico and Europe [35].
-Importantly, along with transportation and industrial activities, electricity generation is a major contributor to ambient air pollution with substantial public health impacts [26, 36, 37].
+Importantly, along with transportation and industrial activities, electricity generation is a key contributor to ambient air pollution with substantial public health impacts [28, 30, 31].
 For example, a recent study [32] shows that, between 1999 and 2020, a total of 460,000 excess deaths were attributed to PM2.5 generated by coal-fired power plants alone in the U.S.
 As highlighted by the U.S.
-EPA [36], despite years of progress, “fossil fuel-based power plants remain a leading source of air, water, and land pollution that affects communities nationwide.” Moreover, according to the U.S.
-Energy Information Administration (EIA) projection [39], the coal consumption by the electricity sector in 2050 will still be about 30% of the 2024 level in the baseline reference case, and the number will exceed 50% in the high zero-carbon technology cost case.
-Indeed, the growing energy demands of AI are already delaying the decommissioning of coal-fired power plants and increasing fossil-fuel plants in the U.S.
-as well as around the world [6, 40, 41].
-The public health outcomes of AI due to its emission of criteria air pollutants lead to various losses, such as hospitalizations, medication usage, emergency room visits, school loss days, and lost workdays.
-Moreover, these losses can be further quantified in economic costs based on epidemiology and economics research for the corresponding health endpoints [42, 22].
-In contrast, the environmental impacts of AI, e.g., carbon emission from fossil fuels and water consumption for data center cooling, often do not cause the same immediate health impacts.
-For instance, while anthropogenic carbon emissions could also pose risks to public health, such impacts are often second- or third-order effects through long-term climate change which can then threaten the human well-being by affecting the food people eat and facilitating the spreading of pests, among others [43].
-Nonetheless, despite their immediate and tangible impacts on public health, the criteria air pollutants of AI have remained under the radar, entirely omitted from today’s AI risk assessments and sustainability reports [44, 45, 10].
-Quantifying the public health costs of AI.
-In this paper, we uncover and quantify the hidden public health impacts of AI.
-We introduce a general methodology to model the emission of criteria air pollutants associated with AI tasks across three distinct scopes: emissions from the maintenance and operation of backup generators (Scope 1), emissions from fossil fuel combustion for electricity generation (Scope 2), and emissions resulting from the manufacturing of server hardware (Scope 3).
-Then, we analyze the dispersion of criteria air pollutants and the resulting public health impacts across different regions.
-Our main results (Section 4) focus on the scope-2 health impacts of U.S.
+EPA [30], despite years of progress, power plants “remain a leading source of air, water, and land pollution that affects communities nationwide.” In Europe, the public health cost of air pollution from power plants is valued at approximately 1% of the gross domestic product (GDP), according to the European Environment Agency’s study in 2024 [33].
+The public health outcomes of data centers due to their emission of criteria air pollutants lead to various losses, such as hospitalizations, medication usage, emergency room visits, school loss days, and lost workdays.
+Nonetheless, despite recent policy efforts [34, 35], the tangible and growing public health impacts of data centers have remained under the radar, almost entirely omitted from today’s risk assessments and sustainability reports [36, 37, 10].
+Quantifying and addressing the public health impacts of data centers.
+In this paper, we uncover and quantify the hidden public health impacts of data centers.
+We introduce a principled methodology to model the emission of criteria air pollutants associated with a computing task and data center across three distinct scopes: emissions from the maintenance and operation of backup generators (Scope 1), emissions from fossil fuel combustion for electricity generation (Scope 2), and emissions resulting from the manufacturing of server hardware (Scope 3).
+Then, we analyze the dispersion of criteria air pollutants and the resulting public health impacts.
+As the U.S.
+hosts nearly half of the world’s data centers [38] and the EPA data excludes other regions [39], our empirical study focuses on the 48 contiguous U.S.
+states plus Washington D.C.
+[^3] If located in countries with higher population densities, more pollutant-intensive electricity mixes, or less stringent air quality standards, the same data centers would likely lead to more premature deaths and other adverse health impacts than in the U.S.
+We recommend further research on the public health impact of non-U.S.
+data centers.
+Our main results (Section 5) focus on the scope-1 and scope-2 health impacts of U.S.
 data centers and, specifically, LLM training.
-[^3] Our study focuses on the 48 contiguous U.S.
-states plus Washington D.C.
-because the EPA data does not include other regions [46].
-If located in countries with higher population densities or less strict air quality standards, the same AI task and data centers would likely contribute to significantly more deaths and other adverse health effects.
-We recommend further research on the public health impact of AI outside the U.S.
-Using the reduced-complexity modeling tool COBRA (CO-Benefits Risk Assessment) provided by the EPA [46], our screening analysis demonstrates that driven by the growing demand for AI, the U.S.
-data centers could contribute to, among others, approximately 600,000 asthma symptom cases and 1,300 premature deaths in 2030, exceeding 1/3 of asthma deaths in the U.S.
+Using the reduced-complexity modeling tool COBRA (CO-Benefits Risk Assessment) provided by the EPA [39], our analysis demonstrates that driven by the growing demand for AI, the U.S.
+data centers could contribute to, among others, approximately 600,000 asthma symptom cases and 1,300 premature deaths in 2028, exceeding 1/3 of asthma deaths in the U.S.
 each year [40].
-The overall public health costs could reach more than \$20 billion, double that of the U.S.
-coal-based steelmaking industry [48], and rival or even top those of on-road emissions of the largest U.S.
+The overall public health costs could reach more than \$20 billion, rival or even top those of on-road emissions of the largest U.S.
 states such as California with $\sim$35 million registered vehicles [41].
 Moreover, depending on the location, training an AI model of the Llama-3.1 scale can produce an amount of air pollutants equivalent to driving a car for more than 10,000 round trips between Los Angeles and New York City (LA-NYC), resulting in a health cost that even exceeds 120% of the training electricity cost.
-Critically, the health costs are unevenly distributed across counties and communities, disproportionately affecting low-income counties (e.g., Meigs County, Ohio) where the per-household health burden could be equivalent to nearly 8 months of electricity bills and more than 200x compared to that in other counties.
-In addition, to highlight the importance of scope-1 and scope-3 health impacts, we consider data center backup generators in Virginia (Scope 1) and semiconductor manufacturing plants in Arizona and Ohio (Scope 3).
-Our analysis shows that, assuming the actual emissions are only 10% of the permitted level, the data center backup generators registered in Virginia (mostly in Loudoun, Prince William, and Fairfax) could already cause 14,000 asthma symptom cases among other health outcomes and a total public health burden of \$220-300 million per year, impacting residents in multiple surrounding states and as far as Florida (Section 2.2.1).
+Importantly, although the public health impact of data centers may be modest at the national level, it is geographically concentrated, with certain regions and communities bearing a disproportionate share.
+In particular, some low-income counties experience significantly greater health costs, with per-household burdens exceeding those in other counties by more than 200-fold.
+Furthermore, to highlight scope-1 health impacts, we examine data center backup generators in Virginia, which hosts one of the largest concentrations of data centers in the world [42].
+Our analysis shows that, assuming the actual emissions are only 10% of the permitted level based on the historical reports and future projections [42, 43, 44], the data center backup generators registered in Virginia (mostly in Loudoun, Prince William, and Fairfax) could already cause 14,000 asthma symptom cases among other health outcomes and a total public health burden of \$220-300 million per year, impacting residents in multiple surrounding states and as far as Florida (Section 3.1).
 If these data centers emit air pollutants at the maximum permitted level, the total public health cost will become 10-fold and reach \$2.2-3.0 billion per year.
-The scope-3 health impact of AI is also substantial.
-For example, just a single semiconductor facility in Arizona can cause an annual public health cost of \$26-39 million, with \$14-21 million attributed to the facility’s on-site emissions of criteria air pollutants (Section 2.2.2).
-Furthermore, relocating the same facility to a planned site in Ohio could almost quadruple the public health cost to \$94-156 million, with \$23-36 million resulting from on-site emissions.
-Finally, we provide recommendations to address the increasing public health impact of AI (Section 5).
-Specifically, we recommend technology companies adopt a standard reporting protocol for criteria air pollutants and public health impacts in their AI model cards and sustainability reports, implement health-informed AI to proactively minimize the adverse health effects of AI data centers, pay attention to all impacted communities, and prioritize reducing the health impact on disadvantaged communities to promote public health equity.
-To summarize, our study sheds light on and quantifies the overlooked public health impact of AI.
-It can inform the public, policymakers, and technology companies in conducting a more comprehensive cost-benefit analysis.
-We also urge further research to comprehensively address the public health implications when developing powerful and truly responsible AI in the future, ensuring that the growth of AI does not exacerbate the health burden or outweigh the potential benefits AI can provide to improve public health.
-2 Background on the Air Quality Impact of AI
-This section presents an overview of AI’s impact on air quality and contribution to criteria air pollutants throughout its lifecycle, beginning with background on criteria air pollutants and U.S.
+To address the growing and uneven distribution of health burdens, we propose health-informed computing that explicitly incorporates the public health risk as a key metric when scheduling data center workloads.
+Specifically, by exploiting the spatial-temporal variations of public health impacts and using spatial load shifting as a case study, we demonstrate that the health-informed approach can significantly reduce the health cost compared to the baseline, while continuing to offer meaningful electricity cost savings and reductions in carbon emissions.
+Finally, we provide broader recommendations to address the increasing public health impact of data centers (Section 7).
+We recommend technology companies adopt a standard reporting protocol for criteria air pollutants and public health impacts in their AI model cards and sustainability reports and pay attention to all impacted communities.
+To summarize, wile AI and data centers offer many societal benefits, our study sheds light on, quantifies, and addresses the often overlooked negative externalities of their resource demand, particularly the public health impact.
+We also urge further research to comprehensively address the public health implications when developing data centers in the future, ensuring that the growth of data centers does not exacerbate the health burden.
+Disclaimer.
+The results presented in this paper are not intended to encourage or discourage the construction of data centers, nor should they be used to support or oppose any specific project, which requires more detailed and context-specific evaluation.
+We do not take a position on decisions related to any specific data centers or the use of AI, but instead provide a quantitative assessment of the potential public health impacts of the data center industry.
+2 Background on Air Pollutants
+This section provides background on criteria air pollutants and U.S.
 air quality policies.
-2.1 Criteria Air Pollutants
+Other countries have similar policies in place to safeguard public health, although their levels of enforcement strictness often differ [33].
 Criteria air pollutants, including PM2.5, SO2 and NO2, are a group of airborne contaminants that are emitted from various sources such as industrial activities and vehicle emissions.
@@
 […]
-While CO2 is broadly classified by the EPA as an air pollutant following the U.S.
-Supreme Court ruling in 2007 [53] and contributes to long-term climate change, it often does not cause the same immediate health impacts as criteria pollutants.
+While the U.S.
+has generally better air quality than many other countries, 4 in 10 people in the U.S.
+still live with unhealthy levels of air pollution, according to the “State of the Air 2024” report published by the American Lung Association [50].
+In 2019 (the latest year of data provided by the World Health Organization, or WHO, as of November 2024), an estimate of 93,886 deaths in the U.S.
+were attributed to ambient air pollution [51].
+In fact, the EPA’s recently tightened standard for PM2.5 sets an annual average limit of 9 $\mu g/m^{3}$, considerably higher than the WHO’s recommended level of 5 $\mu g/m^{3}$ [52, 48].
+Moreover, the EPA projects that 53 U.S.
+counties, including 23 in the already most populous state of California, would fail to meet the revised national annual PM2.5 standard in 2032 [53].
+Although CO2 is broadly classified by the EPA as an air pollutant following the U.S.
+Supreme Court ruling in 2007 [54] and contributes to long-term climate change, it often does not cause the same immediate health impacts as criteria pollutants [55].
 In the U.S., CO2 and other greenhouse gases are subject to different EPA regulations from those for criteria air pollutants.
 Thus, for the sake of presentation in this paper, we use “air pollutants” to solely refer to criteria air pollutants wherever applicable.
-2.2 AI’s Contribution to Air Pollutants
-To understand the impact of AI on air quality, we focus on the three scopes over which AI contributes to criteria air pollutants as well as other toxic materials.
+3 Data Centers’ Contribution to Air Pollutants
+Figure 1: The overview of data centers’ contribution to air pollutants and public health impacts.
+Scope-1 and scope-2 impacts occur during the operation of data centers (“operational”), whereas scope-3 impacts arise from activities across the supply chain (“embodied”).
+This section presents an overview of data centers’ impact on air quality and contribution to criteria air pollutants throughout its lifecycle across three scopes (Fig.
+1).
 The scoping definition in this paper parallels the well-established greenhouse gas protocol [56].
-2.2.1 Scope 1
-The scope-1 public health impact of AI primarily comes from the emission of operating on-site backup generators.
+Specifically, scope-1 and scope-2 air pollutants primarily originate from onsite generators and power plants, collectively referred to as operational emissions, while scope-3 pollutants arise from the supply chain and are referred to as embodied emissions.
+3.1 Scope 1: Onsite Generator
+While the construction phase of a data center directly increases air pollutant emissions, its amortized health impact over a typical 15–20 year lifespan is negligible.
+Therefore, we focus on onsite backup generators as the primary source of scope-1 direct air pollutants.
 Data centers are mission-critical facilities that are designed to operate with high availability and uptime guarantees.
-As a result, to maintain operation during emergencies such as grid outages, AI data centers require highly reliable backup power sources [10, 45].
-Diesel generators are known to emit significant amounts of air pollutants and even hazardous emissions during operation.
+As a result, to maintain operation during emergencies such as grid outages, data centers require highly reliable backup power sources [10, 37].
+Diesel generators are known to emit significant amounts of air pollutants and even hazardous emissions during operation [57].
 For example, they emit 200-600 times more NOx than new or controlled existing natural gas-fired power plants for each unit of electricity produced [58].
-Nonetheless, there is limited experience with cleaner backup alternatives that can provide comparable reliability in real-world settings, as highlighted by the U.S.
+Moreover, capacity redundancy is typically followed for diesel generator installations to ensure high availability [59].
+Nonetheless, there is limited practical experience with cleaner backup alternatives that can provide comparable reliability at scale in real-world settings in the near term, as highlighted by the U.S.
 Department of Energy in its recent recommendations regarding AI data center infrastructures [6].
-Consequently, AI data centers, including those newly built by major technology companies, primarily depend on on-site diesel generators for backup power [6, 10, 45, 56].
-For example, in northern Virginia (mostly in Loudoun, Prince William, and Fairfax), the number of permits for data center diesel generators has increased by about 70% since 2023 compared to the total number of permits issued between 2000 and 2022 [56].
+Consequently, the vast majority of data centers, even including those newly built by major technology companies, depend on onsite diesel generators for backup power [6, 60].
+For example, in Northern Virginia (mostly in Loudoun, Prince William, and Fairfax), the number of permits for data center diesel generators has increased by more than 70% since 2023 compared to the total number of permits issued between 2000 and 2022 [60].
+Importantly, nearly all the diesel generators are Tier 2, which have significantly higher emission rates than Tier 4 units [61, 60].
+The total permitted annual emission limits for these diesel generators are approximately 13,000 tons of NOx, 1,400 tons of VOCs, 50 tons of SO2, and 600 tons of PM2.5, all in U.S.
+short tons.
 (a) Health cost map
@@
 While diesel generators need to comply with air quality regulations and typically do not operate over extended periods of time, regular maintenance and testing are essential to ensure their operational reliability.
-In addition, capacity redundancy is typically followed for diesel generator installations to ensure high availability [58].
-Thus, diesel generators represent a major source of on-site air pollutants for data centers and pose a significant health risk to the public [59].
-For instance, the total permitted annual emission limits for data centers in northern Virginia are approximately 13,000 tons of NOx, 1,400 tons of VOCs, 50 tons of SO2, and 600 tons of PM2.5, all in U.S.
-short tons.
-Assuming that the actual emissions are only 10% of the permitted level, these backup generators could already cause 14,000 asthma symptom cases and 13-19 deaths each year among other health implications, resulting in a total annual public health burden of \$220-300 million throughout the U.S.
+A recent report by the state of Virginia [42] found that the actual air pollutant emissions from backup generators at Virginia’s data centers reached approximately 7% of the total permitted amounts in 2023, primarily for maintenance.
+Likewise, the actual emissions took up 3% to 12% of the permitted levels for some data centers in Quincy, Washington [43].
+Moreover, the U.S.
+EPA recently issued a clarification that would allow data centers to run backup generators for up to 50 hours a year (or roughly 10% of the permits that typically allow 500 hours per year) to participate in demand response—a program designed to reduce grid demand during peak hours, which is increasingly activated as surging data center demand strains grid capacity [44].
+This trend may necessitate extended reliance on backup generators [6].
+What further adds to the public health impact is that many data center generators in a region may operate simultaneously for demand response during grid capacity shortages, potentially resulting in a short-term spike in PM2.5 and NOx emissions that can be particularly harmful [6, 23, 48].
+For example, from June 23 to 25, 2025, some data centers in Loudoun County, Virginia, were instructed to run their on-site diesel generators for demand response, releasing large amounts of air pollutants and “black smoke” [61].
+The high emission rate from onsite generators, combined with extended operation for maintenance and demand response beyond grid outages, could pose serious health risks, especially in regions with a concentration of large data centers.
+To illustrate this point, we consider the data centers’ onsite generators in Virginia.
+Assuming that the actual emissions are 10% of the permitted level as a reference case that reflects both the historical reports and future demand response projections [42, 43, 44],  [^4] If the actual percentage is xx%, our value will be approximately scaled by $\frac{x}{10}$ [22].
+the backup generators could already cause 14,000 asthma symptom cases and 13-19 deaths each year among other health implications, resulting in a total annual public health burden of \$220-300 million throughout the U.S.
 This includes \$190-260 million in Virginia, West Virginia, Maryland, Pennsylvania, Delaware, New Jersey, New York, and Washington D.C.
 We show the county-level health cost and the top-10 counties in Figure 2, while deferring the details of calculations to Appendix A.1.
-Moreover, due to power grid capacity constraints in many U.S.
-states, AI data centers are increasingly pressured to vary their loads subject to the grid’s operating conditions, i.e., grid-integrated data centers [60, 6].
-This trend may necessitate extended reliance on backup generators, e.g., possibly 15 days per year [6].
-Such prolonged usage of diesel generators could substantially elevate AI’s scope-1 air pollution, creating even higher public health costs.
-Concretely, if the backup generators in northern Virginia emit air pollutants at the maximum permitted level, the total public health cost could reach \$2.2-3.0 billion per year.
-What further adds to the public health threat is that many data center generators in a region may operate simultaneously for demand response during grid capacity shortages, potentially resulting in a short-term spike in PM2.5 and NOx emissions that can be particularly harmful [6, 24, 31].
-2.2.2 Scope 2
-While technology companies have started implementing various initiatives — such as purchasing renewable energy credits and nuclear power from small modular reactors [5, 10, 61] — to lower their (market-based) carbon emissions, the vast majority of U.S.
-data centers remain directly powered by local power grids with a substantial portion of fossil fuel-based energy sources [10].
-Thus, just as AI is accountable for scope-2 carbon emissions, it also contributes to scope-2 air pollution through its electricity usage.
-The combustion of fossil fuels for electricity production is a major emitter of criteria air pollutants, releasing large amounts of PM2.5, SO2, NOx, VOCs, and others.
+If the diesel generators in Northern Virginia emit air pollutants at the maximum permitted level, the emission of NOx could even exceed half of the annual total emissions by all sources in the region [42], resulting in a total public health cost of \$2.2-3.0 billion per year.
+3.2 Scope 2: Power Plants
+Just as data centers are accountable for scope-2 carbon emissions, they also contribute to scope-2 air pollution through their electricity usage.
+Along with transportation and manufacturing, the combustion of fossil fuels for electricity production is a leading anthropogenic source of criteria air pollutants, releasing large amounts of PM2.5, SO2, NOx, VOCs, and others [30].
 [^5] Wet cooling towers, including those used by data centers [9, 10] and carbon-free nuclear power plants, rely on water evaporation for heat rejection and produce PM2.5 due to spray drift droplets [63, 64].
 Nonetheless, because of limited data available, we exclude the cooling tower PM2.5 emission from our analysis unless other specified.
-Critically, the growing energy demands of AI are already delaying the decommissioning of coal-fired power plants and increasing fossil-fuel plants in the U.S.
-and other countries [6, 40].
+More alarmingly, the growing energy demands of AI data centers are already delaying the decommissioning of coal-fired power plants and driving the expansion of fossil-fuel-based plants in the U.S.
+and globally [6, 65, 66].
 For example, in addition to keeping 2,099 MW coal generation capacity until 2039 (more than 80% of the 2024 level), Virginia Electric and Power Company plans to install 5,934 MW gas-fired plants to meet the growing energy demand driven by AI data centers [66].
-At the national level, per the EIA’s projection, [39], the 2050 natural gas consumption for U.S.
-electricity generation will be about 80% of the 2024 level in the baseline reference case, and even exceed the 2024 level by 20% if the zero-carbon technology cost is high; for coal consumption by the electricity sector in 2050, the numbers will also be considerably high, about 30% and over 50% of the 2024 level in the baseline reference case and in the high zero-carbon technology cost case, respectively.
-These projections were published by the EIA at the very beginning of the generative AI boom in early 2023.
-More recently, it has been reported that AI data centers could even be primarily powered by coal power plants in some countries [40].
-As a result, AI’s scope-2 air pollution is expected to remain at a high level for a substantially long time into the future.
+Figure 3: Public health costs of electricity generation and on-road emissions in the contiguous U.S.
+in 2023 and 2028 [39].
+The error bars represent high and low estimates returned by COBRA using two different exposure-response functions.
+Based on the emission data projected by the U.S.
+EPA’s COBRA modeling tool [39], we show in Fig.
+3 that the electric power sector’s total public health cost in the contiguous U.S.
+is on track to rise, rivaling that of on-road vehicle emissions by all the registered vehicles (including tailpipe exhausts and brakes) in 2028.
+Looking forward, the U.S.
+Energy Information Administration (EIA) projects that coal consumption by the electricity sector in 2050 will still be approximately 30% of the 2024 level if power plants continue operating under rules existing prior to April 2024 [67].
+At the global scale, the reliance on coal and other fossil fuels for electricity production has seen little change over the past four decades, highlighting the enduring challenges to fully adopt clean energy for powering data centers [68].
+As a result, data centers’ scope-2 air pollution is expected to remain at a high level for a substantially long time into the future.
+Although technology companies have started implementing various initiatives—such as purchasing renewable energy credits and nuclear power from small modular reactors [5, 10, 69]—to lower their (market-based) carbon emissions, the vast majority of U.S.
+data centers remain physically and directly powered by local power grids with a substantial portion of fossil fuel-based energy sources [10].
+While the increasingly stringent transmission line constraint is also driving the co-location of data centers with power plants to speed up the construction process, the onsite plants are often gas-fired [70, 71], raising health concerns [31].
 We also note that the practice of using various credits to offset scope-2 carbon emissions [10] may not be effective for mitigating the scope-2 public health impact.
 The reason is that the public health impact of using grid electricity is highly location-dependent, e.g., the impact in a populated region may not be mitigated by renewable energy generated elsewhere.
-2.2.3 Scope 3
-The surging demand for AI necessitates large quantities of computational hardware, including graphics processing units (GPUs), thus intensifying the supply chain requirements [64].
+3.3 Scope 3: Supply Chain
+The surging demand for AI data centers necessitates large quantities of computational hardware, including graphics processing units (GPUs), thus intensifying the supply chain requirements [72].
 However, semiconductor manufacturing generates various criteria air pollutants, wastewater, toxic materials, and hazardous air emissions [18].
 Moreover, the energy-intensive nature of semiconductor production further contributes to pollutants from power plants.
-Combined with other pollution sources such as transportation and electronic waste recycling [65], the supply chain activities form a large portion of AI’s scope-3 impact on public health.
+Combined with other pollution sources such as transportation and electronic waste recycling [73], the supply chain activities form a large portion of data centers’ scope-3 impact on public health.
 Although semiconductor manufacturing facilities are subject to air quality regulations [74], they still pose significant risks, affecting populations across large regions.
-Maricopa County, AZ, has been an EPA-designated non-attainment area for several years due to its failures to meet federal air quality standards [67].
-The establishment of multiple semiconductor facilities in such areas could further exacerbate air quality issues.
-In 2023–2024, the estimated annual public health impact of a single semiconductor facility was \$26-39 million, with \$14-21 million attributed to direct on-site emissions of air pollutants from the facility, based on COBRA estimates [18, 46].
-Moreover, relocating the facility to a planned site in Licking County, Ohio, could nearly quadruple public health costs to \$94-156 million, with \$23-36 million resulting from direct on-site emissions.
-This increase is partly due to Ohio’s weather conditions and higher reliance on coal-based power [68].
-The details of calculations are available in Appendix A.4.
-Importantly, the global demand for AI chips in 2030 is projected to be tens of times of the overall production capacity of this single facility [69], further magnifying the overall scope-3 public health impact of AI.
-It is also worth noting that additional pollutants, including hazardous air pollutants like hydrogen fluoride, may further elevate public health costs but are not included in this analysis.
-3 Quantifying the Public Health Impact of AI
-To quantify the public health impact of AI, we present a general methodology that quantifies AI’s criteria air pollutants at the emission source, models its dispersed air pollutants at different receptors (i.e., destination regions), and finally obtains the public health impact and cost at each receptor.
-For an AI task (e.g., AI model training), we consider $M$ types of criteria air pollutants, $N$ receptor regions of interest (e.g., all the U.S.
+Importantly, the global demand for AI chips in 2030 is projected to be tens of times of the overall production capacity of this single facility [75], further magnifying the overall scope-3 public health impact of data centers.
+It is also worth noting that additional pollutants, including hazardous air pollutants like hydrogen fluoride, may further elevate public health costs but are not included in this paper.
+However, public data on scope-3 criteria air pollutant emissions from semiconductor manufacturers is limited.
+Thus, we focus on scope-1 and scope-2 emissions, excluding scope-3 impacts from the main analysis.
+A detailed assessment of the health impacts associated with a specific U.S.
+semiconductor manufacturing facility is provided in Appendix A.3.
+4 Quantifying Task-Specific Public Health Impact
+To quantify the public health impact of a specific computing task, we present a principled end-to-end methodology illustrated in Fig.
+1.
+Specifically, the process includes: (1) Quantifying the task’s criteria air pollutants at the emission source; (2) Modeling the dispersed air pollutants at different receptors (i.e., destination regions); (3) Calculating the public health impact and assigning a corresponding monetary value at each receptor.
+For a computing task under consideration (e.g., AI model training), we consider $M$ types of criteria air pollutants, $N$ receptor regions of interest (e.g., all the U.S.
 counties), $H$ types of public health impacts (e.g., mortality, asthma symptoms, school loss days, etc.).
@@
 4.1 Criteria Air Pollutants at the Source
-We first model AI’s criteria air pollutants at the source across the three different scopes in Section 2.2.
-3.1.1 Scope 1
-On-site backup diesel generators are sized based on the data center power capacity and routinely tested to ensure a high availability of the entire data center.
+We first model a computing task’s criteria air pollutants at the source across the three different scopes in Section 3.
+4.1.1 Scope 1.
+Onsite diesel generators are sized based on the data center power capacity, routinely tested to ensure a high availability of the entire data center, and used for demand response.
 Thus, the overall scope-1 air pollutants should be attributed to each computing task based on its power allocation and duration.
 Suppose that the overall scope-1 emission by an AI data center under consideration is $\bar{p}^{s}=(\bar{p}_{1}^{s},\cdots,\bar{p}_{M}^{s})$, for $M$ types of air pollutants, over a timespan of $\overline{T}$ (e.g., one year).
-Considering an AI task that is allocated a fraction of $x\in(0,1]$ of the overall data center power capacity and lasts for a duration of $T$, we express the scope-1 air pollutants attributed to the AI task as
+Considering a task that is allocated a fraction of $x\in(0,1]$ of the overall data center power capacity and lasts for a duration of $T$, we express the scope-1 air pollutants attributed to the task as
 $p^{s}=\frac{x\cdot T}{\overline{T}}\cdot\bar{p}^{s},$ (1)
 which attributes the overall emission $\bar{p}^{s}$ to the task in proportion to its allocated power and duration.
-3.1.2 Scope 2
-AI’s scope-2 air pollutants come from its usage of electricity generated from fossil fuels.
-Suppose that the power grid serving the AI data center has an emission rate of $\gamma=(\gamma_{1},\cdots,\gamma_{M})$ for $M$ types of air pollutants to produce each unit of electricity.
+4.1.2 Scope 2.
+A computing task’s scope-2 air pollutants come from its usage of electricity generated from fossil fuels.
+Suppose that the power grid serving the data center has an emission rate of $\gamma=(\gamma_{1},\cdots,\gamma_{M})$ for $M$ types of air pollutants to produce each unit of electricity.
 In practice, the power grid consists of multiple interconnected power plants to supply electricity to many customers over a wide area (e.g., a balancing area [76]).
@@
 The average emission represents a proportional share of the overall air pollutant emission by an electricity consumer, while the marginal emission is useful for quantifying the additionality of air pollutants due to a consumer’s electricity usage.
-Suppose that the electricity consumption by the AI task is $e$, including the data center overhead captured by the power usage effectiveness. Then, we can write the scope-2 air pollutants as
+Suppose that the electricity consumption by the computing task is $e$, including the data center overhead captured by the power usage effectiveness.
+Then, we write the scope-2 air pollutants as
 $p^{s}=e\cdot\gamma,$ (2)
-which is either based on either average attribution or marginal attribution.
-While the marginal emission is typically associated with a single marginal power plant, the average emission is spread across all the interconnected power plants within a wide area such as a power balancing area [71, 70].
-Thus, when considering the average attribution method, we split the energy consumption $e$ over all the power plants in proportion to their contributions to the grid’s supply and calculate the corresponding per-plant emission.
-In other words, each involved power plant is an individual pollution source, and the air pollutant emission at the $k$-th power plant is $p_{k}^{s}=e\cdot\frac{b_{k}}{\sum_{k}b_{k}}\cdot\gamma_{k}$, where $b_{k}$ is the electricity generation of the $k$-th power plant.
-Since both the average and marginal air pollutant emission rates vary over time and locations to meet the supply-demand balance, we can also refine the calculation of scope-2 air pollutants in (2) by considering the summation of air pollutants over multiple time slots over the AI task’s duration.
-3.1.3 Scope 3
-Following the attribution method for scope-3 carbon emission and water consumption [13, 9], we attribute the AI hardware’s air pollutants during the manufacturing process to a specific task based on the task duration.
-Specifically, let the AI hardware’s expected lifespan be $\overline{T}_{0}$ and the AI task lasts a duration of $T$.
-Considering that the $M$ types of air pollutants for manufacturing the AI hardware are $\bar{p}_{0}^{s}=(\bar{p}_{0,1}^{s},\cdots,\bar{p}_{0,M}^{s})$ and excluding other miscellaneous pollutants (e.g., transportation), we obtain AI’s scope-3 air pollutants as
+which is either based on either average or marginal accounting.
+To evaluate the public health impacts of U.S.
+data centers, we consider the average attribution method unless otherwise noted, which is also the standard methodology of carbon emission accounting used by technology companies in their sustainability reports [10, 78, 37].
+We can also refine the calculation of scope-2 air pollutants in (2) by considering the summation of air pollutants over multiple time slots over the task’s duration.
+Location-based emission.
+There are two types of scope-2 carbon emission accounting associated with electricity consumption: location-based and market-based [10].
+Specifically, location-based carbon emissions refer to the physical carbon emissions attributed to an electricity consumer connected to the power grid, while market-based carbon emissions are net emissions after applying reductions due to contractual arrangements and other credits (e.g., renewable energy credits).
+As noted by a recent study on carbon accounting [79], location-based accounting is considered essential, whereas market-based accounting is valuable.
+Moreover, market-based accounting relies on market instruments whose detailed information is often not publicly disclosed.
+Thus, in this paper, we follow the literature [8] and focus on location-based accounting for scope-2 criteria air pollutants without considering market-based pollution reduction mechanisms.
+We also note that market-based emission reduction is likely less effective to mitigate the public health impact.
+The reason is that, unlike carbon emissions that have a similar effect on climate change regardless of the emission locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source.
+For example, the public health impact of using pollutant-intensive electricity generated from a populated region may not be effectively mitigated by the clean energy credits generated elsewhere.
+4.1.3 Scope 3.
+While our empirical analysis focuses on scope-1 and scope-2 health impacts, we present the scope-3 pollutant attribution method to provide a more complete view.
+Specifically, following the attribution method for scope-3 carbon emission and water consumption [13, 9], we attribute the computing hardware’s air pollutants during the manufacturing process to a specific task based on the task duration.
+Specifically, let the hardware’s expected lifespan be $\overline{T}_{0}$ and the task lasts a duration of $T$.
+Considering that the $M$ types of air pollutants for manufacturing the hardware are $\bar{p}_{0}^{s}=(\bar{p}_{0,1}^{s},\cdots,\bar{p}_{0,M}^{s})$ and excluding other miscellaneous pollutants (e.g., transportation), we obtain the task’s scope-3 air pollutants as
 $p^{s}=\frac{T}{\overline{T}_{0}}\cdot\bar{p}_{0}^{s}.$ (3)
-As an AI server cluster includes multiple hardware components (e.g., GPU and CPU) manufactured in different locations, we apply (3) to estimate the scope-3 air pollutants for each component manufactured in a different location.
+As a server cluster includes multiple hardware components (e.g., GPU and CPU) manufactured in different locations, we apply (3) to estimate the scope-3 air pollutants for each component manufactured in a different location.
 4.2 Air Quality Dispersion Modeling
@@
 Furthermore, the dispersion of air pollutants is influenced by meteorological conditions, such as wind speed and direction.
-The movement of air pollutants can be modeled using mathematical equations to simulate the atmospheric processes governing the dispersion, known as dispersion modeling.
+In practice, dispersion modeling tools are used to track the movement of air pollutants.
+These tools employ complex mathematical equations to simulate the atmospheric processes governing the dispersion.
 By incorporating emission data and meteorological inputs, dispersion modeling can predict pollutant concentrations at selected receptor locations [80].
-We consider a general dispersion model $(p^{r}_{1},\cdots,p^{r}_{N})=D_{\theta}(p^{s})$, which yields the amount of $M$ types of air pollutants $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ at the receptor region $i=1,\cdots,N$.
-The parameter $\theta$ captures the geographical conditions, emission source characteristics (e.g., height), and meteorological data if applicable [73].
+We consider a general dispersion modeling tool $(p^{r}_{1},\cdots,p^{r}_{N})=D_{\theta}(p^{s})$, which yields the amount of $M$ types of air pollutants $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ at the receptor region $i=1,\cdots,N$.
+The parameter $\theta$ captures the geographical conditions, emission source characteristics (e.g., height), and meteorological data [81].
 We apply the dispersion model to each scope of air pollutants (Section 4.1) to estimate the corresponding pollutant concentrations at receptor regions.
-In practice, many dispersion modeling tools are available, including AERMOD, CTDMPLUS, PCAPS and InMAP with a reduced complexity [72, 22, 74, 75].
+Several dispersion modeling tools are available, including AERMOD, PCAPS and InMAP with a reduced complexity [80, 22, 82, 83].
 For example, PCAPS (Pattern Constructed Air Pollution Surfaces), an advanced reduced-complexity model that provides representations of both primarily emitted PM2.5 and secondarily formed PM2.5 and ozone, is used in COBRA as a quick assessment of otherwise lengthy iterations and simulations of various pollution scenarios in terms of the annual average PM2.5 and seasonal average maximum daily average 8-hour ozone [83, 22].
@@
 By assessing pollutant levels $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ and population size at each receptor region $i$, we can estimate the incidences of health outcomes $h_{i}=(h_{i,1},\cdots,h_{i,H})$ and the corresponding public health cost $c_{i}=(c_{i,1},\cdots,c_{i,H})$.
-The relations between $p_{i}^{r}$ and $h_{i}$ and between $h_{i}$ and $c_{i}$ can be established based on epidemiology research [22].
+The relations between $p_{i}^{r}$ and $h_{i}$ and between $h_{i}$ and $c_{i}$ is captured by an exposure-response function and can be established based on epidemiology research [22].
 For example, the premature mortality rate can be modeled as a log-linear function in terms of the PM2.5 level [84].
@@
 Therefore, it is a quantitative scalar measure of the public health impact resulting from a particular pollutant-producing activity.
-3.4 Implementation
-We now briefly describe the specific implementation we use to study the public health impact of U.S. data centers and AI training. The details are available in Appendix A.
-Due to the limited data available for scope-1 and scope-3 impacts, we mainly focus on the scope-2 health impacts from electricity consumption.
-To account for future uncertainties, we use the U.S.
-data center electricity consumption data provided by EPRI [5] and McKinsey [4] under various growth-rate scenarios, excluding cryptocurrency servers.
-Unless otherwise specified, we consider the average attribution method by default, i.e., attribute the overall health impact within an electricity region to data centers in proportion to their electricity consumption.
+4.4 End-to-End Modeling
+Following the end-to-end process shown in Fig.
+1, we now briefly describe our modeling methodology to study the public health impact of U.S.
+data centers and AI training.
+The details are available in Appendix A.
+To quantify data centers’ scope-1 and scope-2 air pollutant emissions, we use air quality permit data for onsite generators [60] and electricity consumption data, including both historical records and 2028 projections, from the Lawrence Berkeley National Laboratory report [4].
 To model the air pollutant dispersion and quantify health impacts, we use the latest COBRA (Desktop v5.1, as of October 2024) provided by the U.S.
@@
 While each health impact model used by COBRA considers 95% confidence intervals, the high-end and low-end estimates provided by COBRA are based on different models instead of the 95% confidence interval of a single model [22].
-COBRA provides data for county-level population, health incidence, and valuation projections in 2030, but the baseline emissions are missing [46].
-Thus, to account for model uncertainties, we estimate the 2030 baseline emission by extrapolating the COBRA data for 2016, 2023, and 2028 using three extrapolation methods (Linear, Exponential, and Unchanged) as detailed in Appendix A.1.
-We only consider the contiguous U.S.
-and simply refer to it as the U.S.
-For consistency with COBRA, cities considered county-equivalents for census purposes are also referred to as “counties” in our paper.
-All our monetary values are for one year (or one AI task if applicable) and in 2023 U.S.
-dollars.
 5 Results
@@
 data centers in aggregate and by training a large generative AI model at specific locations.
-Our results demonstrate that in 2030, the scope-2 pollutants of U.S.
+We focus on the contiguous U.S., which hosts nearly half of the world’s data centers [38], and simply refer to it as the U.S.
+For consistency with COBRA, cities considered county-equivalents for census purposes are also referred to as “counties” in our paper.
+All our monetary values are for one year (or one computing task if applicable) and in 2023 U.S.
+dollars as used by COBRA.
+Our results demonstrate that in 2028, the total scope-1 and scope-2 pollutants of U.S.
 data centers alone could cause, among others, approximately 600,000 asthma symptom cases and 1,300 premature deaths, exceeding 1/3 of asthma deaths in the U.S.
 each year [40].
-The overall public health costs of U.S.
-data centers could rival or even exceed those of on-road emissions of the largest U.S.
-states such as California.
+The overall public health costs could reach more than \$20 billion, rival or even top those of on-road emissions of the largest U.S.
+states such as California with $\sim$35 million registered vehicles [41].
+Importantly, the health costs are unevenly distributed across counties and communities, particularly affecting low-income counties that could experience approximately 200x per-household health costs than others.
 Moreover, depending on the locations, training an AI model of the Llama-3.1 scale can produce an amount of air pollutants equivalent to driving a passenger car for more than 10,000 LA-NYC round trips, resulting in a health cost that even exceeds 120% of the training electricity cost.
-Importantly, the health costs are disproportionately distributed across counties and communities, particularly affecting low-income counties that could experience more than 200x per-household health costs than others.
-4.1 Public Health Impact of U.S. Data Centers in 2023
-We first show in Table 1 the public health cost of U.S.
-data centers in 2023 as a reference.
+5.1 Public Health Impact of U.S. Data Centers
+We now present our public health impact analysis for U.S.
+data centers in aggregate, beginning with the historical analysis from 2019 to 2023 and highlighting the uneven distribution of health impacts, followed by a 2028 projection.
+The overall trend is shown in Fig.
+4, which demonstrates a significant increase in the public health impact of U.S.
+data centers from 2023 to 2028.
+Specifically, the surging demand for AI data centers in the U.S.
+has outweighed the power plant emission efficiency improvement, potentially tripling the public health cost from 2023 to 2028.
+We present the details in Table 1 and see that scope-2 health cost dominates the scope-1 cost.
 [^6] We use the “mid (low, high)” format to represent the midrange, low and high estimates offered by COBRA.
 When presenting a single value or a ratio (e.g., health-to-electricity cost ratio), we use the midrange by default.
+This suggests that while using alternative fuels for onsite generators can help improve local air quality around data centers, greater health benefits can be achieved by powering data centers with pollutant-free electricity sources.
+Mobile sources, including vehicles, marine engines, and generators, collectively account for more than half of the air pollutants in the U.S., with vehicles being a primary contributor [89, 90].
+Thus, to contextualize the public health costs of data centers, we compare them to on-road emissions in the three largest U.S.
+states (California, Texas, and Florida).
+In particular, California, which has about 35 million registered vehicles, exhibits the highest public health cost from on-road emissions among all U.S.
+states [41, 39].
+In the COBRA model, on-road emissions are categorized under the “Highway Vehicles” sector and include both tailpipe exhaust and tire and brake wear.
+The details of calculating on-road emissions and the corresponding health costs are provided in Appendix A.
+As shown in Fig.
+4, in 2023, the total public health cost of U.S.
+data centers is 42% of that from California’s on-road emissions.
+Due to the tightening air pollutant regulations [91], the health costs of on-road emissions have generally decreased from 2019 to 2028.
+However, with rapid growth, the public health impact of U.S.
+data centers is projected to rival or even surpass that of California’s on-road emissions by 2028, underscoring the need for greater attention to the growing public health impacts of the data center industry.
+Figure 4: The public health costs of U.S.
+data centers and top-3 state on-road emissions from 2019 to 2023 and the 2028 projection based on the Lawrence Berkeley National Lab’s report [4].
+The cost for U.S.
+data centers includes scope-1 and scope-2 impacts.
+The “High” and “Low” represent the high and low growth rates considered in [4].
+Table 1: The public health cost of U.S. data centers from 2019 to 2023 and projection in 2028
+Year Electricity (TWh) Electricity Cost (billion \$) Scope Mortality Health Cost (billion \$) Per-Household Health Cost (\$) % of CA On-road Health Cost
+2019 90.28 7.73 Scope-1 7 (6, 8) 0.11 (0.10, 0.13) 0.84 (0.72, 0.97) 1%
+Scope-2 189 (138, 240) 2.92 (2.18, 3.66) 21.51 (16.05, 26.97) 14%
+Total 196 (144, 248) 3.03 (2.27, 3.79) 22.36 (16.77, 27.94) 15%
+2020 105.56 9.04 Scope-1 9 (8, 11) 0.15 (0.13, 0.17) 1.11 (0.94, 1.27) 1%
+Scope-2 233 (171, 296) 3.60 (2.68, 4.51) 26.30 (19.62, 32.98) 18%
+Total 243 (179, 307) 3.75 (2.81, 4.69) 27.41 (20.56, 34.25) 19%
+2021 127.78 10.94 Scope-1 13 (11, 15) 0.20 (0.17, 0.24) 1.48 (1.26, 1.70) 1%
+Scope-2 280 (205, 355) 4.31 (3.22, 5.41) 31.24 (23.31, 39.17) 23%
+Total 293 (216, 370) 4.52 (3.39, 5.64) 32.72 (24.57, 40.88) 24%
+2022 151.39 12.97 Scope-1 21 (17, 24) 0.33 (0.28, 0.38) 2.40 (2.04, 2.76) 2%
+Scope-2 330 (242, 418) 5.08 (3.79, 6.37) 36.44 (27.18, 45.70) 29%
+Total 351 (259, 443) 5.41 (4.07, 6.75) 38.84 (29.22, 48.46) 31%
+2023 176.39 15.11 Scope-1 32 (26, 37) 0.51 (0.43, 0.59) 3.65 (3.08, 4.21) 3%
+Scope-2 401 (294, 508) 6.16 (4.59, 7.73) 43.83 (32.69, 54.97) 39%
+Total 433 (320, 546) 6.67 (5.03, 8.32) 47.48 (35.77, 59.19) 42%
+2019 651.40 55.79 Scope-1 82 (68, 95) 1.32 (1.12, 1.52) 9.49 (8.05, 10.92) 1%
+to Scope-2 1434 (1050, 1818) 22.07 (16.46, 27.67) 159.32 (118.85, 199.80) 24%
+2023 Total 1516 (1118, 1913) 23.38 (17.58, 29.19) 168.81 (126.90, 210.72) 25%
+2028 (Low) 325.00 27.84 Scope-1 54 (46, 63) 0.90 (0.77, 1.03) 6.11 (5.22, 7.00) 5%
+Scope-2 650 (483, 818) 10.78 (8.14, 13.41) 73.29 (55.37, 91.21) 61%
+Total 705 (529, 880) 11.67 (8.91, 14.44) 79.40 (60.59, 98.21) 67%
+2028 (High) 580.00 49.68 Scope-1 97 (82, 112) 1.61 (1.37, 1.84) 10.94 (9.35, 12.53) 9%
+Scope-2 1165 (865, 1464) 19.29 (14.58, 24.01) 131.23 (99.14, 163.31) 110%
+Total 1262 (947, 1576) 20.90 (15.95, 25.85) 142.16 (108.49, 175.84) 119%
+5.1.1 Historical analysis: 2019-2023.
+Table 1 shows the public health cost of U.S.
+data centers from 2019 to 2023 as a reference.
 Even at the beginning of the generative AI boom, the U.S.
-data centers have already resulted in a total public health cost of about \$5.6 billion, or \$39.7 per household, in 2023.
-This is equivalent to 43% of the data centers’ total electricity cost.
-By considering marginal attribution, the U.S.
-data centers’ public health cost increases to about \$7.6 billion in 2023, due to the heavy reliance on fossil fuels by many marginal generators [70].
-This suggests that, by powering the U.S.
-data centers using alternative energy sources (e.g., geothermal) off the main grid, the U.S.
-could have seen a public health benefit of \$7.6 billion in 2023.
-Additional results can be found in Appendix B, including county-wide total and per-household health costs that demonstrate the uneven distribution of health impacts across different communities.
-Table 1: The public health cost of U.S. data centers in 2023.
-Attribution Method Electricity (TWh) Electricity Cost (billion \$) Mortality Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$) Months of Electricity Bill % of CA On-road Health Cost
-Average 152.1 13.0 360 (270, 460) 5.6 (4.2, 7.0) 43% 39.7 (29.6, 49.8) 0.29 (0.22, 0.36) 35%
-Marginal 152.1 13.0 490 (360, 620) 7.6 (5.7, 9.4) 58% 53.8 (40.5, 67.2) 0.39 (0.30, 0.49) 47%
-Mobile sources, including vehicles, marine engines, and generators, collectively account for more than half of the air pollutants in the U.S., with vehicles being a primary contributor [80, 81].
-Thus, we contextualize the data centers’ public health cost by comparing it to that of on-road emissions of California, which has about 35 million registered vehicles and exhibits the highest public health cost resulting from on-road emissions among all the U.S.
-states [49, 46].
-On-road emissions are categorized as the “Highway Vehicles” sector in COBRA and include both tailpipe exhaust and tire and brake wear.
-The details of calculating on-road emissions and the corresponding health costs are available in Appendix A.1.
-We see from Table 1 that in 2023, the total public health cost of U.S.
-data centers exceeds 1/3 of that of California’s on-road emissions.
-4.2 Public Health Impact of U.S. Data Centers in 2030
-This section presents our projections of the public health cost of the U.S. data centers in 2030.
-We first show in Fig.
-2 the health costs of U.S.
-data centers and compare them with top-3 state on-road emissions in 2030 by using different extrapolation methods.
-More detailed results are available in Table 2.
-Due to the tightening air pollutant regulations [82], the health costs of on-road emissions — a primary source of air pollutants in the U.S.
-— have generally decreased from 2016 to 2030.
-In contrast, the surging demand for AI data centers in the U.S.
-has outweighed the power plant emission efficiency improvement, potentially quadrupling the public health cost from 2023 to 2030.
-Under McKinsey’s projection with a medium growth rate, the scope-2 pollutants of U.S.
-data centers in 2030 alone could cause, among others, approximately 600,000 asthma symptom cases and 1,300 deaths, exceeding 1/3 of asthma deaths in the U.S.
-each year [47].
-Importantly, the public health costs of U.S.
-data centers could rival or even exceed those of on-road emissions of the largest U.S.
-states including California, suggesting a need for urgent attention to the health impact of U.S.
-data centers beyond on-road emissions.
-(a) Linear extrapolation
-(b) Exponential extrapolation
-(c) 2028-to-2030 unchanged
-Figure 2: The health costs of U.S. data centers and top-3 state on-road emissions from 2016 to 2030 based on different extrapolations for 2030 baseline emissions.
-Table 2: The public health cost of U.S. data centers in 2030 based on EPRI’s energy demand projection [5]. “” denotes McKinsey’s projection under a medium growth rate [4].
-Extrapolation Method Projected Growth Electricity (TWh) Electricity Cost (billion \$) Mortality Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$) Months of Electricity Bill % of CA On-road Health Cost
-Unchanged Low 196.3 16.8 490 (360, 610) 8.3 (6.3, 10.3) 49% 55.3 (41.9, 68.8) 0.40 (0.31, 0.50) 43%
-Moderate 214.0 18.3 530 (390, 660) 9.0 (6.8, 11.1) 49% 59.9 (45.4, 74.5) 0.44 (0.33, 0.54) 47%
-High 296.4 25.4 710 (530, 890) 12.1 (9.2, 15.1) 48% 80.9 (61.3, 100.6) 0.59 (0.45, 0.74) 63%
-Higher 403.9 34.6 940 (700, 1180) 16.0 (12.1, 19.8) 46% 106.7 (80.8, 132.6) 0.78 (0.59, 0.97) 84%
-Medium 519.3 44.5 1210 (900, 1510) 20.5 (15.5, 25.5) 46% 137.2 (103.9, 170.5) 1.00 (0.76, 1.25) 107%
-Linear Low 196.3 16.8 500 (370, 630) 8.5 (6.4, 10.5) 50% 56.5 (42.6, 70.5) 0.41 (0.31, 0.51) 61%
-Moderate 214.0 18.3 540 (400, 680) 9.2 (6.9, 11.4) 50% 61.3 (46.2, 76.3) 0.45 (0.34, 0.56) 66%
-High 296.4 25.4 730 (540, 920) 12.4 (9.3, 15.4) 49% 82.7 (62.4, 103.1) 0.60 (0.46, 0.75) 90%
-Higher 403.9 34.6 960 (710, 1210) 16.3 (12.3, 20.3) 47% 109.0 (82.2, 135.9) 0.80 (0.60, 0.99) 118%
-Medium 519.3 44.5 1230 (920, 1550) 21.0 (15.8, 26.1) 47% 140.2 (105.7, 174.7) 1.02 (0.77, 1.28) 152%
-Exponential Low 196.3 16.8 510 (380, 640) 8.7 (6.6, 10.8) 52% 58.0 (43.8, 72.3) 0.42 (0.32, 0.53) 53%
-Moderate 214.0 18.3 550 (410, 700) 9.4 (7.1, 11.7) 51% 62.9 (47.4, 78.3) 0.46 (0.35, 0.57) 57%
-High 296.4 25.4 750 (560, 940) 12.7 (9.6, 15.8) 50% 84.9 (64.1, 105.7) 0.62 (0.47, 0.77) 78%
-Higher 403.9 34.6 990 (730, 1240) 16.7 (12.6, 20.9) 48% 111.9 (84.4, 139.4) 0.82 (0.62, 1.02) 102%
-Medium 519.3 44.5 1270 (940, 1590) 21.5 (16.2, 26.8) 48% 143.9 (108.6, 179.2) 1.05 (0.79, 1.31) 132%
-(a) Per-household health cost map
-(b) CDF of county-level per-household health cost
-State County Per-household Health Cost (\$) Months of Electricity Bills County-to-nation Income Ratio WV Mason 986.4 (782.3, 1190.5) 7.1 (5.6, 8.6) 0.71 OH Meigs 981.1 (744.5, 1217.7) 7.9 (6.0, 9.8) 0.62 WV Marion 977.0 (783.2, 1170.9) 7.1 (5.7, 8.5) 0.80 WV Marshall 974.0 (759.0, 1189.0) 7.0 (5.5, 8.6) 0.77 OH Gallia 939.6 (710.7, 1168.6) 7.5 (5.7, 9.4) 0.74 WV Taylor 890.1 (722.9, 1057.2) 6.4 (5.2, 7.6) 0.70 PA Fayette 874.9 (694.0, 1055.7) 6.1 (4.8, 7.4) 0.74 WV Brooke 861.6 (660.5, 1062.7) 6.2 (4.8, 7.7) 0.69 PA Greene 837.9 (687.2, 988.5) 5.9 (4.8, 6.9) 0.88 WV Jackson 753.2 (610.5, 895.9) 5.4 (4.4, 6.5) 0.73
-(c) Top-10 counties by per-household health cost
-County Per-household Health Cost (\$) Months of Electricity Bills County-to-nation Income Ratio Emporia City 510.4 (369.2, 651.6) 3.6 (2.6, 4.6) 0.55 Lancaster 509.5 (405.5, 613.6) 3.6 (2.9, 4.3) 0.83 Staunton City 502.7 (386.3, 619.2) 3.5 (2.7, 4.4) 0.79 Buena Vista City 490.5 (378.8, 602.3) 3.5 (2.7, 4.3) 0.65 Highland 466.0 (367.9, 564.1) 3.3 (2.6, 4.0) 0.76 Buchanan 463.2 (362.3, 564.0) 3.3 (2.6, 4.0) 0.53 Fairfax City 462.8 (344.9, 580.8) 3.3 (2.4, 4.1) 1.71 Dickenson 454.0 (354.2, 553.9) 3.2 (2.5, 3.9) 0.53 Tazewell 441.8 (344.8, 538.8) 3.1 (2.4, 3.8) 0.62 Colonial Heights City 435.5 (340.3, 530.7) 3.1 (2.4, 3.7) 0.96
-(d) Top-10 counties in Virginia by per-household health cost
+data centers have already resulted in a total public health cost of about \$6.7 billion, or \$47.5 per household, in 2023.
+This is equivalent to approximately 44% of the data centers’ total electricity cost.
+Next, we show in Fig.
+5 the county-level total public health cost of U.S.
+data centers from 2019 to 2023, which exhibits significant spatial variability.
+In particular, populated counties located downwind of power plants supplying electricity to data centers tend to experience higher health costs, reflecting the transport of air pollutants across regions.
+The cumulative distribution function (CDF) in Fig.
+5b highlights that while many counties incur relatively low health costs, a small fraction of counties bear substantially higher impacts.
+Table 5c further identifies the top-10 counties with the highest total health costs, illustrating how local population density and proximity to power generation infrastructure combine to amplify public health risks in specific communities.
+State County Health Cost (million \$) IL Cook 358.5 (245.7, 471.3) PA Allegheny 322.4 (246.3, 398.6) TX Harris 306.7 (208.5, 404.8) OH Hamilton 195.5 (147.8, 243.1) PA Philadelphia 193.7 (137.5, 250.0) MI Wayne 188.6 (133.2, 244.1) OH Cuyahoga 174.7 (125.7, 223.8) TX Dallas 174.4 (131.2, 217.6) OH Franklin 166.2 (121.8, 210.6) NY Kings 164.5 (102.7, 226.3)
+Figure 5: The county-level total health cost of U.S. data centers from 2019 to 2023. (a) Health cost map; (b) CDF of county-level health cost; (c) Top-10 counties by total health cost.
+5.1.2 Uneven distribution of data centers’ public health impacts.
+Next, Fig.
+6 presents the county-level per-household total health costs attributable to U.S.
+data centers from 2019 to 2023.
+The results reveal a highly disproportionate distribution of health impacts across counties, with low-income communities particularly affected.
+The ratio of the highest to lowest county-level per-household health cost reaches approximately 200.
+Notably, all of the top 10 counties with the highest per-household costs have median household incomes below the national median.
+The high degree of disparity across different communities in terms of the public health cost suggests that we need to carefully examine the local and regional health impacts of data centers and improve public health equity to enable truly responsible computing.
+Furthermore, many of the hardest-hit communities neither host large data centers nor directly benefit economically from AI data centers, such as through tax revenues.
+For example, several counties in West Virginia are among the most affected, because many coal-fired power plants in West Virginia are supplying electricity to data centers in the neighboring state of Virginia [5, 4].
+By contrast, despite hosting a large number of data centers and incurring high total health costs due to its large population, California’s clean power grid results in some of the lowest per-household health impacts in the country.
+State County Per-household Health Cost  (\$) IR WV Marion 1218.3 (978.0, 1458.5) 0.80 WV Mason 1139.0 (897.6, 1380.4) 0.71 OH Meigs 1123.1 (840.8, 1405.5) 0.62 OH Gallia 1107.6 (828.1, 1387.2) 0.74 WV Marshall 1083.8 (831.2, 1336.5) 0.77 WV Taylor 1052.5 (853.1, 1252.0) 0.70 PA Fayette 992.6 (782.0, 1203.3) 0.74 PA Greene 944.9 (770.3, 1119.5) 0.88 WV Brooke 918.8 (693.2, 1144.4) 0.69 WV Jackson 871.9 (703.8, 1039.9) 0.73
 Figure 6: The county-level per-household health cost of U.S.
-data centers in 2030 based on exponential extrapolation of baseline emissions (McKinsey’s medium-growth forecast).
-The income data is based on the 2018-2022 American Community Survey 5-year estimates provided by [83].
-Next, we show in Fig.
-3 the county-level per-household health cost of U.S.
-data centers in 2030 based on exponential extrapolation under McKinsey’s medium-growth forecast.
-We see that the health cost is highly disproportionately distributed across different counties and communities, particularly affecting low-income communities.
-The ratio of the highest county-level per-household health cost to the lowest cost could be more than 200.
-Crucially, all the top-10 counties in the U.S.
-and 9 out of top-10 counties in Virginia (which has the largest concentration of data centers in the U.S.
-[5, 4]) have lower median household incomes than the national median value.
-Moreover, many of the hardest-hit communities do not have large data centers or directly receive economic benefits from AI data centers such as tax revenues.
-Yet, compared to the national average of about 1 month of electricity bill per year, the households in these communities could each suffer from health impacts equivalent to up to $\sim$8 months of their electricity bills.
-The high degree of disparity across different communities in terms of the public health cost suggests that we must examine the local and regional health impacts of AI data centers and improve public health equity to enable truly responsible AI.
-(a) Health cost map
-(b) CDF of county-level health cost
-State County Health Cost (million \$) IL Cook 309.6 (214.4, 404.7) PA Allegheny 266.4 (204.7, 328.1) TX Harris 262.0 (182.9, 341.1) MI Wayne 171.8 (123.0, 220.6) PA Philadelphia 169.4 (118.0, 220.8) OH Hamilton 166.5 (127.3, 205.7) NY Kings 165.2 (105.2, 225.2) OH Franklin 152.0 (112.2, 191.8) OH Cuyahoga 151.1 (110.6, 191.5) NY Queens 138.6 (90.7, 186.5)
-(c) Top-10 counties by total health cost
-Figure 4: The county-level total health cost of U.S. data centers in 2030 based on exponential extrapolation of baseline emissions (McKinsey’s medium-growth forecast).
-We also show the county-level total public health cost in Fig.
-4.
-Compared to the per-household health cost distribution in Fig.
-3, the county-level total health cost distribution is more aligned with the population distribution — despite the low per-household health cost, populous counties in California have a high total health cost.
-Nonetheless, some less populous counties (e.g., Hamilton County, Ohio) near coal and/or natural gas power plants are still significantly impacted and even more so than those (e.g., Loudoun County, Virginia) that have high concentrations of data centers.
-Table 3: The public health cost of training an AI model of the Llama-3.1 scale in Meta’s U.S. data centers.
-Location Electricity Price (¢/kWh) Electricity Cost (million \$) Health Cost (million \$) % of Electricity Cost Emission (Metric Ton)
+data centers from 2019 to 2023.
+(a) Per-household health cost map; (b) CDF of county-level per-household health cost; (c) Top-10 counties by per-household health cost.
+IR represents “County-to-Nation Per-Household Median Income Ratio.”
+5.1.3 Projection for 2028.
+According to a recent Lawrence Berkeley National Laboratory (LBNL) report [4], the U.S.
+data center electricity consumption is expected to increase from 4.4% of the total national electricity use in 2023 to 6.7–12.0% in 2028, depending on the growth trajectory of AI adoption.
+At the same time, the projected rise in peak power demand is accompanied by massive installations of onsite backup diesel generators to ensure reliability during grid contingencies [60].
+This substantial growth in electricity demand and onsite generation is expected to offset, and in fact outweigh, the gradual pollution emission intensity reductions anticipated from the power sector.
+As a result, the total public health costs attributable to data center operations are projected to potentially triple from 2023 to 2028.
+Quantitatively, based on the low- and high-growth scenarios considered in [4], the total public health impact of U.S.
+data centers is estimated to reach \$11.7 billion and \$20.9 billion in 2028, respectively.
+Under the high-growth scenario, the resulting health burden could rival or exceed that of on-road emissions in the largest U.S.
+state.
+This highlights the growing health externalities of U.S.
+data centers at a national scale.
+Table 2: The public health cost of training a large AI model in selected U.S. data centers.
+Location Electricity Price (¢/kWh) Electricity (million \$) Health Cost (million \$) % of Electricity Cost Emission (Metric Ton)
 PM2.5 (LA-NYC) NOx (LA-NYC) SO2
@@
 5.2 Public Health Impact of Generative AI Training
-We now study the health impact of training a generative AI model.
+We now study the health impact of a specific computing task.
 Specifically, we consider the training of an LLM and assume that the electricity consumption is the same as training Llama-3.1 recently released by Meta [92].
+As the scope-2 impact is dominant and the power allocated to train Llama-3.1 is unknown to determine scope-1 impacts, we focus on scope-2 health costs associated with the electricity consumption.
 While we use Meta’s Llama-3.1 training electricity consumption and U.S.
@@
 Combined with the spatial flexibility of model training, they suggest that AI model developers should take into account potential health impacts when choosing data center locations for training.
-5 Our Recommendations
-We provide our recommendations to address the increasing public health impact of AI.
-Recommendation 1: Standardization of Reporting Protocols
-Despite their immediate and tangible impacts on public health, criteria air pollutants have been entirely overlooked in AI model cards and sustainability reports published by technology companies [45, 44, 10].
-The absence of such critical information adds substantial challenges to accurately identifying specific AI data centers as a key root cause of public health burdens and could potentially pose hidden risks to public health.
-To enhance transparency and lay the foundation for truly responsible AI, we recommend standardization of reporting protocols for criteria air pollutants and the public health impacts across different regions.
-Concretely, criteria air pollutants can be categorized into three different scopes (Section 2.2), and reported following the greenhouse gas protocol widely adopted by technology companies [45, 10, 85].
-Just as addressing scope-2 and scope-3 carbon emissions is important for mitigating climate change, it is equally crucial to address scope-2 and scope-3 criteria air pollutants to promote public health throughout the power generation and hardware manufacturing processes in support of AI.
-For instance, power plants are dispatched based on real-time energy demand to ensure grid stability.
-As a result, only focusing on regulating scope-2 air pollutants at the power plant level fails to address the root cause — electricity consumption — and overlooks the potential of demand-side solutions.
-In contrast, recognizing scope-2 air pollutants and their associated public health impacts enables novel opportunities for health-informed AI, which, as detailed below, taps into demand-side flexibilities to holistically reduce AI’s adverse public health impacts.
-Recommendation 2: Health-informed AI
+5.3 Location-Dependent Public Health Impacts of Two Technology Companies
+We further highlight locational dependency of public health impacts by considering two major technology companies’ U.S.
+data center locations in 2023, excluding their leased colocation data centers whose locations are proprietary.
+We name these two companies A and B, respectively.
+These two companies do not have same data center locations.
+While company B discloses its per-location electricity usage [37], company A does not.
+Thus, we uniformly distribute company B’s North America electricity consumption over its U.S.
+data center locations based on its latest sustainability report [10].
+We consider location-based emission accounting without taking into account renewable energy credits these two companies apply to offset their grid electricity consumption (see “Location-based emission” in Section 4.1.2).
+We see from Fig.
+7 that the two companies have significant differences in terms of the per-household health cost distribution and most-affected counties.
+This is primarily due to the two companies’ different data center locations, and highlights the locational dependency of public health impacts.
+That is, unlike carbon emissions that have a similar effect regardless of the emission source locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source.
+Thus, technology companies should account for public health impacts when deciding where they build data centers, where they get electricity for their data centers, and where they install onsite renewables in order to best mitigate the adverse health effects.
+(a) Per-household health cost (Company A)
+(b) Per-household health cost (Company B)
+Figure 7: The county-level per-household health cost of two U.S. technology companies in 2023.
+6 Health-Informed Computing: Addressing Data Centers’ Public Health Impact
+In this section, we present Health-Informed Computing, a framework that explicitly incorporates public health impacts as a key optimization objective and strategically manages data center workloads to minimize adverse health outcomes while supporting broader sustainability goals.
+To mitigate the public health impact of computing, one straightforward approach is to focus solely on reducing the energy consumption (e.g., reducing AI model sizes [93, 94]).
+While reducing energy consumption is beneficial, overlooking the downstream public health impact of where and when energy is produced does not necessarily lead to minimized health burdens.
+For example, Table 2 demonstrates a 10x difference in health costs for training the same AI model across different locations.
+This highlights that health-informed and energy-aware computing, when combined, offer complementary benefits, leading to better public health outcomes.
+6.1 Opportunities for Health-Informed Computing
 Data centers, including those operated by major technology companies [10, 37], predominantly rely on grid electricity due to the practical challenges of installing on-site low-pollutant and low-carbon energy sources at scale.
 However, the spatial-temporal variations of scope-2 health costs (Fig.
-5) open up new opportunities to reduce the public health impact by exploiting the high scheduling flexibilities of AI training and inference workloads.
-For example, as further supported by EPRI’s recent initiative on maximizing data center flexibility for demand response [11], AI training can be scheduled in more than one data center, while multiple AI models with different resource-performance tradeoffs are often available to serve AI inference requests.
+8) open up new opportunities to reduce the public health impact by exploiting the high scheduling flexibilities of computing workloads (e.g., AI training).
+For example, as further supported by EPRI’s recent initiative on maximizing data center flexibility for demand response [11], AI training can be scheduled in more than one data center, while multiple AI models with different sizes are often available to serve AI inference requests, offering flexible resource-performance tradeoffs.
 To date, the existing data centers have mostly exploited such scheduling flexibilities for reducing electricity costs [95], carbon emissions [15], water consumption [96], and/or environmental inequity [97].
@@
 (c) The Pearson correlation between the per-region yearly average marginal health cost and carbon intensity is 0.292.
-To further confirm this point, we analyze the scope-2 marginal carbon intensity and public health cost for each unit of electricity generation across all the 114 U.S.
-regions between October 1, 2023, and September 30, 2024, provided by [71].
+To further confirm this point and highlight the potential of health-informed data center load shifting, we analyze the scope-2 marginal carbon intensity and public health cost for each unit of electricity generation across all the 114 U.S.
+regions between October 1, 2023, and September 30, 2024, provided by WattTime [77].
 [^7] The health cost signal provided by [77] only considers mortality from PM2.5, while COBRA includes a variety of health outcomes including asthma, lung cancer, and mortality from ozone, among others [22].
 The time granularity for data collection is 5 minutes.
+Here, we focus on marginal health impacts and carbon emissions for two main reasons: first, WattTime provides only real-time marginal health impact estimates [98]; and second, marginal signals are generally considered more useful for guiding energy load adjustments [99], which also explain why the EPA reports marginal health benefits per kWh (i.e., marginal health price) to inform energy demand changes [100].
 We show in Fig.
-5a the region-wise normalized interquartile ranges (IQR divided by the yearly average) for both public health costs and carbon emissions.
+8a the region-wise normalized interquartile ranges (IQR divided by the yearly average) for both public health costs and carbon emissions.
 The normalized IQR measures the spread of the time-varying health and carbon signals.
@@
 Likewise, in Fig.
-5b, the greater temporal variation of health costs is also supported by its greater normalized standard deviation (STD divided by the yearly average) in 90 out of the 114 U.S.
+8b, the greater temporal variation of health costs is also supported by its greater normalized standard deviation (STD divided by the yearly average) in 90 out of the 114 U.S.
 regions (79%).
 Next, we show in Fig.
-5c the weak spatial correlation (Pearson correlation coefficient: 0.292) between the yearly average health cost and carbon intensity across the 114 regions.
+8c the weak spatial correlation (Pearson correlation coefficient: 0.292) between the yearly average health cost and carbon intensity across the 114 regions.
 Furthermore, the normalized IQR of the health cost spatial distribution is 3.62x that of carbon emission spatial distribution (1.05 vs.
@@
 0.19).
-In other words, the health cost has a greater spatial spread than the carbon emission.
-These findings highlight that leveraging spatial-temporal variations in a health-aware manner could significantly reduce AI’s public health costs while still maintaining climate benefits.
-As a result, we advocate for a new research direction — health-informed AI.
-Specifically, decisions regarding the siting of AI data centers and the runtime scheduling of AI tasks should explicitly address their public health impacts.
-By judiciously accounting for and exploiting the spatial-temporal diversity of health costs, AI data centers can be optimized to minimize adverse public health impacts while supporting sustainability goals.
-Additionally, as the public health awareness serves as an effective implicit incentive (e.g., as demonstrated in the context of residential energy conservation [89]), AI data center operators can also leverage this approach by informing end users about the public health impacts of their AI usage.
-This may help extract additional user-side demand flexibilities as part of the recent efforts to maximize the overall data center load flexibility [11].
-Recommendation 3: Attention to All
-Counties and communities located near AI data centers or supplying electricity to them often experience most significant health burdens.
-Nonetheless, these health impacts can extend far beyond the immediate vicinity, affecting communities hundreds of miles away [34, 33].
-For example, the health impact of backup generators in northern Virginia can affect several surrounding states (Fig.
-1a) and even reach as far as Florida.
-While the health impact on communities where data centers operate is increasingly recognized, there has been very little, if any, attention paid to other impacted communities that bear substantial public health burdens.
-This disconnect leaves those communities to shoulder the public health cost of AI silently without receiving adequate support.
-To fulfill their commitment to social responsibility, we recommend technology companies holistically evaluate the cross-state public health burden imposed by their operations on all impacted communities, when deciding where they build data centers, where they get electricity for their data centers, and where they install renewables.
-Additionally, to quantify the health effects on impacted communities with greater accuracy for potential regulatory actions, we recommend further interdisciplinary research such as cross-state air quality dispersion, health economics, and health-informed computing.
-Recommendation 4: Promoting Public Health Equity
-The public health impact of AI is highly unevenly distributed across different counties and communities in the U.S., often disproportionately affecting low-income communities and potentially exacerbating socioeconomic inequities [90, 37].
-For example, as shown in Table 3c and 3d, all the top-10 counties in the U.S.
-and 9 out of top-10 counties in Virginia have lower median household incomes than the national median value.
-The ratio of the highest county-level per-household health cost to the lowest cost could be more than 200.
-Critically, minimizing the total health cost without considering equity can even reinforce existing inequities, similar to the way environmental inequities have been amplified [88].
-Therefore, it is imperative to address the substantial health impact disparities across communities and ensure that AI fosters public health equity rather than exacerbating inequities.
-6 Conclusion
-In this paper, we uncover and quantify the overlooked public health impact of AI.
-We present a general methodology to model air pollutant emissions across AI’s lifecycle, from chip manufacturing to data center operation.
-Our findings demonstrate that under McKinsey’s projection with a medium-growth scenario, the U.S.
-data centers in 2030 could contribute to nearly 1,300 deaths annually, resulting in a public health burden of more than \$20 billion which could even exceed that of on-road emissions of California.
-Importantly, these public health costs are unevenly distributed and disproportionately impact low-income communities, where the per-household health burden could be equivalent to nearly 8 months of electricity bills and 200x compared to other less-impacted counties.
-We recommend adopting a standard reporting protocol for criteria air pollutants and public health costs, paying attention to impacted communities, and implementing health-informed AI to mitigate these effects while promoting public health equity.
-Our study provides novel insights for the public, policymakers, and technology companies, enabling a more comprehensive cost-benefit analysis of AI’s impacts on society.
-We also call for further research to fully address the public health implications when developing powerful and responsible AI in the future.
-It is crucial to prioritize public health and ensure that the growth of AI does not exacerbate health burdens or negate the potential benefits AI can bring in improving public health.
-Acknowledgement
-The authors would like to thank Susan Wierman for providing comments on an initial draft of this paper.
-Appendix
-Appendix A Implementation Details
-We describe the evaluation methodology used for our empirical analysis.
-We use the latest COBRA (Desktop v5.1, as of October 2024) provided by the U.S.
-EPA [46] to study the public health impact of U.S.
-data centers in both 2023 and 2030.
-While COBRA uses a reduced-complexity air quality dispersion model based on a source-receptor matrix for rapid evaluation, its accuracy has been validated and the same or similar model has been commonly adopted in the literature for large-area air quality and health impact analysis [76, 78, 74, 79].
-We consider county-level air pollutant dispersion throughout the contiguous U.S., which is the area currently supported by COBRA [46].
-Note that cities considered county-equivalents for census purposes are also referred to as “counties” in COBRA.
-Throughout the paper, we use “county” without further specification.
-All the monetary values are presented in the 2023 U.S.
-dollars unless otherwise stated.
-We set the discount rate as 2% in COBRA as recommended by the EPA based on the U.S.
-Office of Management and Budget Circular No.
-A-4 guidance [46].
-When presenting a single value or a ratio (e.g., health-to-electricity cost ratio) if applicable, we use the midrange of the low and high estimates provided by COBRA.
-A.1 Estimation of 2030 Baseline Emissions
-For estimates in 2030, COBRA provides data for county-level population, health incidence, and valuation, but the baseline emissions are missing [46].
-Thus, we estimate the 2030 baseline emission by extrapolating the data for 2016, 2023, and 2028 provided by COBRA.
-Specifically, we consider three different extrapolation methods as follows.
-$\bullet$ Linear: For each pollutant type (e.g., PM2.5, SO2, and NOx) at each source, we apply a linear model $y=a\cdot t+b$, where $t$ is the year, to fit the 2016, 2023, and 2028 values and use the linear model to estimate the 2030 value.
-We also calculate the coefficient of determination, or $R^{2}$ score for each linear model.
-If $R^{2}$ is less than 0.5, we set the predicted 2030 value equal to the 2028 value.
-In addition, if the value is missing for a pollutant type at a source for any of the three years (2016, 2023, and 2028), we directly use the 2028 value as the 2030 value.
-$\bullet$ Exponential: The exponential extrapolation method is similar to the linear method.
-When the model $y=a\cdot(1+r)^{t}$ shows an exponentially decreasing trend from 2016 to 2028 (i.e., $r<0$), we apply the model to estimate the 2030 value.
-Nonetheless, when the trend from 2016 to 2028 is increasing (i.e., $r>0$), we roll back to a linear model for conservative estimates to avoid over-estimates resulting from an exponential model.
-$\bullet$ Unchanged: We directly apply the 2028 baseline emission data to 2030.
-We show in Table 4 and Table 5 the estimated total baseline emissions of air pollutants for electricity generation and on-road traffic in 2030 using different extrapolation methods.
-We also show the baseline emissions for 2016, 2023, and 2028 as provided by COBRA [46].
-By reducing a state’s on-road emissions to zero in COBRA, we obtain the corresponding public health cost in that state.
-Table 4: U.S. electricity generation baseline emissions from 2016 to 2030
-Year Electricity Generation Emission (Metric Ton)
-NOx SO2 PM2.5 VOC
-2016 1100575.41 1369417.44 111604.62 30250.76
-2023 711746.94 717409.25 110878.22 34311.54
-2028 695495.34 733437.11 110279.40 34446.71
-2030 (Linear) 682541.75 726267.77 119326.10 36903.77
-2030 (Exponential) 707846.63 751245.61 120870.45 37488.27
-Table 5: U.S. and California on-road baseline emissions from 2016 to 2030
-Year U.S. On-road Emission (Metric Ton) California On-road Emission (Metric Ton)
-NOx SO2 PM2.5 VOC NOx SO2 PM2.5 VOC
-2016 3293579.05 25001.53 106828.36 1680342.17 202427.66 1438.07 10197.26 89087.60
-2023 1588423.83 11325.07 65742.16 996965.92 98095.76 1280.27 8144.83 54141.57
-2028 1130369.84 10616.37 53455.43 758508.40 86573.30 1154.27 8276.27 44586.45
-2030 (Linear) 594848.10 6402.84 40555.54 545983.16 52560.36 1109.71 7583.01 33840.13
-2030 (Exponential) 925971.64 9009.44 47978.65 653737.61 68881.37 1122.56 7910.51 38536.35
-On-road emissions are categorized as the “Highway Vehicles” sector in COBRA and include both tailpipe exhaust and tire and brake wear.
-Thus, following the EPA and U.S.
-Department of Transportation classification [22, 91], PM2.5 resulting from road dust is not counted as emissions of highway vehicles in our study.
-If the PM2.5 from paved road dust (categorized as “Miscellaneous $\to$ Other Fugitive Dust $\to$ Paved Roads” in COBRA) is considered, California is still projected to have the highest state-wide public health cost of on-road vehicles among all the U.S.
-states in 2030.
-For example, by assuming exponential extrapolation and including \$7.6 billion attributed to paved road dust PM2.5, California is projected to have a total health cost of \$23.9 billion.
-Nonetheless, even by including paved road dust PM2.5, our finding still indicates that the public health cost of U.S.
-data centers (e.g., \$21.5 billion based on McKinsey’s projection) could be comparable to that of California’s on-road emissions in 2030.
-A.2 Evaluation of AI’s Public Health Impact (Scope 2)
-Due to the limited data available for scope-1 and scope-3 impacts, we mainly focus on the scope-2 health impacts unless otherwise specified.
-Thus, the locations of emission sources depend on the power plants supplying electricity to data centers.
-To evaluate the public health impacts of U.S.
-data centers, we consider both average attribution and marginal attribution methods for 2023.
-Nonetheless, since it is difficult, if not impossible, to obtain the marginal emission rate without knowing the actual dispatch decisions for the future, we only use the average attribution method for 2030.
-The two attribution methods are described as follows.
-$\bullet$ Average attribution: We first calculate the total data center electricity consumption $e_{DC}$ and the overall electricity consumption (including non-data center loads) $e_{Total}$ within each electricity region.
-The U.S.
-electricity grid is divided into 14 regions following the AVoided Emissions and geneRation Tool (AVERT, the latest version v4.3 as of October 2024) provided by the EPA [70].
-We use the state-level electricity consumption data for 2023 and 2030 provided by EPRI [5], and distribute state-level electricity consumption to relevant electricity regions following the state-to-region electricity apportionment used by AVERT.
-Note that the actual state-to-region electricity apportionment in 2030 may vary from the assumption in AVERT.
-Thus, we also consider an alternative apportionment to further evaluate the public health impact of U.S.
-data centers.
-Specifically, we consider a state-level electricity apportionment scenario in which each state is viewed as an electricity region.
-The evaluation results are shown in Appendix C and further reinforce our key finding that the health impact of U.S.
-data centers could rival that of on-road emissions in some of the largest U.S.
-states such as California.
-We calculate the percentage $x\%=\frac{e_{DC}}{e_{Total}}$ of the data center electricity consumption with respect to the overall electricity consumption for each electricity region.
-The relationship between the health impact and emission reduction in COBRA is approximately linear.
-Thus, we apply a reduction by $x\%$ to the baseline emissions of all the power plants within the respective electricity region in COBRA and estimate the corresponding county-level health impacts, including health outcomes and costs.
-When assessing the health impact of generative AI training, we follow the same approach, except for changing the total data center electricity consumption to the AI model training electricity consumption.
-Assuming a medium growth rate, McKinsey projects that the U.S.
-data center electricity demand (excluding cryptocurrency) will reach 606 TWh, or 11.7% of the U.S.
-national electricity demand, in 2030 [4].
-When using McKinsey’s projection, we only use its projected percentage of 11.7%.
-That is, we consider the EPRI’s projection of non-data center loads and scale up the EPRI’s projection of data center electricity demand to match the percentage of 11.7%.
-As a result, the 2030 U.S.
-data center electricity demand is 519 TWh, instead of 606 TWh, in our study under McKinsey’s projection.
-Nonetheless, as we apply a reduction by $x\%$ to the baseline emissions in COBRA, what matters most is the percentage, rather than the absolute electricity consumption by data centers.
-$\bullet$ Marginal attribution: We only consider marginal attribution for 2023.
-Specifically, we use the state-level data center electricity consumption [5] and run AVERT to calculate the resulting county-level marginal air pollutant reduction [70].
-AVERT allows a maximum of 15% electricity reduction within an electricity region during each hour.
-For regions where the data center electricity demand exceeds the 15% reduction threshold for certain hours in 2023, we cap the reduction at 15%, which results in a conservative estimate (i.e., the actual health impact of data centers is slightly higher).
-The county-level emission reduction data provided by AVERT is then applied to COBRA to estimate the county-level health outcomes and costs.
-Electricity price.
-When estimating the electricity cost for data centers in 2023 and 2030, we use the state-level average price for industrial users in [92].
-The projected U.S.
-nominal electricity price for industrial users remains nearly the same from 2023 to 2030 (24.96 \$/MMBtu in 2023 vs.
-23.04 \$/MMBTu in 2030) in the baseline case per the EIA’s Energy Outlook 2023 [39].
-Thus, our estimated health-to-electricity cost ratio will be even higher if we further adjust inflation.
-Similarly, to estimate the household electricity bills, we use the state-level average price for residential users and county-level average household electricity consumption in [92].
-Location-based emission.
-There are two types of scope-2 carbon emissions associated with electricity consumption: location-based and market-based [10].
-Specifically, location-based carbon emissions refer to the physical carbon emissions attributed to an electricity consumer connected to the power grid, while market-based carbon emissions are net emissions after applying reductions due to contractual arrangements and other credits (e.g., renewable energy credits).
-In this paper, similar to location-based carbon emissions commonly studied in the literature [8], we focus on criteria air pollutants for AI data centers without considering market-based pollution reduction mechanisms.
-While data centers, including large technology companies, often use various credits to reduce their market-based carbon emissions [10], it is likely less effective to apply this practice to mitigate the public health impact.
-The reason is that, unlike carbon emissions that have a similar effect on climate change regardless of the emission source locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source.
-For example, the public health impact of using grid power from a populated region may not be effectively mitigated by the renewable energy credits generated elsewhere.
-A.3 Public Health Impact of Backup Generators in Virginia
-Virginia has issued a total of 174 air quality permits for data center backup generators as of December 1, 2024 [56].
-More than half of the data center sites are within Loudoun County.
-We collect a dataset of the air quality permits: permits issued before January 1, 2023, from [57], and permits issued between January 1, 2023 and December 1, 2024, from [56].
-The total permitted site-level annual emission limits are approximately 13,000 tons of NOx, 1,400 tons of VOCs, 50 tons of SO2, and 600 tons of PM2.5, all in U.S.
-short tons.
-By assuming that the actual emissions are 10% of the permitted level, the data centers in Virginia could already cause approximately 14,000 asthma symptom cases and 13-19 deaths each year, among other health implications, resulting in a total annual public health burden of \$220-300 million, including \$190-260 million incurred in Virginia, West Virginia, Maryland, Pennsylvania, New York, New Jersey, Delaware, and Washington D.C., as estimated by COBRA under the “Fuel Combustion: Industrial” sector.
-A.4 Public Health Impact of a Semiconductor Facility
-We consider a semiconductor manufacturing facility located in Ocotillo, a neighborhood in Chandler, Arizona [93].
-By averaging the rolling 12-month air pollutant emission levels listed in the recent air quality monitoring report (as of October, 2024) [18], we obtain the annual emissions as follows: 150.4 tons of NOx, 82.7 tons of VOCs, 1.1 tons of SO2, and 28.9 tons of PM2.5.
-By applying these on-site emissions to COBRA under the “Other Industrial Processes” sector, we obtain a total public health cost of \$14-21 million.
-Additionally, the total annual energy consumption by the facility is 2074.88 million kWh as of Q2, 2024 [93].
-Assuming 84.2% of the energy comes from the electricity based on the company’s global average [94], we obtain the facility’s annual electricity consumption as 1746.63 million kWh.
-By using the average attribution method, we further obtain an estimated health cost of \$12-17 million associated with the electricity consumption.
-Thus, the total health cost of the facility is \$26-39 million.
-By relocating the facility from Chandler, Arizona, to a planned site in Licking County, Ohio, and assuming the same emission level and electricity consumption, we can obtain the total health cost of \$94-156 million, including \$23-36 million attributed to direct on-site emissions and \$70-120 million attributed to electricity consumption.
-A.5 Energy Consumption for Training a Generative AI Model
-We consider Llama-3.1 as an example generative AI model.
-According to the model card [44], the training process of Llama-3.1 (including 8B, 70B, and 405B) utilizes a cumulative of 39.3 million GPU hours of computation on H100-80GB hardware, and each GPU has a thermal design power of 700 watts.
-Considering Meta’s 2023 PUE of 1.08 [45] and excluding the non-GPU overhead for servers, we estimate the total training energy consumption as approximately 30 GWh.
-A.6 Average Emission for Each LA-NYC Round Trip by Car
-We use the 2023 national average emission rate for light-duty vehicles (gasoline) provided by the U.S.
-Department of Transportation [91].
-The emission rate accounts for tailpipe exhaust, tire wear and brake wear.
-Specifically, the average PM2.5 emission rate is 0.008 grams/mile (including 0.004 grams/mile for exhaust, 0.003 grams/mile for brake wear, and 0.001 grams/mile for tire wear), and the average NOx emission rate is 0.199 grams/mile for exhaust.
-We see that half of PM2.5 for light-duty vehicles comes from brake and tire wear (0.004 gram/miles), which are also produced by other types of vehicles including electric vehicles.
-The distance for a round-trip between Los Angeles, California, and New York City, New York, is about 5,580 miles.
-Thus, the average auto emissions for each LA-NYC round trip are estimated as 44.64 grams of PM2.5 and 1110.42 grams of NOx.
-Appendix B Public Health Impact of U.S. Data Centers in 2023
-We show in Fig. 6 the state-wide data center electricity consumption in 2023 [5]. It can be seen that Virginia, Texas and California have the highest data center electricity consumption in 2023.
-Next, we show in Fig.
-7 the county-level per-household (scope-2) health cost caused by the U.S.
-data centers in 2023.
-We see that the health cost is highly disproportionately distributed across different counties and communities, particularly affecting low-income communities.
-The ratio of the highest county-level per-household health cost to the lowest cost is more than 100.
-Crucially, all the top-10 counties in the U.S.
-have lower median household incomes than the national median value.
-Moreover, by comparing Fig.
-7 and Fig.
-6, we see that many of the hardest-hit communities do not have large data centers or directly receive economic benefits from AI data centers such as tax revenues.
-We also show in Fig.
-8 the county-level total health costs of U.S.
-data centers in 2023.
-(a) Electricity consumption map
-State Electricity Consumption (TWh) State Electricity Consumption (TWh) State Electricity Consumption (TWh) VA 33.85 OH 2.36 ID 0.15 TX 21.81 SC 2.02 WI 0.15 CA 9.33 WY 1.86 MD 0.10 IL 7.45 KY 1.62 LA 0.08 OR 6.41 CO 1.51 SD 0.07 AZ 6.25 AL 1.49 ME 0.03 IA 6.19 FL 1.38 NH 0.02 GA 6.18 TN 1.33 RI 0.02 WA 5.17 OK 1.23 KS $<$ 0.01 PA 4.59 MA 1.06 AR $<$ 0.01 NY 4.07 MO 0.97 DE $<$ 0.01 NJ 4.04 MN 0.82 DC $<$ 0.01 NE 3.96 MT 0.58 MS $<$ 0.01 ND 3.92 MI 0.53 VT $<$ 0.01 NV 3.42 NM 0.40 WV $<$ 0.01 NC 2.67 CT 0.26 UT 2.56 IN 0.19
-(b) Electricity consumption by state (descending order)
-Figure 6: State-level electricity consumption of U.S. data centers in 2023 [5].
-(a) Per-household health cost map
-(b) CDF of county-level per-household health cost
-State County Per-household Health Cost  (\$) Months of Electricity Bills County-to-nation Income Ratio WV Marion 306.0 (244.9, 367.1) 2.2 (1.8, 2.6) 0.80 WV Mason 299.4 (235.6, 363.1) 2.2 (1.7, 2.6) 0.71 OH Meigs 294.4 (220.0, 368.8) 2.4 (1.8, 3.0) 0.62 OH Gallia 289.9 (216.5, 363.3) 2.3 (1.7, 2.9) 0.74 WV Marshall 280.6 (215.6, 345.7) 2.0 (1.6, 2.5) 0.77 WV Taylor 266.6 (215.4, 317.7) 1.9 (1.6, 2.3) 0.70 PA Fayette 256.1 (201.9, 310.3) 1.8 (1.4, 2.2) 0.74 PA Greene 245.4 (200.2, 290.5) 1.7 (1.4, 2.0) 0.88 WV Brooke 235.7 (177.9, 293.5) 1.7 (1.3, 2.1) 0.69 WV Jackson 227.1 (183.3, 270.9) 1.6 (1.3, 2.0) 0.73
-(c) Top-10 counties by per-household health cost
-Figure 7: The county-level per-household health cost of U.S. data centers in 2023.
-(a) Health cost map
-(b) CDF of county-level health cost
-State County Health Cost (million \$) IL Cook 81.9 (55.8, 107.9) PA Allegheny 80.1 (61.1, 99.2) TX Harris 68.3 (46.4, 90.2) OH Hamilton 49.0 (37.1, 60.9) PA Philadelphia 48.1 (33.4, 62.7) MI Wayne 46.8 (33.0, 60.6) OH Cuyahoga 43.1 (31.1, 55.2) NY Kings 41.7 (26.1, 57.3) OH Franklin 41.5 (30.3, 52.7) NY Queens 35.2 (22.5, 47.8)
-(c) Top-10 counties by health cost
-Figure 8: The county-level health cost of U.S. data centers in 2023.
-We show in Fig.
-9 the per-household health cost of U.S.
-data centers in 2023 by considering the marginal attribution method.
-The health cost using marginal attribution means the public health burden resulting from the additional loads of the U.S.
-data centers connected to the grid in 2023.
-In other words, if the U.S.
-data centers had been powered using off-grid sources (e.g., on-site renewables) in 2023, the per-household public health benefit would be valued at up to \$319 and the total public health benefit would be \$7.6 billion.
-Table 6: The public health cost of U.S.
-data centers in 2030.
-“” denotes McKinsey’s projection under a medium growth rate (excluding energy consumption for cryptocurrency) [4].
-State-level electricity apportionment.
-Extrapolation Method Projected Growth Electricity (TWh) Electricity Cost (billion \$) Mortality Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$) Months of Electricity Bill % of CA On-road Health Cost
-Unchanged Low 196.3 16.8 370 (270, 460) 6.3 (4.8, 7.8) 37% 42.0 (31.8, 52.1) 0.31 (0.23, 0.38) 33%
-Moderate 214.0 18.3 400 (300, 500) 6.8 (5.1, 8.4) 37% 45.4 (34.4, 56.3) 0.33 (0.25, 0.41) 36%
-High 296.4 25.4 530 (400, 660) 9.0 (6.9, 11.2) 36% 60.4 (45.8, 75.0) 0.44 (0.33, 0.55) 47%
-Higher 403.9 34.6 690 (510, 860) 11.8 (8.9, 14.6) 34% 78.6 (59.6, 97.6) 0.57 (0.44, 0.71) 62%
-Medium 519.3 44.5 890 (660, 1110) 15.1 (11.5, 18.8) 34% 101.0 (76.6, 125.4) 0.74 (0.56, 0.92) 79%
-Linear Low 196.3 16.8 370 (280, 460) 6.3 (4.8, 7.8) 38% 42.2 (32.0, 52.4) 0.31 (0.23, 0.38) 46%
-Moderate 214.0 18.3 400 (300, 500) 6.8 (5.2, 8.5) 37% 45.6 (34.5, 56.6) 0.33 (0.25, 0.41) 49%
-High 296.4 25.4 530 (400, 670) 9.1 (6.9, 11.3) 36% 60.7 (46.0, 75.5) 0.44 (0.34, 0.55) 66%
-Higher 403.9 34.6 690 (520, 870) 11.8 (9.0, 14.7) 34% 79.0 (59.8, 98.2) 0.58 (0.44, 0.72) 86%
-Medium 519.3 44.5 890 (660, 1120) 15.2 (11.5, 18.9) 34% 101.6 (76.9, 126.2) 0.74 (0.56, 0.92) 110%
-Exponential Low 196.3 16.8 380 (290, 480) 6.5 (4.9, 8.1) 39% 43.6 (33.0, 54.2) 0.32 (0.24, 0.40) 40%
-Moderate 214.0 18.3 410 (310, 520) 7.0 (5.3, 8.8) 38% 47.1 (35.7, 58.5) 0.34 (0.26, 0.43) 43%
-High 296.4 25.4 550 (410, 690) 9.4 (7.1, 11.7) 37% 62.7 (47.5, 78.0) 0.46 (0.35, 0.57) 57%
-Higher 403.9 34.6 720 (530, 900) 12.2 (9.2, 15.2) 35% 81.6 (61.8, 101.4) 0.60 (0.45, 0.74) 75%
-Medium 519.3 44.5 920 (690, 1160) 15.7 (11.9, 19.5) 35% 104.9 (79.4, 130.4) 0.77 (0.58, 0.95) 96%
-(a) Per-household health cost map
-(b) CDF of county-level per-household health cost
-State County Per-household Health Cost (\$) Moths of Electricity Bills County-to-nation Income Ratio WV Mason 319.8 (258.1, 381.5) 2.3 (1.9, 2.8) 0.71 OH Meigs 308.5 (235.2, 381.8) 2.5 (1.9, 3.1) 0.62 OH Gallia 299.6 (229.0, 370.2) 2.4 (1.8, 3.0) 0.74 WV Brooke 285.3 (213.3, 357.2) 2.1 (1.5, 2.6) 0.69 WV Marshall 270.7 (204.2, 337.2) 2.0 (1.5, 2.4) 0.77 PA Fayette 254.3 (195.8, 312.8) 1.8 (1.4, 2.2) 0.74 WV Marion 252.8 (194.8, 310.8) 1.8 (1.4, 2.2) 0.80 WV Jackson 252.2 (206.8, 297.7) 1.8 (1.5, 2.1) 0.73 WV Hancock 252.1 (193.1, 311.1) 1.8 (1.4, 2.2) 0.77 WV Roane 241.4 (196.8, 286.0) 1.7 (1.4, 2.1) 0.55
-(c) Top-10 counties by per-household health cost
-Figure 9: The county-level per-household health cost of U.S. data centers in 2023. Marginal attribution.
-Appendix C Public Health Impact of U.S. Data Centers in 2030 (State-level Electricity Apportionment)
-AVERT [70] divides the U.S.
-electricity grid into 14 regions.
-Since the actual state-to-region electricity apportionment in 2030 may vary from the assumption in AVERT, we now consider an alternative apportionment to further evaluate the public health impact of U.S.
-data centers in 2030.
-Specifically, we hypothesize a state-level electricity apportionment scenario in which each state is viewed as an electricity region (i.e., data centers are powered by in-state electricity).
-We show the results in Table 6, Fig.
-10, and Fig.
-11.
-While the actual values slightly differ from those in Section 4.2, the key message remains the same: the health impact of U.S.
-data centers could rival that of on-road emissions in some of the largest U.S.
-states such as California, and disproportionately affect low-income communities.
-As we consider in-state electricity to power data centers, 9 out of 10 most-effected counties in terms of the per-household public health burden are in Virginia which has the largest concentration of data centers [5].
-(a) Linear extrapolation
-(b) Exponential extrapolation
-(c) 2028-to-2030 unchanged
-Figure 10: The health costs of U.S.
-data centers and top-3 state on-road emissions from 2016 to 2030 based on different extrapolations for 2030 baseline emissions.
-(State-level electricity apportionment.)
-(a) Per-household health cost map
-(b) CDF of county-level per-household health cost
-State County Per-household Health Cost (\$) Months of Electricity Bills County-to-nation Income Ratio VA Emporia City 882.6 (623.3, 1141.8) 6.2 (4.4, 8.1) 0.55 VA Colonial Heights City 661.4 (516.2, 806.5) 4.7 (3.6, 5.7) 0.96 VA Brunswick 621.2 (441.8, 800.6) 4.4 (3.1, 5.7) 0.70 VA Greensville 607.1 (408.1, 806.1) 4.3 (2.9, 5.7) 0.69 VA Hopewell City 603.8 (466.9, 740.7) 4.3 (3.3, 5.2) 0.67 VA Hanover 593.7 (455.7, 731.6) 4.2 (3.2, 5.2) 1.39 VA Lancaster 580.3 (465.9, 694.7) 4.1 (3.3, 4.9) 0.83 VA Petersburg City 576.0 (429.3, 722.7) 4.1 (3.0, 5.1) 0.62 ND McLean 565.7 (451.1, 680.2) 4.8 (3.8, 5.8) 1.07 VA Sussex 563.1 (397.7, 728.4) 4.0 (2.8, 5.1) 0.79
-(c) Top-10 counties by per-household health cost
-Figure 11: The county-level per-household health cost of U.S.
-data centers in 2030 based on exponential extrapolation of baseline emissions (McKinsey’s medium-growth forecast).
-The income data is based on the 2018-2022 American Community Survey 5-year estimates provided by [83].
-(State-level electricity apportionment.)
-Appendix D Health-informed AI
-We now provide additional results to highlight the importance of health-informed AI.
-D.1 Correlation Analysis of Marginal Carbon Intensity and Health Impact for Google’s U.S. Data Center Locations
-In addition to the analysis in Section 5, we study the scope-2 marginal carbon intensity and public health cost for each unit of electricity generation across Google’s U.S.
-data center between October 1, 2023, and September 30, 2024, provided by [71].
-The health cost signal provided by [71] only considers mortality from PM2.5, while COBRA includes a variety of health outcomes including asthma, lung cancer, and mortality from Ozone, among others [22].
-The time granularity for data collection is 5 minutes.
-We present the results Table 7, further confirming that carbon intensities and health impacts are not always aligned and that health impacts vary more significantly than carbon intensities in almost all the locations.
-This suggests that, by judiciously accounting for and exploiting the spatial-temporal diversity of health costs, AI data centers can be optimized to minimize adverse public health impacts while supporting sustainability goals.
+In other words, the health cost could have a greater spatial spread than the carbon emission.
 Location Pearson Correlation Normalized IQR Normalized STD
@@
 Mayes County, OK 0.617 0.122 0.049 2.495 0.171 0.222 0.772
-Table 7: Correlation analysis of marginal carbon emissions and health impacts for Google’s U.S.
+Table 3: Correlation analysis of marginal carbon emissions and health impacts for a technology company’s U.S.
 data center locations between October 1, 2023, and September 30, 2024 [77].
 According to the region classification of WattTime [98], the two data centers in Storey County, NV, and Henderson, NV, belong to the same power grid region, and so do those in Jackson County, AL, and Montgomery County, TN.
-D.2 Location-dependent Public Health Impact
-We now show the location-dependent public health impacts of two technology companies based on Google’s and Meta’s U.S.
-data center locations in 2023, excluding their leased colocation data centers whose locations are proprietary.
-Due to the lack of information about the per-data center electricity consumption, we uniformly distribute Google’s North America electricity consumption over its U.S.
-data center locations based on Google’s latest sustainability report [10].
-Meta discloses its per-location electricity usage [45].
-We consider criteria air pollutants without accounting for renewable energy credits these two companies apply to offset their grid electricity consumption (see “Location-based emission” in Appendix A.2).
-As a consequence, although we consider the U.S.
-data center locations of Google and Meta, our results should not be interpreted as a quantitative evaluation of these two specific companies’ actual public health impacts.
-We also emphasize that our goal is to highlight the locational dependency of public health impacts and to motivate the need for health-informed siting of data centers.
-In our results, we refer to Google and Meta as Company A and Company B, respectively, to avoid potential misunderstandings.
-We first see from Table 8 that while the two companies have different public health costs due to their different electricity consumption, their health-to-electricity cost ratios are similar at the national level.
-Nonetheless, we notice from Fig.
-12 that the two companies have significant differences in terms of the per-household health cost distribution and most-affected counties.
-This is primarily due to the two companies’ different data center locations, and highlights the locational dependency of public health impacts.
-That is, unlike carbon emissions that have a similar effect on climate change regardless of the emission source locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source.
-Thus, technology companies should account for public health impacts when deciding where they build data centers, where they get electricity for their data centers, and where they install renewables in order to best mitigate the adverse health effects while promoting equity.
-Table 8: The public health costs based on two technology companies’ U.S. data center electricity consumption in 2023.
-Company (Attribution) Electricity (TWh) Electricity Cost (billion \$) Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$)
-A (Average) 18.5 1.4 0.63 (0.47, 0.78) 45% 4.5 (3.4, 5.5)
-A (Marginal) 18.5 1.4 0.97 (0.75, 1.20) 70% 6.9 (5.3, 8.6)
-B (Average) 10.6 0.8 0.38 (0.29, 0.48) 51% 2.7 (2.0, 3.4)
-B (Marginal) 10.6 0.8 0.53 (0.41, 0.66) 71% 3.8 (2.9, 4.7)
-(a) Per-household health cost (Company A)
-(b) Per-household health cost (Company B)
-(c) CDF of per-household health cost (Company A)
-(d) CDF of per-household health cost (Company B)
-State County Per-household Health Cost (\$) County-to-nation Income Ratio TX Marion 33.8 (27.2, 40.4) 0.64 VA Mecklenburg 23.7 (19.3, 28.1) 0.68 VA Halifax 23.5 (18.8, 28.1) 0.65 NC Person 23.5 (19.4, 27.6) 0.81 VA Martinsville City 22.5 (18.6, 26.5) 0.52 VA Danville City 21.7 (17.3, 26.1) 0.55 TX Cass 21.2 (16.5, 25.8) 0.72 GA Pickens 21.1 (17.3, 24.9) 0.97 WV Marion 20.5 (16.3, 24.6) 0.80 VA Henry 20.4 (15.9, 24.8) 0.58
-(e) Top-10 counties by per-household health cost (Company A)
-State County Per-household Health Cost (\$) County-to-nation Income Ratio TX Marion 21.1 (17.0, 25.3) 0.64 TX Cass 13.3 (10.4, 16.3) 0.72 WV Marion 12.4 (9.9, 15.0) 0.80 GA Pickens 12.4 (10.1, 14.6) 0.97 WV Marshall 12.1 (9.2, 14.9) 0.77 WV Mason 12.0 (9.4, 14.6) 0.71 TX Gregg 12.0 (9.6, 14.5) 0.85 TX Harrison 12.0 (9.7, 14.3) 0.84 TX Morris 12.0 (9.3, 14.6) 0.69 OH Gallia 11.9 (8.9, 14.9) 0.74
-(f) Top-10 counties by per-household health cost (Company B)
-Figure 12: The county-level per-household health cost of two companies in 2023.
-The income data is based on the 2018-2022 American Community Survey 5-year estimates provided by [83].
-Average attribution.
+In addition to analysis for all U.S.
+regions, we turn to specific regions where a large technology company builds its U.S.
+data centers.
+We present the results Table 3, further confirming that carbon intensities and health impacts are not always aligned and that health impacts vary more significantly than carbon intensities in almost all the locations.
+We further analyze the Pearson correlation coefficients between hourly marginal health prices and carbon emission rates throughout 2023 for the U.S.
+regions that have complete health and carbon data provided by WattTime [77].
+The CDF of the correlation coefficients is shown in Figure 9a.
+We see that nearly 70% of the regions have a weak or moderate correlation, with a carbon-health correlation coefficient of less than 0.60.
+This implies that despite having fossil fuels as the common source, health costs and carbon emissions are different and can exhibit trade-offs.
+Moreover, due to their additional dependence on population distribution and meteorological conditions, health prices demonstrate more pronounced temporal fluctuations than carbon emissions.
+These findings highlight that leveraging spatiotemporal variations in a health-aware manner can substantially reduce the public health costs of data center operations.
+Moreover, the observed distinctions between health impacts and carbon emissions suggest the need to optimize data center decisions by explicitly accounting for and exploiting the spatiotemporal heterogeneity of health impacts.
+6.2 Benefits of Health-Informed Computing
+To improve system performance and reliability, technology companies typically operate data centers over a variety of geographically distributed regions and dynamically distribute computing workloads through a process known as geographical load balancing (GLB).
+Here, we leverage the unique spatial load flexibility of geographically distributed data centers to demonstrate the benefits of health-informed computing as a proof of concept.
+Health-Informed GLB.
+Specifically, we study health-informed GLB (HI-GLB) to address the public health burden of data center operation.
+We consider a discrete-time model of duration $T$ and measure the workloads in terms of their energy demand.
+For the sake of examining the impact of spatial flexibility, we assume that the energy loads (i.e., workloads) can be flexibly distributed across a set of $N$ data centers denoted as $\mathcal{N}=\{1,2,\ldots,N\}$.
+In each time $t\in\{1,2,\ldots,T\}$, the total energy demand is $M_{t}$, and $w_{i,t}$ represents the load assigned to data center $i$.
+We use $l_{i}$ to represent the default load capacity of data center $i$, and introduce a slackness parameter $\lambda\geq 1$ to represent the ability of each data center to accept loads in excess of its default capacity.
+Thus, the greater the value of $\lambda$, the more spatial flexibility the operator has.
+In each time $t$, we use $p_{i,t}^{e}$ and $p_{i,t}^{h}$ to denote the electricity price and health price at data center $i$, respectively.
+The health price $p_{i,t}^{h}$ depends on the emission source of criteria air pollutants (e.g., power plants’ emission rates and their locations), air pollutant dispersion, and estimates of adverse health outcomes and resulting costs attributed to the increased air pollutant concentration in each region [22].
+Thus, the health price quantifies the ultimate health burden imposed on affected populations and is measured in terms of economic costs for each unit of electricity consumption.
+It varies over time due to fluctuations in the grid’s generation mix and changing meteorological conditions.
+Third-party organizations such as WattTime [98] provide real-time estimates of the marginal health price of electricity across 114 power balancing regions in the U.S., while the EPA [100] reports annualized average health prices for electricity in 14 broader regions nationwide.
+Our goal is to minimize the sum of electricity costs $\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i,t}^{e}\cdot w_{i,t}$ and health costs $\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i,t}^{h}\cdot w_{i,t}$ across all regions.
+Thus, HI-GLB is formulated as follows:
+$\min_{{w}}$ $\qquad\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i,t}^{e}\cdot w_{i,t}+\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i,t}^{h}\cdot w_{i,t}$ (4a)
+$\mathrm{s.t.}$ $\qquad\sum_{i=1}^{N}w_{i,t}=M_{t},\quad\forall t\in\mathcal{T}$ (4b)
+$\qquad 0\leq w_{i,t}\leq\lambda\cdot l_{i},\quad\forall i\in\mathcal{N},\forall t\in\mathcal{T}$ (4c)
+where the constraint (4b) means that all loads must be dispatched to a data center (with no loads dropped) and the constraint (4c) encodes the maximum workload capacity of each data center.
+Our formulation can be easily extended to incorporate additional considerations such as load dispatching distance constraints and other metrics such as carbon emissions.
+Additionally, it can also include long-term, per-region health impact constraints, rather than focusing solely on national-level total health costs.
+For the sake of clarity, we set these extensions aside to focus on the novel metric of health cost for data center resource management.
+[…]
+For each data center location, we use the health price provided by WattTime [77] and the industry electricity price from the EIA [102].
+To quantify the carbon emissions of each location, we consider the marginal carbon emissions rate provided by WattTime [77].
+We specifically use marginal carbon emissions for a fair comparison with the marginal health price, and since this metric often better reflects the actual impact on system-wide carbon emissions resulting from demand-side load shifting [99].
+More details can be found in Appendix B.1.
+Table 4: Comparison between health-informed and carbon-aware GLB ($\lambda=1.5$)
+Metric Baseline Carbon-Aware GLB Health-Informed GLB
+$p^{c}=$ \$0/ton $p^{c}=$ \$5/ton $p^{c}=$ \$200/ton $p^{c}=$ $\infty$/ton HI-GLB ($p_{i,t}^{e}=0$) HI-GLB
+Health (Million \$) 393.23 416.29 (5.86%) 416.29 (5.86%) 383.32 (-2.52%) 404.26 (2.80%) 289.67 (-26.34%) 291.94 (-25.76%)
+Energy (Million \$) 756.50 714.99 (-5.49%) 714.99 (-5.49%) 734.77 (-2.87%) 765.68 (1.21%) 741.49 (-1.98%) 733.66 (-3.02%)
+Carbon (Million Ton) 6.60 6.67 (1.02%) 6.67 (1.02%) 6.20 (-6.01%) 6.12 (-7.23%) 6.54 (-0.89%) 6.51 (-1.38%)
+Figure 9: Correlation analysis.
+(a) CDF of correlation coefficients between hourly health prices and marginal carbon emission rates for all the U.S.
+regions; (b) Scatter plot of health price and marginal carbon emission rate (annual average in 2023) across Meta’s U.S.
+data center locations.
+Health-Informed vs.
+Carbon-Aware GLB.
+We show results comparing GLB with health and carbon costs in Table 4.
+The results highlight that optimizing solely for environmental metrics like carbon does not effectively mitigate health impacts.
+Concretely, as the carbon price varies, the health cost ranges between \$380 million and \$420 million, whereas the baseline health cost is approximately \$393 million.
+When $p_{c}=\mathdollar 200$/ton, the health cost decreases by 2.52% relative to the baseline.
+However, for all other carbon prices, the health cost increases by using carbon-aware GLB.
+Among these, the algorithm that focuses solely on minimizing carbon ($p_{c}=\infty$) leads to a 2.80% increase in health cost compared to the baseline.
+This stands in sharp contrast to HI-GLB, which achieves a substantial $\sim$26% reduction in health cost compared to the baseline, demonstrating that environmental metrics such as carbon are insufficient for mitigating health impacts unless public health considerations are explicitly integrated into the optimization process.
+Moreover, compared to the pure health-informed GLB that sets a zero electricity price ($p^{e}=0$), HI-GLB only slightly increases the health cost from \$290 million to \$292 million, further demonstrating that considering the health price can effectively drive healthier GLB decisions while maintaining low cost.
+At the same time, the pure carbon-aware GLB algorithm achieves a maximum reduction in carbon reduction of 7.23% while increasing the health burden and electricity cost relative to the baseline.
+In contrast, HI-GLB achieves 26% health cost reduction, 3% electricity cost saving, and more than 1% carbon reduction.
+This highlights the necessity of adopting health-informed algorithms for health burden reduction while offering co-benefits of cost saving and carbon reduction.
+We also vary the capacity slackness $\lambda$ to change the spatial flexibility and obtain similar insights. The full details are available in Appendix B.2.
+Spatial correlation analysis.
+To further explain the results, we analyze the correlation between health costs and carbon emissions across the 13 regions where Meta operates its U.S.
+data centers.
+Figure 9b presents a scatter plot illustrating the correlation between the annual average carbon emission rate and health price across various regions.
+The Pearson correlation coefficient between different locations’ health prices and carbon emissions is approximately -0.35, indicating a negative relationship.
+This suggests a potential conflict between efforts to optimize carbon emissions and those aimed at improving health outcomes via GLB.
+Furthermore, as highlighted by the spatial patterns in Figure 9b, the health prices exhibit a significantly higher degree of spatial variability compared to carbon emissions across regions.
+This discrepancy further reinforces the point that focusing solely on optimizing environmental factors, such as carbon emissions, may not effectively reduce health costs, underscoring the importance of integrating health-informed optimization strategies to achieve more comprehensive benefits.
+7 Our Recommendations
+We provide additional recommendations to address the growing public health impact of data centers.
+Recommendation 1: Standardization of Reporting Protocols
+Despite their immediate and tangible impacts on public health, criteria air pollutants have been entirely overlooked in AI model cards and sustainability reports published by technology companies [37, 36, 10].
+The absence of such critical information adds substantial challenges to accurately identifying specific AI data centers as a key root cause of public health burdens and could potentially pose hidden risks to public health.
+To enhance transparency and lay the foundation for truly responsible AI, we recommend standardization of reporting protocols for criteria air pollutants and the public health impacts across different regions.
+Concretely, criteria air pollutants can be categorized into three different scopes (Section 3), and reported following the greenhouse gas protocol widely adopted by technology companies [37, 10, 78].
+Just as addressing scope-2 and scope-3 carbon emissions is important for mitigating climate change, it is equally crucial to address scope-2 and scope-3 criteria air pollutants to promote public health throughout the power generation and hardware manufacturing processes in support of AI.
+For instance, power plants are dispatched based on real-time energy demand to ensure grid stability.
+As a result, only focusing on regulating scope-2 air pollutants at the power plant level fails to address the root cause — electricity consumption — and overlooks the potential of demand-side solutions.
+In contrast, recognizing scope-2 air pollutants and their associated public health impacts enables novel opportunities for health-informed AI, which, as detailed below, taps into demand-side flexibilities to holistically reduce AI’s adverse public health impacts.
+Recommendation 2: Attention to All
+Counties and communities located near AI data centers or supplying electricity to them often experience most significant health burdens.
+Nonetheless, these health impacts can extend far beyond the immediate vicinity, affecting communities hundreds of miles away [25, 24].
+For example, the health impact of backup generators in northern Virginia can affect several surrounding states (Fig.
+2a) and even reach as far as Florida.
+While the health impact on communities where data centers operate is increasingly recognized, there has been very little, if any, attention paid to other impacted communities that bear substantial public health burdens.
+This disconnect leaves those communities to shoulder the public health cost of AI silently without receiving adequate support.
+To fulfill their commitment to social responsibility, we recommend technology companies holistically evaluate the cross-state public health burden imposed by their operations on all impacted communities, when deciding where they build data centers, where they get electricity for their data centers, and where they install renewables.
+Additionally, to quantify the health effects on impacted communities with greater accuracy for potential regulatory actions, we recommend further interdisciplinary research such as cross-state air quality dispersion, health economics, and health-informed computing.
+Recommendation 3: Promoting Public Health Equity
+The public health impact of AI is highly unevenly distributed across different counties and communities, disproportionately affecting certain (often low-income) communities [103, 31].
+For example, as shown in Table 6c, all the top-10 most impacted counties in the U.S.
+have lower median household incomes than the national median value.
+The ratio of the highest county-level per-household health cost to the lowest cost is approximately 200.
+Therefore, it is imperative to address the substantial health impact disparities across communities.
+8 Conclusion
+In this paper, we quantify and address the overlooked public health impact of data centers.
+We introduce a principled methodology to model these lifecycle pollutant emissions and quantify their associated public health impacts.
+Our findings suggest that the total annual public health burden of U.S.
+data centers could exceed \$20 billion by 2028, approaching or even surpassing the impacts of on-road vehicle emissions in California.
+Importantly, these health costs are not evenly distributed: disadvantaged communities bear a disproportionate share, with per-household impacts potentially up to 200 times higher than in less-affected areas.
+This highlights the need for targeted mitigation strategies.
+To this end, we propose health-informed computing, a framework that explicitly incorporates public health risk as a key metric when scheduling data center workloads, enabling more informed and equitable operational decisions.
+More broadly, we recommend the adoption of standardized reporting protocols for the public health costs of data centers, alongside policies that ensure attention to all impacted communities, thereby supporting responsible, sustainable, and inclusive deployment of AI infrastructure.
+Acknowledgement
+The authors would like to thank Susan Wierman for providing comments on an initial draft of this paper.
+Appendix
+Appendix A Modeling Details
+We describe the evaluation methodology used for our empirical analysis.
+We use the latest COBRA (Desktop v5.1, as of October 2024) provided by the U.S.
+EPA [39] to study the public health impact of U.S.
+data centers in both 2019-2023 and 2028.
+While COBRA uses a reduced-complexity air quality dispersion model based on a source-receptor matrix for rapid evaluation, its accuracy has been validated and the same or similar model has been commonly adopted in the literature for large-area air quality and health impact analysis [85, 87, 82, 88].
+We consider county-level air pollutant dispersion throughout the contiguous U.S., which is the area currently supported by COBRA [39].
+Note that cities considered county-equivalents for census purposes are also referred to as “counties” in COBRA.
+Throughout the paper, we use “county” without further specification.
+All the monetary values are presented in the 2023 U.S.
+dollars unless otherwise stated.
+We set the discount rate as 2% in COBRA as recommended by the EPA based on the U.S.
+Office of Management and Budget Circular No.
+A-4 guidance [39].
+When presenting a single value or a ratio (e.g., health-to-electricity cost ratio) if applicable, we use the midrange of the low and high estimates provided by COBRA.
+COBRA provides data for county-level population, health incidence, valuation, and baseline emissions for 2016, 2023, and 2028 [39].
+For the data from 2019 to 2022, we use linear interpolation as recommended by the EPA’s COBRA team.
+We show in Table 5 and Table 6 the total baseline emissions of air pollutants for electricity generation and on-road traffic provided by COBRA [39].
+By reducing a state’s on-road emissions to zero in COBRA, we obtain the corresponding public health cost in that state.
+Table 5: U.S. electricity generation baseline emissions from 2016 to 2028
+Year Electricity Generation Emission (Metric Ton)
+NOx SO2 PM2.5 VOC
+2016 1100575.41 1369417.44 111604.62 30250.76
+2023 711746.94 717409.25 110878.22 34311.54
+2028 695495.34 733437.11 110279.40 34446.71
+Table 6: U.S. and California on-road baseline emissions from 2016 to 2028
+Year U.S. On-road Emission (Metric Ton) California On-road Emission (Metric Ton)
+NOx SO2 PM2.5 VOC NOx SO2 PM2.5 VOC
+2016 3293579.05 25001.53 106828.36 1680342.17 202427.66 1438.07 10197.26 89087.60
+2023 1588423.83 11325.07 65742.16 996965.92 98095.76 1280.27 8144.83 54141.57
+2028 1130369.84 10616.37 53455.43 758508.40 86573.30 1154.27 8276.27 44586.45
+On-road emissions are categorized as the “Highway Vehicles” sector in COBRA and include both tailpipe exhaust and tire and brake wear.
+Thus, following the EPA and U.S.
+Department of Transportation classification [22, 104], PM2.5 resulting from road dust is not counted as emissions of highway vehicles in our study.
+If the PM2.5 from paved road dust (categorized as “Miscellaneous $\to$ Other Fugitive Dust $\to$ Paved Roads” in COBRA) is considered, California is still projected to have the highest state-wide public health cost of on-road vehicles among all the U.S.
+states.
+Electricity price.
+When estimating the electricity cost for data centers in 2023 and 2038, we use the state-level average price for industrial users in [102].
+The projected U.S.
+nominal electricity price for industrial users remains nearly the same from 2023 to 2030 (24.96 \$/MMBtu in 2023 vs.
+23.04 \$/MMBTu in 2030) in the baseline case per the EIA’s Energy Outlook 2023 [105].
+Thus, our estimated health-to-electricity cost ratio will be even higher if we further adjust inflation.
+A.1 Public Health Impact of Backup Generators in Virginia
+Virginia has issued a total of 174 air quality permits for data center backup generators as of December 1, 2024 [60].
+More than half of the data center sites are within Loudoun County.
+We collect a dataset of the air quality permits: permits issued before January 1, 2023, from [62], and permits issued between January 1, 2023 and December 1, 2024, from [60].
+The total permitted site-level annual emission limits are approximately 13,000 tons of NOx, 1,400 tons of VOCs, 50 tons of SO2, and 600 tons of PM2.5, all in U.S.
+short tons.
+By assuming that the actual emissions are 10% of the permitted level, the data centers in Virginia could already cause approximately 14,000 asthma symptom cases and 13-19 deaths each year, among other health implications, resulting in a total annual public health burden of \$220-300 million, including \$190-260 million incurred in Virginia, West Virginia, Maryland, Pennsylvania, New York, New Jersey, Delaware, and Washington D.C., as estimated by COBRA under the “Fuel Combustion: Industrial” sector.
+The scope-1 emission information for data centers in other states is not always publicly available.
+Thus, when estimating scope-1 emissions for data centers in another state, we apply the emission rate (tons/MWh) derived from Virginia’s data and multiply it by the energy consumption of data centers in that states.
+Although this approach may introduce some estimation errors, the impact is expected to be limited because scope-2 health costs are substantially more dominant.
+A.2 Data Centers’ Scope-2 Public Health Impact
+The locations of emission sources depend on the power plants supplying electricity to data centers.
+To evaluate the public health impacts of U.S.
+data centers, we focus on average attribution method, which is also the standard methodology of carbon emission accounting used by technology companies in their sustainability reports [10, 78, 37].
+We first calculate the total data center electricity consumption $e_{DC}$ and the overall electricity consumption (including non-data center loads) $e_{Total}$ within each electricity region.
+The U.S.
+electricity grid is divided into 14 regions following the AVoided Emissions and geneRation Tool (AVERT, the latest version v4.3 as of October 2024) provided by the EPA [76].
+We use the state-level data center electricity consumption distribution for 2023 provided by EPRI [5], scale it by the U.S.
+total data center electricity consumption in 2019-2023 and for the 2028 projection based on data provided by [4], and then distribute state-level electricity consumption to relevant electricity regions following the state-to-region electricity apportionment used by AVERT.
+We calculate the percentage $x\%=\frac{e_{DC}}{e_{Total}}$ of the data center electricity consumption with respect to the overall electricity consumption for each electricity region.
+The relationship between the health impact and emission reduction in COBRA is approximately linear.
+Thus, we apply a reduction by $x\%$ to the baseline emissions of all the power plants within the respective electricity region in COBRA and estimate the corresponding county-level health impacts, including health outcomes and costs.
+When assessing the health impact of generative AI training, we follow the same approach, except for changing the total data center electricity consumption to the AI model training electricity consumption.
+A.3 Public Health Impact of a Semiconductor Facility
+Although semiconductor manufacturing facilities are subject to air quality regulations [74], they still pose significant risks, affecting populations across large regions.
+Maricopa County, AZ, has been an EPA-designated non-attainment area for several years due to its failures to meet federal air quality standards [106].
+The establishment of multiple semiconductor facilities in such areas could further exacerbate air quality issues.
+We consider a semiconductor manufacturing facility located in Ocotillo, a neighborhood in Chandler, Arizona [107].
+By averaging the rolling 12-month air pollutant emission levels listed in the recent air quality monitoring report (as of October, 2024) [18], we obtain the annual emissions as follows: 150.4 tons of NOx, 82.7 tons of VOCs, 1.1 tons of SO2, and 28.9 tons of PM2.5.
+By applying these on-site emissions to COBRA under the “Other Industrial Processes” sector, we obtain a total public health cost of \$14-21 million.
+Additionally, the total annual energy consumption by the facility is 2074.88 million kWh as of Q2, 2024 [107].
+Assuming 84.2% of the energy comes from the electricity based on the company’s global average [108], we obtain the facility’s annual electricity consumption as 1746.63 million kWh.
+By using the average attribution method, we further obtain an estimated health cost of \$12-17 million associated with the electricity consumption.
+Thus, the total health cost of the facility is \$26-39 million.
+By relocating the facility from Chandler, Arizona, to a planned site in Licking County, Ohio, and assuming the same emission level and electricity consumption, we can obtain the total health cost of \$94-156 million, including \$23-36 million attributed to direct on-site emissions and \$70-120 million attributed to electricity consumption.
+A.4 Energy Consumption for Training a Generative AI Model
+We consider Llama-3.1 as an example generative AI model.
+According to the model card [36], the training process of Llama-3.1 (including 8B, 70B, and 405B) utilizes a cumulative of 39.3 million GPU hours of computation on H100-80GB hardware, and each GPU has a thermal design power of 700 watts.
+Considering Meta’s 2023 PUE of 1.08 [37] and excluding the non-GPU overhead for servers, we estimate the total training energy consumption as approximately 30 GWh.
+Our estimation method follows Meta’s guideline [36] and is conservative, as it excludes the substantial non-GPU energy overheads (e.g., CPUs) associated with server operations.
+A.5 Average Emission for Each LA-NYC Round Trip by Car
+We use the 2023 national average emission rate for light-duty vehicles (gasoline) provided by the U.S.
+Department of Transportation [104].
+The emission rate accounts for tailpipe exhaust, tire wear and brake wear.
+Specifically, the average PM2.5 emission rate is 0.008 grams/mile (including 0.004 grams/mile for exhaust, 0.003 grams/mile for brake wear, and 0.001 grams/mile for tire wear), and the average NOx emission rate is 0.199 grams/mile for exhaust.
+We see that half of PM2.5 for light-duty vehicles comes from brake and tire wear (0.004 gram/miles), which are also produced by other types of vehicles including electric vehicles.
+The distance for a round-trip between Los Angeles, California, and New York City, New York, is about 5,580 miles.
+Thus, the average auto emissions for each LA-NYC round trip are estimated as 44.64 grams of PM2.5 and 1110.42 grams of NOx.
+A.6 State-wide Electricity Consumption by U.S. Data Centers in 2023
+We show in Fig.
+10 the state-wide data center electricity consumption in 2023 [5].
+It can be seen that Virginia, Texas and California have the highest data center electricity consumption in 2023.
+The total national electricity consumption reported by EPRI is slightly lower than the values in [4], and we scale it up accordingly in our calculations to ensure consistency.
+(a) Electricity consumption map
+State Electricity Consumption (TWh) State Electricity Consumption (TWh) State Electricity Consumption (TWh) VA 33.85 OH 2.36 ID 0.15 TX 21.81 SC 2.02 WI 0.15 CA 9.33 WY 1.86 MD 0.10 IL 7.45 KY 1.62 LA 0.08 OR 6.41 CO 1.51 SD 0.07 AZ 6.25 AL 1.49 ME 0.03 IA 6.19 FL 1.38 NH 0.02 GA 6.18 TN 1.33 RI 0.02 WA 5.17 OK 1.23 KS << 0.01 PA 4.59 MA 1.06 AR << 0.01 NY 4.07 MO 0.97 DE << 0.01 NJ 4.04 MN 0.82 DC << 0.01 NE 3.96 MT 0.58 MS << 0.01 ND 3.92 MI 0.53 VT << 0.01 NV 3.42 NM 0.40 WV << 0.01 NC 2.67 CT 0.26 UT 2.56 IN 0.19
+(b) Electricity consumption by state (descending order)
+Figure 10: State-level electricity consumption of U.S. data centers in 2023 [5].
+Appendix B Additional Results for Health-Informed GLB
+B.1 Details of the experiment setup
+We use Meta’s electricity consumption for each U.S.
+data center location in 2023 [37] for our experiments.
+Table 7 summarizes the baseline annual energy load $W_{i}$ for each data center $i$.
+Since data centers are mostly stable loads in practice, the hourly workload for each location is calculated as $l_{i}=\frac{W_{i}}{T}$, where $T$ represents the total number of hours in the study period.
+The total hourly workload is then computed as $M_{t}=\sum_{i=1}^{N}l_{i}$.
+We use the annual average industrial electricity prices in different states provided by the EIA [109].
+The health price $p_{i,t}^{h}$ (\$/MWh) and carbon emission rate $r_{i,t}^{c}$ (ton/MWh) are based on data provided by WattTime [77].
+WattTime divides the U.S.
+into more than 100 regions and provides marginal health prices and carbon emission rates for each region.
+These values are updated every 5 minutes.
+In our study, we use the WattTime data from 0:00 on January 1, 2023, to 23:55 on December 31, 2023.
+The actual time interval for our experiment is 1 hour, where the 5-minute data points are averaged hourly to compute $p_{i,t}^{h}$ and $r_{i,t}^{c}$, $\forall t=\{1,2,\ldots,T\}$, with $T=8760$.
+B.2 Sensitivity analysis of capacity slackness $\lambda$
+The parameter $\lambda\geq 1$ represents the capacity slackness to accept additional loads: The greater $\lambda$, the more spatial flexibility there is.
+Here, we vary $\lambda$ and report the results for $\lambda=1.2$ and $\lambda=2.0$ in Tables 8 and 9, respectively.
+Similar to Table 4, the potential reduction in health costs is significant.
+An increase in $\lambda$ corresponds to an expansion in the spatial flexibility, leading to a larger reduction in health cost.
+Specifically, when $\lambda=2.0$, the health-informed algorithm achieves a reduction in health cost of over 40% compared to the baseline.
+In contrast, under $\lambda=2.0$, the pure carbon-aware algorithm, which focuses exclusively on optimizing carbon emissions with an infinite carbon price $p^{c}=\infty$, can achieve a reduction in carbon emissions of approximately 12% relative to the baseline.
+This comparison highlights the great potential for health cost reduction and reinforces the importance of designing HI-GLB.
+Table 7: Information about Meta’s U.S. data center locations in 2023
+Location Energy (MWh) Electricity Price $p^{e}$ (\$/MWh) Health Price $p^{h}$ (\$/MWh) Carbon Intensity $r^{c}$(ton/MWh)
+Huntsville, AL 614198 71.0 49.49 0.59
+Stanton Springs, GA 968565 68.8 19.35 0.58
+DeKalb, IL 138965 82.0 51.72 0.58
+Altoona, IA 1243306 69.1 57.86 0.65
+Sarpy, NE 1148091 76.3 36.72 0.56
+Los Lunas, NM 1110100 57.5 18.96 0.68
+Forest City, NC 507068 71.5 54.64 0.66
+New Albany, OH 793063 70.3 54.32 0.61
+Prineville, OR 1375321 75.2 13.31 0.65
+Gallatin, TN 116520 62.3 49.49 0.59
+Fort Worth, TX 1029570 66.0 44.56 0.51
+Eagle Mountain, UT 787740 69.9 21.93 0.74
+Henrico, VA 805061 89.2 54.37 0.61
+Table 8: Comparison between health-informed and carbon-aware GLB ($\lambda=1.2$)
+Metric Baseline Carbon-Aware GLB Health-Informed GLB
+$p^{c}=$ \$0/ton $p^{c}=$ \$5/ton $p^{c}=$ \$200/ton $p^{c}=$ $\infty$/ton HI-GLB ($p_{i,t}^{e}=0$) HI-GLB
+Health (Million \$) 393.23 374.20 (-4.84%) 374.44 (-4.78%) 395.28 (0.52%) 404.72 (2.92%) 345.26 (-12.20%) 349.29 (-11.17%)
+Energy (Million \$) 756.50 732.06 (-3.23%) 732.30 (-3.20%) 747.88 (-1.14%) 761.20 (0.62%) 752.26 (-0.56%) 738.49 (-2.38%)
+Carbon (Million Ton) 6.60 6.68 (1.22%) 6.56 (-0.60%) 6.40 (-3.04%) 6.37 (-3.54%) 6.54 (-0.87%) 6.57 (-0.44%)
+Table 9: Comparison between health-informed and carbon-aware GLB ($\lambda=2.0$)
+Metric Baseline Carbon-Aware GLB Health-Informed GLB
+$p^{c}=$ \$0/ton $p^{c}=$ \$5/ton $p^{c}=$ \$200/ton $p^{c}=$ $\infty$/ton HI-GLB ($p_{i,t}^{e}=0$) HI-GLB
+Health Cost (Million \$) 393.23 368.10 (-6.39%) 409.67 (4.18%) 374.82 (-4.68%) 408.92 (3.99%) 221.38 (-43.70%) 223.96 (-43.05%)
+Energy Cost (Million \$) 756.50 702.16 (-7.18%) 702.69 (-7.11%) 723.42 (-4.37%) 767.59 (1.47%) 734.29 (-2.94%) 728.19 (-3.74%)
+Carbon (Million Ton) 6.60 6.72 (1.73%) 6.52 (-1.17%) 5.95 (-9.91%) 5.84 (-11.49%) 6.61 (0.19%) 6.55 (-0.71%)